### v1

# Sampling 3D Molecular Conformers with Diffusion Transformers

Abstract
Diffusion Transformers (DiTs) have demonstrated strong performance in generative modeling, particularly in image synthesis, making them a compelling choice for molecular conformer generation. However, applying DiTs to molecules introduces novel challenges, such as integrating discrete molecular graph information with continuous 3D geometry, handling Euclidean symmetries, and designing conditioning mechanisms that generalize across molecules of varying sizes and structures. We propose DiTMC, a framework that adapts DiTs to address these challenges through a modular architecture that separates the processing of 3D coordinates from conditioning on atomic connectivity. To this end, we introduce two complementary graph-based conditioning strategies that integrate seamlessly with the DiT architecture. These are combined with different attention mechanisms, including both standard non-equivariant and SO(3)-equivariant formulations, enabling flexible control over the trade-off between between accuracy and computational efficiency. Experiments on standard conformer generation benchmarks (GEOM-QM9, -DRUGS, -XL) demonstrate that DiTMC achieves state-of-the-art precision and physical validity. Our results highlight how architectural choices and symmetry priors affect sample quality and efficiency, suggesting promising directions for large-scale generative modeling of molecular structures. Code available at https://github.com/ML4MolSim/dit_mc.
Keywords Generative Modeling  $\cdot$ Flow Matching  $\cdot$ Diffusion Transformers  $\cdot$ Geometric Deep Learning  $\cdot$ Molecular Simulation  $\cdot$ Conformer Generation
1 Introduction
The three-dimensional arrangement of atoms in a molecule, known as conformation, determines its biological activity and physical properties, making it fundamental to applications such as computational drug discovery and material design. Accurately predicting the most energetically favorable conformers (i.e., stable conformations) for large molecular systems is a highly non-trivial task. Traditional techniques, such as Molecular Dynamics and Markov Chain Monte Carlo, attempt to explore the conformational space by simulating physical movement or probabilistic sampling. However, these methods often require many simulation steps to move from one conformer to another, making them computationally expensive. Generative machine learning (ML) models offer a more targeted approach by allowing to directly sample from the space of promising conformations.
Recent years have seen significant progress, enabled by the development of specialized architectures for the generation of molecules [1, 2, 3, 4, 5, 6, 7, 8, 9] and materials [10, 11, 12]. This is in contrast to image and video synthesis, where the more generalized diffusion transformer (DiT) architecture [13] has become a leading model, consistently delivering strong performance and efficiency across diverse applications [14, 15, 16, 17]. Adapting DiTs, which were originally developed for grid-structured image data, to continuous, irregular molecular geometries poses unique challenges, which need to be addressed to unlock the potential of this powerful architecture for molecular conformer generation and leverage prior advances from vision research. Key design questions include how to encode molecular connectivity and incorporate Euclidean symmetries, such as translational and rotational invariance/equivariance.
In this work, we address these conceptual challenges and propose DiTMC, a new DiT-style architecture for molecular conformer generation. We introduce novel conditioning strategies based on molecular graphs, enabling 3D conformer generation. Our modular architectural design allows us to systematically investigate the impact of different self-attention mechanisms within the DiT architecture on conformer generation quality and efficiency. We conduct a comparative study including standard (non-equivariant) self-attention with both absolute and relative positional embeddings and an explicitly $\mathrm{SO}(3)$-equivariant variant. While exact equivariance can positively impact performance, it also incurs significant computational costs. We find that simpler attention mechanisms are highly scalable and still perform competitively. Multiple of the tested DiT variants achieve state-of-the-art (SOTA) precision on established conformer generation benchmarks. Further, the ensembles of molecular structures generated by our models exhibit increased realism, as evidenced by the high accuracy of physical properties extracted from them. To summarize, our work contains the following main contributions:
  • We propose two complementary conditioning strategies based on trainable conditioning tokens for (pairs of) atoms extracted from molecular graphs, which are designed to align with the architectural principles of DiTs.
  • We investigate the impact of different self-attention mechanisms, including standard (non-equivariant) and SO(3)-equivariant formulations, on model accuracy and performance. We find that including symmetries can improve the fidelity of generated samples at the price of increased computational cost during training and inference.
  • Based on our insights, we present a simple, non-equivariant, yet expressive DiT architecture that achieves state-of-the-art (SOTA) precision and physical validity on established benchmarks. Its performance improves with model scaling, making it a promising candidate for large-scale molecular conformer generation.
Figure 1: (A) Diffusion transformer for molecular conformer generation (DiTMC), with interchangeable self-attention blocks and positional embeddings (PEs); we evaluate various combinations as detailed in the main text. (B) DiTMC predicts a velocity per atom, used to model a probability flow ODE, which samples from the probability distribution $p(\boldsymbol{x}|\mathcal{G})$, where $\mathcal{G}$ is a molecular graph.
2 Related Work
Generative Modeling
Generative models create diverse, high-quality samples from an unknown data distribution. They are widely used in image generation [18, 19], text synthesis [20] and the natural sciences [21, 22, 23]. Most recent approaches learn a probabilistic path from a simple prior to the data distribution, enabling both efficient sampling and likelihood estimation [24, 25, 26, 27, 28]. This is achieved by different modeling paradigms, including flow matching [26] or denoising diffusion [24].
(Diffusion) Transformers
Originally introduced for natural language processing [29], transformer architectures have become SOTA in domains such as computer vision [18], and recently also found widespread adoption in quantum chemistry, e.g., for protein prediction [30, 31], 3D molecular generation [32, 33] or molecular dynamics simulations [34, 35, 36]. In the context of generative modeling, diffusion transformers [13] (DiTs) have emerged as powerful tools incorporating conditioning tokens, e.g., to prompt image generation [13] or to design molecules and materials with desirable properties [37]. This work applies prototypical DiTs to molecular conformer generation.
Molecular Conformer Generation
Molecular conformer generation aims to find atomic arrangements (Cartesian coordinates) consistent with a given molecular graph. Numerous ML approaches have been proposed for sampling conformers, all aiming to improve upon conventional methods. While early approaches were based on RDKit [38] or variational auto encoders [39, 40, 41], more recent advances employ diffusion and flow-based models [42, 43], including E-NFs [9], CGCF [44], GeoDiff [45], TorsionDiff [46], MCF [47], ET-Flow [32], and DMT [48].
3 Preliminaries
3.1 Molecular Conformers
A molecule can be associated with a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where the nodes $\mathcal{V}$ correspond to atoms and the edges $\mathcal{E}$ represent chemical bonds between them. The nodes and edges contain information about their types, bond orders, and additional structural features such as branches, rings, and stereochemistry. What is missing is the exact spatial arrangement of the $N=\lvert\mathcal{V}\rvert$ atoms as a 3D point cloud $\boldsymbol{x}\in\mathbb{R}^{N\times 3}$ in Euclidean space. Only the relative distances between atoms are relevant, as translating or rotating the entire point cloud $\boldsymbol{x}$ does not change the identity of the conformer. We frame conformer prediction as sampling from the conditional probability distribution $p(\boldsymbol{x}\,|\,\mathcal{G})$, which will guide the design of our model architectures (Sec. 4).
3.2 Conditional Flow Matching
Starting from an easy-to-sample base distribution $q_{0}:\mathbb{R}^{d}\mapsto\mathbb{R}_{\geq 0}$, a generative process creates samples from a target distribution $q_{1}:\mathbb{R}^{d}\mapsto\mathbb{R}_{\geq 0}$ [27]. Here, $q_{1}$ models the molecular conformer data with $d=N\times 3$. We aim to learn a time-dependent vector field $u_{t}(\boldsymbol{x}):[0,1]\times\mathbb{R}^{N\times 3}\mapsto\mathbb{R}^{N\times 3}$, which defines an ordinary differential equation (ODE) whose solution pushes samples $\boldsymbol{x}_{0}\in\mathbb{R}^{N\times 3}$ from the prior to samples $\boldsymbol{x}_{1}\in\mathbb{R}^{N\times 3}$ from the data distribution.
We describe this transformation in terms of a stochastic interpolant $\boldsymbol{x}_{t}$ [49, 26, 50]. A noisy sample at time $t\in[0,1]$ is defined as
$\boldsymbol{x}_{t}=(1-t)\cdot\boldsymbol{x}_{0}+t\cdot\boldsymbol{x}_{1}+\sigma\cdot\boldsymbol{\epsilon},$ (1)
where $\boldsymbol{\epsilon}\in\mathbb{R}^{N\times 3}$ is drawn from the standard normal distribution $\mathcal{N}(0,\boldsymbol{I})$ and scaled by a constant $\sigma\in\mathbb{R}_{\geq 0}$. We remark that $t$ represents progress along this interpolation path, not physical time. Notably, stochastic interpolants enable transformations between arbitrary distributions and allow us to assess the performance of the generative process under varying prior distributions $q_{0}$. This contrasts with, e.g., score based diffusion methods [24, 51], which typically assume an isotropic Gaussian prior.
The stochastic interpolant induces a deterministic trajectory of densities $p_{t}(\boldsymbol{x})$, governed by an ODE known as the probability flow:
$\mathrm{d}\boldsymbol{x}=u_{t}(\boldsymbol{x})\,\mathrm{d}t.$ (2)
If the vector field $u_{t}(\boldsymbol{x})$ was tractable to sample, the weights of a neural network (NN) $\boldsymbol{v}^{\theta}(\boldsymbol{x},t):[0,1]\times\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ could be optimized directly by minimizing
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1),\boldsymbol{x}\sim p_{t}(\boldsymbol{x})}\ \Big{|}\Big{|}u_{t}(\boldsymbol{x})-\boldsymbol{v}^{\theta}(\boldsymbol{x},t)\Big{|}\Big{|}.$ (3)
The learned vector field $\boldsymbol{v}^{\theta}$ could then be used to generate new samples from the target distribution by starting from $\boldsymbol{x}_{0}\sim q_{0}$ and integrating the probability flow ODE (Eq. 2), for example, using a numerical scheme such as Euler’s method, i.e., $\boldsymbol{x}_{t+\Delta t}=\boldsymbol{x}_{t}+\boldsymbol{v}^{\theta}(\boldsymbol{x}_{t},t)\Delta t$ for time step $\Delta t$.
However, for arbitrary distributions $q_{0}$ and $q_{1}$, the objective in Eq. 3 is computationally intractable [52]. Instead, we consider the expectation over interpolated point pairs from the two distributions. Eq. 1 defines a conditional probability distribution $p_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}|(1-t)\cdot\boldsymbol{x}_{0}+t\cdot\boldsymbol{x}_{1},\sigma^{2})$, with conditional vector field $\boldsymbol{u}_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})=\boldsymbol{x}_{1}-\boldsymbol{x}_{0}$ [27]. The ability to directly sample from the conditional probability via Eq. 1 allows formulating the conditional flow matching (CFM) objective
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1),\boldsymbol{x}_{0}\sim q_{0},\boldsymbol{x}_{1}\sim q_{1},\boldsymbol{x}\sim p_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})}\Big{|}\Big{|}\boldsymbol{u}_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})-\boldsymbol{v}^{\theta}(\boldsymbol{x},t)\Big{|}\Big{|}^{2}.$ (4)
As shown in Ref. [26], the gradients of the two losses coincide, $\nabla_{\theta}\mathcal{L}_{\text{FM}}=\nabla_{\theta}\mathcal{L}_{\text{CFM}}$, thereby recovering the vector field that defines the probability flow ODE in Eq. 2. Following prior work [53, 37], we reparametrize the training objective to predict noise-free data $\boldsymbol{x}_{0}^{\theta}$ directly. During inference, we invert the reparametrization to obtain $\boldsymbol{v}^{\theta}$ for sampling (see Appendix C).
4 A New Diffusion Transformer for Molecular Conformer Sampling
We now describe the methodological advances of our work. We propose DiTMC, a new DiT-style architecture for conformer generation by learning a vector field using the loss in Eq. 4. As outlined in Sec. 3.1, this involves sampling from a conditional probability $p(\boldsymbol{x}|\mathcal{G})$, where $\mathcal{G}$ is the molecular graph representing atomic connectivity. Therefore, we choose our model to be a function $\boldsymbol{v}^{\theta}(\boldsymbol{x},t,\mathcal{G})$.
We begin by initializing the node features using learnable embeddings $\boldsymbol{e}(z_{i})\in\mathbb{R}^{H}$, where $z_{i}\in\mathbb{N}_{+}$ denotes the atomic number of atom $i$ [35]. These embeddings are then processed through multiple DiT blocks. Each DiT block transforms the current node features (input tokens) =$\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$ to a set of output tokens $\mathcal{H}^{\prime}$, which serve as input for the next block. Additional inputs consist of positional embeddings $\mathcal{P}$ (representing the position of nodes in Euclidean space), as well as conditioning tokens $\mathcal{C}$, for example for encoding the time ($\boldsymbol{c}^{t}$) and graph structure ($\boldsymbol{c}^{\mathcal{G}}$).
The central operation in each DiT block is a self-attention update,
$\boldsymbol{h}_{i}=\boldsymbol{h}_{i}+\text{ATT}(\mathcal{H},\mathcal{P},\mathcal{C}^{\text{Pair}})_{i}\,,$ (5)
followed by a node-wise Multi-Layer Perceptron (MLP). Here, $\mathcal{C}^{\text{Pair}}$ are special conditioning tokens acting on pairs of atoms (see Sec. 4.1). Similar to applications of DiTs in image synthesis, adaptive layer norm (AdaLN) and adaptive scale (AdaScale) [13] are used for conditioning (see Fig. 1A) based on per-atom bias and scaling parameters
$\alpha_{1i},\beta_{1i},\gamma_{1i},\alpha_{2i},\beta_{2i},\gamma_{2i}=\text{MLP}(\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}})\,.$ (6)
The final output is a 3D vector (“velocity”) per atom, which is extracted from a readout layer. The whole architecture is summarized in Fig. 1A. In the following, we describe the individual components of the architecture in more detail. Training details are given in Appendix A.
4.1 Conditioning
Each DiT block receives conditioning signals that encode information about the time step $t$, as well as atom- and pair-wise information derived from a molecular graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$. Pair-wise information is used during the self-attention update (see Eq. 5 and Sec. 4.3), whereas atom-wise information and the time conditioning signal are injected via adaptive layer norms (see Eq. 6).
Time conditioning
The current time $t$ of the latent state $\boldsymbol{x}_{t}$ is encoded via a two-layer MLP as
$\boldsymbol{c}^{t}=\text{MLP}(t)\,.$ (7)
Atom-wise conditioning
Atom-wise conditioning tokens are obtained from a GNN inspired by the processor module of the MeshGraphNet (MGN) framework [54]
$\boldsymbol{c}_{i}^{\mathcal{G}}=\text{GNN}_{\text{node}}(\mathcal{V},\mathcal{E})\,,$ (8)
where $\boldsymbol{c}_{i}^{\mathcal{G}}$ denotes the final node representation for atom $i$. See Appendix I for details on the GNN.
Pair-wise conditioning
In addition to node representations, the GNN also produces edge-level representations, which can be used to define pair-wise conditioning tokens:
[…]
These tokens only capture interactions between bonded atoms, i.e., when $(i,j)\in\mathcal{E}$. Conditioning tokens for non-bonded pairs are set to a learnable vector $\boldsymbol{\bar{c}}^{\mathcal{G}}$. Since self-attention operates on all atom pairs $(i,j)$, even if they are not connected by a chemical bond, we also define an alternative pair-wise conditioning method inspired by the Graphormer architecture [55]
[…]
where $s(i,j)$ denotes the graph geodesic (i.e., the shortest path between atoms $i$ and $j$ via edges in $\mathcal{G}$). This formulation allows conditioning on all atom pairs, even if they are not directly connected. In the following, we refer to these two variants as bond-pair (Eq. 9) and all-pair (Eq. 10) conditioning, respectively (see Sec. 5 for empirical comparisons between both methods).
4.2 Positional Embeddings ( PE )
We inject information about the current atomic positions =$\mathcal{R}=\{\vec{r}_{1},\dots\vec{r}_{N}\,|\,\vec{r}_{i}\in\mathbb{R}^{3}\}$ of the latent state $\boldsymbol{x}_{t}$ via positional embeddings (PEs). To that end, we assume the positions to have zero center of mass (see section A.1). As mentioned in Sec. 3.1, conformer identity is determined only by the relative positions of atoms (i.e., it is invariant under rigid translations and rotations), but ensuring that these symmetries are preserved within the network often incurs computational overhead.
Therefore, we examine a representative range of PEs that vary in the number of Euclidean symmetries they respect by construction, which affects how the latent representations transform under translations and rotations. This includes absolute and relative PEs, motivated by a similar distinction made in other application domains [56, 57]. Inspired by recent advances in geometric deep learning [58, 59, 60], we also propose a PE strategy, obeying all Euclidean symmetries. Our aim is to find an acceptable tradeoff between model accuracy, efficiency and scalability, by scanning the design space of PE strategies.
Absolute Positional Embeddings ( aPE )
Following Refs. [31, 37] they are calculated as
$\boldsymbol{p}_{i}^{\text{aPE}}=\text{MLP}(\vec{r}_{i})\,,$ (11)
where $\vec{r}_{i}\in\mathbb{R}^{3}$ is the Cartesian position of the $i$-th atom. This kind of PE is neither rotationally nor translationally invariant, and serves as a baseline without any symmetry constraints.
Relative Positional Embedding ( rPE )
We use displacements vectors $\vec{r}_{ij}=\vec{r}_{i}-\vec{r}_{j}$ to build pairwise relative PEs as
$\boldsymbol{p}_{ij}^{\text{rPE}}=\text{MLP}(\vec{r}_{ij})\,.$ (12)
This formulation ensures translational invariance but not rotational invariance.
Euclidean Positional Embedding ( PE(3) )
Adapting ideas from equivariant message passing neural networks like PaiNN [61] or NequIP [62], we construct SO(3)-equivariant pairwise PEs as a concatenation of $L+1$ components
$\boldsymbol{p}_{ij}^{\text{PE(3)}}=\bigoplus_{\ell=0}^{L}\boldsymbol{\phi}_{\ell}(r_{ij})\,\odot\boldsymbol{Y}_{\ell}(\hat{r}_{ij})\,,$ (13)
where $\boldsymbol{\phi}_{\ell}:\mathbb{R}\mapsto\mathbb{R}^{1\times H}$ is a radial filter function, $\hat{r}=\vec{r}/r$, and $\boldsymbol{Y}_{\ell}\in\mathbb{R}^{(2\ell+1)\times 1}$ are spherical harmonics of degree $\ell=0\dots L$. The element-wise multiplication $\odot$’ between radial filters and spherical harmonics is understood to be “broadcasting” along axes with size $1$, such that $(\boldsymbol{\phi}_{\ell}\,\odot\boldsymbol{Y}_{\ell})\in\mathbb{R}^{(2\ell+1)\times H}$ and (after concatenation) $\boldsymbol{p}_{ij}^{\text{PE(3)}}\in\mathbb{R}^{(L+1)^{2}\times H}$. Under rotation of the input positions, these PEs transform equivariantly (see Appendix B.2.2). Moreover, because displacement vectors are used as inputs, the embeddings are also invariant to translations. As a result, they respect the full set of Euclidean symmetries relevant to molecular geometry.
4.3 Self-Attention Operation
For ease of notation, we only describe self-attention with a single head, but employ multi-head attention [29] with $n_{\text{heads}}$ heads in our experiments. All self-attention blocks rely on query, key and value vectors, which are obtained from the input tokens =$\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\ |\ \boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$ as
$\boldsymbol{q}=\boldsymbol{W}_{q}\boldsymbol{\tilde{h}}\,,\hskip 25.0pt\boldsymbol{k}=\boldsymbol{W}_{k}\boldsymbol{\tilde{h}}\,,\hskip 25.0pt\boldsymbol{v}=\boldsymbol{W}_{v}\boldsymbol{\tilde{h}}\,,$ (14)
where $\boldsymbol{W}_{q},\boldsymbol{W}_{k},\boldsymbol{W}_{v}\in\mathbb{R}^{H\times H}$ are trainable weight matrices and $\boldsymbol{\tilde{h}}$ is either identical to $\boldsymbol{h}$ or combines it with a PE (see below). We define a slightly modified similarity kernel
$\text{sim}(\boldsymbol{q},\boldsymbol{k},\boldsymbol{u})=\exp\Bigg{(}\frac{\boldsymbol{q}^{\intercal}\cdot(\boldsymbol{k}\,\odot\boldsymbol{u})}{\sqrt{H}}\Bigg{)},$ (15)
where $\boldsymbol{u}\in\mathbb{R}^{H}$ is used to inject additional information, e.g., conditioning signals and/or positional embeddings, and $\odot$’ denotes element-wise multiplication. Depending on the subset of Euclidean symmetries we aim to incorporate, we adopt slightly different formulations of the self-attention mechanism, as detailed below.
Standard Self-Attention
For absolute and relative PEs, we slightly modify standard self-attention to allow injecting pair-wise information into the values in addition to using our modified similarity kernel:
$\text{ATT}(\mathcal{H})_{i}=\frac{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})\cdot(\boldsymbol{v}_{j}\odot\boldsymbol{u}_{ij})}{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})}\,.$ (16)
Queries, keys, and values are obtained with Eq. 14 from different (position-encoded) tokens $\boldsymbol{\tilde{h}}_{i}$ depending on the chosen PEs; further, the injected pair-wise information $\boldsymbol{u}_{ij}$ differs:
[…]
Here $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$ are pair-wise graph conditioning tokens (see Eqs. 9 and 10) and $\boldsymbol{p}_{i}^{\text{aPE}}\in\mathbb{R}^{H}$ and $\boldsymbol{p}_{ij}^{\text{rPE}}\in\mathbb{R}^{H}$ are the absolute and relative PEs described above (see Eqs. 11 and 12).
SO(3)-Equivariant Self-Attention
To preserve all Euclidean symmetries throughout the network, every token must transform equivariantly. One way to achieve this is by separating out the rotational degrees of freedom, encoding them with irreducible representations of the rotation group $\mathrm{SO(3)}$. This introduces a “degree-axis” of size $(L+1)^{2}$, which encodes angular components of increasing order. The maximum degree $L$ is chosen to ensure high fidelity at a reasonable computational cost. For example, setting $L=1$ restricts the representation to scalars and vectors, as used in models like PaiNN [61] or TorchMDNet [34]. An $\mathrm{SO(3)}$-equivariant formulation of self-attention is then given as
$\text{ATT}_{\text{SO(3)}}(\mathcal{H})_{i}=\frac{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})\cdot(\hat{\boldsymbol{u}}_{ij}\otimes\boldsymbol{v}_{j})}{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})}\,,$ (18)
where equivariant queries, keys and values can be calculated similarly to Eq. 14 and $\otimes$’ denotes a Clebsch-Gordan (CG) tensor product contraction [60]. The dot-product in the similarity measure is taken along both feature and degree axes, such that the overall update preserves equivariance (see Appendix for details). Tokens and scaling vectors are calculated as
[…]
where $\boldsymbol{\phi}(r_{ij})\in\mathbb{R}^{(L+1)^{2}\times H}$ is a radial filter, and the element-wise products with the pair-wise conditioning tokens $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{1\times H}$ are broadcast along the degree axis. Importantly, the $12\ell+$ 1 subcomponents of the radial filter for degree $\ell$ are obtained by repeating per-degree filter functions $\boldsymbol{\phi}_{\ell}(r_{ij})\in\mathbb{R}^{1\times H}$ along the degree axis to preserve equivariance (see also Eq. 13).
5 Experiments
Datasets and Metrics
We conduct our experiments on the GEOM dataset [63], comprising QM9 (133,258 small molecules) and AICures (304,466 drug-like molecules). Reference conformers are generated using CREST [64]. Drug-like molecules exhibit greater structural diversity, including more rotatable bonds and multiple stereocenters. Data splits are taken from Ref. [65].
We evaluate our models’ ability to generate accurate and diverse conformers using Average Minimum RMSD (AMR) and Coverage (COV), measuring Recall (ground-truth coverage) and Precision (generation accuracy). A generated conformer is considered valid if it falls within a specified RMSD threshold of any reference conformer ($\delta=0.5\mbox{\AA}$ for GEOM-QM9 and $\delta=0.75\mbox{\AA}$ for GEOM-DRUGS). Following prior work, we generate $2K$ conformers per test molecule with $K$ reference structures. Appendix D.3 provides further details on the calculation of metrics. Following ET-Flow [32] and GeomMol [65] we also apply chirality correction (see Appendix D.4).
Conditioning Strategies
To assess the impact of different graph conditioning strategies, we compare “node only”, “node & bond-pair”, and “node & all-pair” conditioning (see Sec. 4.1 for details) against a baseline without graph conditioning. We find all variants to be effective, with “node & all-pair” conditioning performing best (see Appendix Tab. A8). It outperforms both, “node only” and “node & bond-pair” conditioning, highlighting the strength of our all-pair-conditioning strategy based on graph geodesics. This underlines the importance of deriving conditioning tokens for all atom pairs, not just those connected by edges in the molecular graph, which lack global information about the graph structure.
Table 1: Results on GEOM-QM9 for different generative models (parameter counts in parentheses). -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined; our models are marked with an asterisk (*). Our results are averaged over three random seeds. See Tab. A9 for results including standard deviations.
Method COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
Mean Median Mean Median Mean Median Mean Median
GeoMol (0.3M) 91.5 100.0 0.225 0.193 86.7 100.0 0.270 0.241
GeoDiff (1.6M) 76.5 100.0 0.297 0.229 50.0 33.5 0.524 0.510
Tors. Diff. (1.6M) 92.8 100.0 0.178 0.147 92.7 100.0 0.221 0.195
MCF-B (64M) 95.0 100.0 0.103 0.044 93.7 100.0 0.119 0.055
DMT-B (55M) 95.2 100.0 0.090 0.036 93.8 100.0 0.108 0.049
ET-Flow (8.3M) 96.5 100.0 0.073 0.030 94.1 100.0 0.098 0.039
*DiTMC+aPE-B (9.5M) 96.1 100.0 0.073 0.030 95.4 100.0 0.085 0.037
*DiTMC+rPE-B (9.6M) 96.3 100.0 0.070 0.027 95.7 100.0 0.080 0.035
*DiTMC+PE(3)-B (8.6M) 95.7 100.0 0.068 0.021 93.4 100.0 0.089 0.032
Figure 2: Analysis of equivariant (PE(3)) and non-equivariant (aPE, rPE) model formulations. (A) Mean coverage recall (COV-R) versus threshold $\delta$ (root mean square deviation (RMSD) to any reference conformer). (B) Histogram of the minimal RMSD per generated sample. (C) Loss as a function of latent time $t$ relative to PE(3) loss.
Ablating Self-Attention and Positional Embedding Strategies
The modular structure of DiTMC allows efficient exploration of the design space through variations in the PE strategy and associated attention blocks (see Sec. 4). All DiTMC models share hyperparameters, differing only in the choice of PEs and self-attention formulation. For the architecture using PE(3), we reduce the number of heads to match the parameter count across models. Tab. 1 shows results on GEOM-QM9. All DiTMC models produce diverse, high quality samples, outperforming the current SOTA in all AMR metrics and COV-P. This demonstrates the broad applicability of our modular design and conditioning. We use the harmonic prior introduced in Ref. [66] throughout, which yields improved results (see Appendix Tab. A15). Therefore, we use it in all following experiments.
Probing the effect of Euclidean symmetries
Our PEs form a hierarchy based on the extent of Euclidean symmetry incorporated by construction. This enables a systematic evaluation of how incorporating symmetry affects model behavior. We summarize our findings below.
Equivariance improves the fidelity of samples. We compute COV-R as a function of threshold $\delta$ for different PEs and self-attention blocks (Fig. 2A). SO(3)-equivariant attention with PE(3) outperforms the other variants at low $\delta$, indicating that many of the generated geometries closely match the ground-truth structures. This appears as a leftward shift in the distribution of the minimal RMSD found per generated structure (Fig. 2B) and aligns with the observation that the PE(3) model gives better AMR-R and AMR-P values (see Tab. 1). To better understand this behavior, we analyze the loss over time $t$ and find that models with non-equivariant PEs exhibit higher error near the data distribution ($t=1$) (Fig. 2C). The increase in error towards the end of the generation trajectory results in noisier structures and reduced fidelity.
Equivariance increases the computational cost for models of similar size. The benefit of higher fidelity comes at increased computational cost. During training, the equivariant PE(3) model is approximately $3.5$ times slower than models using aPE or rPE, while at inference time, the factor is $\sim 3$ (see Appendix Tab. A5). All models use the same number of layers and differ only in the number of heads per layer in order to match the total parameter count.
Table 2: Results on GEOM-DRUGS for different generative models (parameter counts in parentheses). -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined; our models are marked with an asterisk (*). Our results are averaged over three random seeds. See Tab. A10 for results including standard deviations.
Method COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
Mean Median Mean Median Mean Median Mean Median
GeoMol (0.3M) 44.6 41.4 0.875 0.834 43.0 36.4 0.928 0.841
GeoDiff (1.6M) 42.1 37.8 0.835 0.809 24.9 14.5 1.136 1.090
Tors. Diff. (1.6M) 72.7 80.0 0.582 0.565 55.2 56.9 0.778 0.729
MCF-L (242M) 84.7 92.2 0.390 0.247 66.8 71.3 0.618 0.530
DMT-L (150M) 85.8 92.3 0.375 0.346 67.9 72.5 0.598 0.527
ET-Flow - SS (8.3M) 79.6 84.6 0.439 0.406 75.2 81.7 0.517 0.442
*DiTMC+aPE-L 79.2 84.4 0.432 0.386 77.8 85.7 0.470 0.387
*DiTMC+PE(3)-L 80.8 85.6 0.415 0.376 76.4 82.6 0.491 0.414
Figure 3: (A) Time per gradient step for the DiTMC S, M and L variants using aPE and PE(3), measured on an Nvidia H100 GPU (batch size 128 on Geom QM9). (B) Coverage precision and recall as function of threshold $\delta$ and (C) mean AMR metrics for base (“B”) and large (“L”) model of DiTMC+aPE.
Drug-like Molecules
Finally, we evaluate our architecture on the challenging GEOM-DRUGS dataset. We analyze its behavior when scaling model capacity, assess physical validity via ensemble property prediction, and examine out-of-distribution performance. We focus on the models using aPE and PE(3), as they represent the extreme cases of not including any, and including all Euclidean symmetries in the architecture design, respectively.
Model Scaling. To analyze how DiTMC behaves as model capacity scales, we define a small base (“B”) and a large (“L”) model variant (see Appendix B). The compute cost per gradient step for the equivariant PE(3) model is substantially higher than for the aPE model, despite having a similar number of parameters (Fig. 3A). Even the base PE(3) model requires more compute than the large aPE model. This enables training aPE models for more epochs under the same compute budget (see Appendix Fig. A4). Scaling the capacity of aPE-L yields substantial gains in precision metrics (Fig. 3B and C). For recall, sample fidelity improves with larger model capacity (as indicated by larger COV-R at small thresholds) but differences start to vanish at $\delta=0.75\mbox{\AA}$ (Fig. 3B). Extrapolation performance also improves for the larger model (see Tab. 4). Similar trends are observed for both base and large variants of PE(3) (see Appendix Fig. A4). This allows our aPE-L model to achieve SOTA results on GEOM-DRUGS for all precision metrics (see Tab. 2). For recall, aPE-L achieves better results than ET-Flow, but worse than DMT-L and MCF-L.
Table 3: Median absolute error of ensemble properties between generated and reference conformers for GEOM-DRUGS. $E$, $\Delta\epsilon$, $E_{\text{min}}$ are in kcal/mol, and $\mu$ is in debye. Best results in bold, second best underlined; our models are marked with an asterisk (*). Results for MCF, ET-Flow, and ours are averaged over three random seeds. See Tab. A11 for results including standard deviations.
Method $E$$\downarrow$ $\mu$$\downarrow$ $\Delta\epsilon$$\downarrow$ $E_{\text{min}}$$\downarrow$
OMEGA 0.68 0.66 0.68 0.69
Torsional Diff. 0.22 0.35 0.54 0.13
MCF-L 0.68 0.28 0.63 0.04
ET-Flow 0.18 0.18 0.35 0.02
*DiTMC+aPE-B 0.17 0.16 0.27 0.01
*DiTMC+aPE-L 0.16 0.14 0.27 0.01
Table 4: Out-of-distribution generalization results on GEOM-XL for models trained on GEOM-DRUGS reporting median RMSD in Å. Best results in bold, second best underlined; our models are marked with an asterisk (*). Our results are averaged over three random seeds. See Tab. A13 for results including standard deviations.
Method AMR-R $\downarrow$ AMR-P $\downarrow$
102 molecules
MCF - L (242M) 1.60 2.43
*DiTMC+aPE-B 1.60 2.49
*DiTMC+aPE-L 1.51 2.30
77 molecules
MCF - L (242M) 1.51 2.26
*DiTMC+aPE-B 1.47 2.24
*DiTMC+aPE-L 1.28 2.14
Ensemble properties. To complement RMSD-based geometric evaluation with a chemically meaningful assessment, we report the median absolute error (MAE) of ensemble properties between generated and reference conformers. The RMSD metric can penalize conformers that differ due to rotations or atom reordering, but are otherwise chemically equivalent. For our analysis we follow the protocol of MCF [47] and ET-Flow [32] (see Appendix E.3) and compare energy $E$, dipole moment $\mu$, HOMO-LUMO gap $\Delta\epsilon$, and minimum energy $E_{\text{min}}$. Our aPE-L model predicts ensemble properties more accurately than all baselines, highlighting the physical validity of our generated structures. In particular, MCF-L, which shows better performance for recall metrics, is outperformed by a large margin (up to a factor of $4$ for energy).
Generalization performance. We assess the generalization of our model to larger and unseen molecules using the GEOM-XL dataset [46]. It comprises 102 molecules of size $N>100$ atoms, whereas the molecules in the training data contain only $N=44$ atoms on average. Following MCF [47] we report the generalization performance for all 102 molecules and a subset of 77 molecules. Our models perform on par or better to the previously best-performing method MCF-L while using only a fraction of the parameters (Tab. 4). Other baselines (like ET-Flow) are outperformed by a larger margin (see Appendix Tab. A13).
6 Summary and Limitations
We propose a framework for molecular conformer generation that incorporates conditioning strategies tailored to the architectural design principles of DiTs. This modular framework enables a rigorous exploration of different positional embedding and self-attention strategies, allowing us to identify scalable generative architectures that perform competitively with prior methods on standard benchmarks. Our models achieve SOTA results on GEOM-QM9 and GEOM-DRUGS, excelling in both precision and physical validity. Through ablation studies, we assess the impact of incorporating Euclidean symmetries into DiTs. While such symmetries improve performance, they also increase computational cost. Notably, simpler non-equivariant variants remain highly effective. These findings allow us to develop an efficient, accurate, and scalable DiT architecture suitable for large-scale conformer generation.
Nonetheless, some limitations persist. Our evaluation is currently restricted to small and medium-sized molecules, with larger, more flexible compounds left for future work. Moreover, the training process depends on high-quality ground-truth conformers, which may be unavailable in some cases. Finally, while our analysis advances understanding of equivariance within DiT-based generative models, drawing broader conclusions would require further study.
7 Acknowledgements
JTF, WR, KRM, and SC acknowledge support by the German Ministry of Education and Research (BMBF) for BIFOLD (01IS18037A). Further, this work was in part supported by the BMBF under Grants 01IS14013A-E, 01GQ1115, 01GQ0850, 01IS18025A, 031L0207D, and 01IS18037A. KRM was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grants funded by the Korea government (MSIT) (No.2019-0-00079, Artificial Intelligence Graduate School Program, Korea University and No. 2022-0-00984, Development of Artificial Intelligence Technology for Personalized Plug-and-Play Explanation and Verification of Explanation). We further want to thank Romuald Elie for helpful discussion.
Appendix A Training
Algorithm 1 describes the computation of the training loss for our flow matching objective. We start by sampling from the prior $\boldsymbol{x}_{0}\sim p_{0}(\boldsymbol{x})$, from the data $\boldsymbol{x}_{1}\sim p_{1}(\boldsymbol{x})$, and from a Gaussian distribution $\boldsymbol{\epsilon}\sim N(\boldsymbol{x};0,\boldsymbol{I})$. The interpolant is then constructed as
$\boldsymbol{x}_{t}=(1-t)\cdot\boldsymbol{x}_{0}+t\cdot\boldsymbol{x}_{1}+\sigma\cdot\boldsymbol{\epsilon}\,,$ (A20)
where $\sigma\in\mathbb{R}_{>0}$ is a non-zero noise scaling parameter.
Instead of predicting the conditional vector field $\boldsymbol{u}_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})$ directly, we choose to reparametrize the network such that it predicts the clean sample $\boldsymbol{x}_{1}$. Similar to Ref. [53], we add a weighting term $1/(1-t)^{2}$ to encourage the model to accurately capture fine details close to the data distribution. This gives raise to the following loss function
$\mathcal{L}=\frac{1}{(1-t)^{2}}\|\boldsymbol{x}_{1}^{\theta}(\boldsymbol{x}_{t},t,\mathcal{G})-\boldsymbol{x}_{1}\|^{2}\,,$ (A21)
where $t\in(0,1)$ denotes the point of time in the interpolant $\boldsymbol{x}_{t}\in\mathbb{R}^{N\times 3}$, $\boldsymbol{x}_{1}\in\mathbb{R}^{N\times 3}$ is the clean geometry and $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denotes the molecular graph. The full algorithm is summarized in Algorithm 1.
Geometry Alignment
Given a set of vectors $\mathcal{U}=\{\vec{u}_{1},\dots,\vec{u}_{N}\,|\,\vec{u}_{i}\in\mathbb{R}^{3}\}$ associated with a point cloud $\boldsymbol{x}\in\mathbb{R}^{N\times 3}$ we define a centering operation for the $i$-th row
$\mathrm{Center}(\boldsymbol{x})_{i}=\vec{u}_{i}-\sum_{j=1}^{N}\vec{u}_{j}\,,$ (A22)
which removes global drift in $\boldsymbol{x}$. Given two point clouds $\boldsymbol{x}_{A}\in\mathbb{R}^{N\times 3}$ and $\boldsymbol{x}_{B}\in\mathbb{R}^{N\times 3}$ with positions $\mathcal{R}_{A}=\{\vec{r}_{1A}\,\dots,\vec{r}_{NA}\}$ and $\mathcal{R}_{B}=\{\vec{r}_{1B}\,\dots,\vec{r}_{NB}\}$, we define a rotational alignment operation
$\mathrm{RotationAlign}(\boldsymbol{x}_{A},\boldsymbol{x}_{B})_{i}=\boldsymbol{R}_{\text{opt}}\boldsymbol{x}_{iA},$ (A23)
where $\boldsymbol{R}_{\text{opt}}\in\mathbb{R}^{3\times 3}$ is the optimal rotation matrix, minimizing the root mean square deviation (RMSD) between the positions of point clouds A and B. Here, we employ the Kabsch-Algorithm. For full geometry alignment “$\mathrm{GeometryAlign}(\boldsymbol{x}_{A},\boldsymbol{x}_{B})$”, we do the following
$\boldsymbol{x}_{A}\leftarrow\mathrm{Center}(\boldsymbol{x}_{A})$ (A24)
$\boldsymbol{x}_{B}\leftarrow\mathrm{Center}(\boldsymbol{x}_{B})$ (A25)
$\boldsymbol{x}_{A}\leftarrow\mathrm{RotationAlign}(\boldsymbol{x}_{A},\boldsymbol{x}_{B})$ (A26)
In words, the operations from above first center both point clouds at the origin and then align them as much as possible via a rotation. This procedure also minimizes the path length of a linear interpolation between the point clouds A and B.
Algorithm 1 Conditional Flow Matching Training Loss
1:Graph $\mathcal{G}$, target $\boldsymbol{x}_{1}$, noise level $\sigma$, Model $\boldsymbol{x}_{1}^{\theta}$
2:
3:$\boldsymbol{x}_{0}\sim p_{0}$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, $t\sim\mathcal{U}(0,1)$, $\boldsymbol{R}\sim\mathrm{SO}(3)$
4:$\boldsymbol{x}_{0},\boldsymbol{x}_{1}\leftarrow\mathrm{GeometryAlign}(\boldsymbol{x}_{0},\boldsymbol{x}_{1})$triangleright This centers $\boldsymbol{x}_{0}$ and $\boldsymbol{x}_{1}$ and rotation-aligns $\boldsymbol{x}_{0}$ to $\boldsymbol{x}_{1}$.
5:$\boldsymbol{\epsilon}\leftarrow\mathrm{Center}(\boldsymbol{\epsilon})$
6:$\boldsymbol{x}_{t}\leftarrow(1-t)\boldsymbol{x}_{0}+t\boldsymbol{x}_{1}+\sigma\boldsymbol{\epsilon}$
7:$\boldsymbol{x}_{t}\leftarrow\mathrm{ApplyRotation}(\boldsymbol{R},\boldsymbol{x}_{t})$
8:$\boldsymbol{x}_{1}\leftarrow\mathrm{ApplyRotation}(\boldsymbol{R},\boldsymbol{x}_{1})$
9:$\hat{\boldsymbol{x}}_{1}\leftarrow\boldsymbol{x}_{1}^{\theta}(\boldsymbol{x}_{t},t,\mathcal{G})$
10:$\hat{\boldsymbol{x}}_{1}\leftarrow\mathrm{Center}(\hat{\boldsymbol{x}}_{1})$
11:
12:return $\frac{1}{(1-t)^{2}}\|\hat{\boldsymbol{x}}_{1}-\boldsymbol{x}_{1}\|^{2}$
Data Augmentation
One can construct an $\mathrm{SE}(3)$-invariant density by learning an $\mathrm{SO}(3)$-equivariant vector field on centered $\mathrm{SE}(3)$ (see section A.1). However, only DiTMC with PE(3) is $\mathrm{SO}(3)$-equivariant, whereas aPE and rPE violate $\mathrm{SO}(3)$-equivariance. Therefore, we learn equivariance approximately during training, using data augmentation. Specifically, we randomly sample rotation matrices $\boldsymbol{R}$ (orthogonal matrices with determinant +1) and apply them as
$\mathrm{ApplyRotation}(\boldsymbol{R},\boldsymbol{x})_{i}=\boldsymbol{R}\vec{r}_{i},$ (A27)
where $\vec{r}_{i}$ denotes the positions of the $i$-th atom, i.e., the $i$-th row in the point cloud $\boldsymbol{x}\in\mathbb{R}^{N\times 3}$.
Noise Scaling Parameter
We ablated the noise scaling parameter $\sigma$ for both GEOM QM9 and drugs experiments, comparing a larger value of $0.5$ and a smaller value of $0.05$. We set the $\sigma$ parameter to the value that empirically worked best for each dataset: $0.05$ for Geom QM9 and $0.5$ for GEOM drugs.
Optimizer and Hyperparameters
We use the AdamW optimizer (weight decay 0.01) with initial learning rate of $\mu_{0}=10^{-5}$. It is increased up to $\mu_{\text{max}}$ via a linear learning rate warmup up over the first 1% of training steps. Afterwards, it is decreased to $\mu_{\text{min}}$ via a cosine decay schedule. We use $\mu_{\text{max}}=3\times 10^{-4}$ and $\mu_{\text{min}}=0$ for GEOM-QM9 and $\mu_{\text{max}}=1\times 10^{-4}$ and $\mu_{\text{min}}=1\times 10^{-5}$ for GEOM-DRUGS. We use a batch size of 128 for all data sets and models. All models on QM9 are trained for 250 epochs in total. For GEOM-DRUGS we determine the maximal number of epochs from the epochs the larger model “L” can perform within a fixed compute budget of 9 days. Thus, we train the PE(3) variants for 10 epochs and the aPE variants for 50 epochs. Fig. A4A shows the number of gradient steps per day that can be performed with the different DiTMC variants.
A.1 From SE(3) to SO(3) Invariance
The target data distribution of molecular conformers $p_{1}(\boldsymbol{x})$ is $\mathrm{SE}(3)$-invariant, i.e. it does not change under translations and rotations of the input. Following Ref. [67], one can define an $\mathrm{SE}(3)$-invariant measure on $\mathrm{SE}(3)^{N}$ by keeping the center of mass fixed at zero, which can be achieved via the centering operation from Eq. A22. This defines a subgroup $\mathrm{SE}(3)^{N}_{0}$, called centered $\mathrm{SE}(3)$. It can then be shown, that one can define an $\mathrm{SE}(3)$-invariant measure on $\mathrm{SE}(3)^{N}_{0}$ by constructing an $\mathrm{SO}(3)$-invariant (rotationally invariant) measure on $\mathrm{SE}(3)^{N}_{0}$.
As a consequence, it is then sufficient to learn an $\mathrm{SO}(3)$-equivariant vector field on the space of centered input positions (see also Ref. [68]). This is achieved by centering $\boldsymbol{x}_{0}$, $\boldsymbol{x}_{1}$ and $\boldsymbol{z}$ for the calculation of the interpolant. Moreover, the neural network output (predicted velocities) and the clean target $\boldsymbol{x}_{1}$ must be centered to have zero center of mass. We discuss the implications for sampling in section C.
A.2 Compute and Training Times
Tab. A5 compares training and inference time for our models (see also Fig. 3A for a visual comparison). Despite having a similar parameter count, the equivariant PE(3)-B model is approximately $3.5$ times slower than models using aPE or rPE, while at inference time, the factor is $\sim 3$. The effect is even more pronounced for our large models, where the equivariant model is almost 5 times slower during training compared to its non-equivariant counterpart with a similar number of parameters. Even our large non-equivariant model is significantly faster to train and sample from than the equivariant base model with far fewer parameters.
All our models on GEOM QM9 are trained for 250 epochs, which requires 2 days of training on Nvidia H100 GPU for aPE and rPE models and almost 4 days for PE(3). Due to compute constraints, our non-equivariant (DiTMC+aPE) models are trained for 50 epochs on GEOM drugs. The equivariant models are trained for 10 epochs on GEOM drugs resulting in a similar compute budget for training in terms of GPU hours. Our maximal compute budget per model is 9 days on a single Nvidia H100 GPU.
Table A5: Architectural details as well as training and inference time (in ms) for different PE strategies on QM9. Times are measured with batch size 128 on a single Nvidia H100 GPU. $T$ means number of transformer layers, $n_{\text{heads}}$ number of heads, $d_{\text{head}}$ number of features per head in the attention update, and $T_{MGN}$ number of layers for the conditioning mesh graph net. Thus, total feature dimension is given as $H=n_{\text{heads}}\cdot d_{\text{head}}$.
Model $T$ $n_{\text{heads}}$ $d_{\text{head}}$ $T_{MGN}$ $H$ Train [ms] Infer [ms]
DiTMC+aPE-B (9.5M) 6 8 32 2 256 19.2 8.1
DiTMC+rPE-B (9.6M) 6 8 32 2 256 19.7 8.3
DiTMC+PE(3)-B (8.6M) 6 6 32 2 192 70.0 25.9
DiTMC+aPE-L (28.2M) 8 12 32 3 384 32.7 9.5
DiTMC+PE(3)-L (31.1M) 8 10 32 3 320 151.6 41.8
Appendix B Architecture Details
Tab. A5 shows the architectural details for our base and large model variants. Tab. A6 contains details about the MLPs we use throughout our architecture.
Table A6: Architecture details for MLPs used in the model. The feature dimension is given as $H=n_{\text{heads}}\cdot d_{\text{head}}$ where $n_{\text{head}}$ is the number of heads and $d_{\text{head}}$ is the number of features per head. The use of gated GELU ensures equivariance (see section Sec. B.4).
Name Layers Hidden Dim Out Dim Activation Input
DiT Block 2 $4H$ $H$ GELU Tokens
SO(3) DiT Block 2 $4H$ $H$ gated GELU Tokens
Time and Atom Cond. 1 $6H$ SiLU $\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}}$
Bond pair 2 $H$ $H$ SiLU $\boldsymbol{c}_{ij}^{\mathcal{G}}$
Time embedding 2 $H$ $H$ SiLU $t\in[0,1]$
Shortest-hop embedding 2 $H$ $H$ SiLU Hop distance
aPE embedding 2 $H$ $H$ SiLU Abs. ositions $\vec{r}_{i}$
rPE embedding 2 $H$ $H$ SiLU Rel. positions $\vec{r}_{ij}$
GNN embedding 2 $H$ $H$ SiLU Node/edge features
We first describe the building blocks of our non-equivariant DiTMC implementation and then its equivariant counterpart. Both architectures rely on conditioning tokens $\mathcal{C}$, for the time $\boldsymbol{c}^{t}\in\mathbb{R}^{H}$, per-node $\boldsymbol{c}_{i}^{\mathcal{G}}\in\mathbb{R}^{H}$, per bond-pair $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$ and for all-pairs $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$. See main text section 4.1 for more details. Following Ref. [13], we use adaptive layer norm (AdaLN) and adaptive scale (AdaScale) to include per-node conditioning based on time $t$ and molecular graph information. To that end, we construct conditioning tokens
$\boldsymbol{c}_{i}=\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}}.$ (A28)
B.1 Non Equivariant DiT
In the non-equivariant DiTMC formulations with aPE and rPE, we have the following set of tokens =$\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$.
B.1.1 Adaptive Layer Normalization and Adaptive Scale
In the standard, non-equivariant setting, we can follow the standard approach of other DiT architectures. We define adaptive layer norm as
$\text{AdaLN}(\boldsymbol{h},\boldsymbol{\alpha},\boldsymbol{\beta})=\text{LN}(\boldsymbol{h})\odot(1+\boldsymbol{\alpha})+\boldsymbol{\beta},$ (A29)
where LN is a standard layer normalization without trainable scale and bias, and $\odot$” denotes entry-wise product.
Adaptive Scale is defined as
$\text{AdaScale}(\boldsymbol{h},\boldsymbol{\gamma})=\boldsymbol{h}\odot\boldsymbol{\gamma}.$ (A30)
In each DiTMC block, we calculate
$\boldsymbol{\alpha}_{1i},\boldsymbol{\beta}_{1i},\boldsymbol{\gamma}_{1i},\boldsymbol{\alpha}_{2i},\boldsymbol{\beta}_{2i},\boldsymbol{\gamma}_{2i}=\boldsymbol{W}\big{(}\text{SiLU}(\boldsymbol{c}_{i})\big{)},$ (A31)
where $\boldsymbol{W}\in\mathbb{R}^{6H\times H}$ and the output is split into six equally sized vectors $\boldsymbol{\alpha}_{1i},\boldsymbol{\beta}_{1i},\boldsymbol{\gamma}_{1i},\boldsymbol{\alpha}_{2i},\boldsymbol{\beta}_{2i},\boldsymbol{\gamma}_{2i}\in\mathbb{R}^{H}$. The weight matrix $\boldsymbol{W}$ is initialized to all zeros, such that “AdaLN” behaves like identity at initialization. “AdaScale” damps all input tokens to zero at initialization such that the whole DiT block behaves like the identity function at initialization.
B.1.2 Readout
Given final tokens $\boldsymbol{h}$ after performing updates via $T$ DiTMC blocks, we use a readout layer to predict the atomic positions of the clean data sample $\boldsymbol{x}_{1}$. As in the DiTMC blocks, we employ adaptive LN and therefore calculate
$\boldsymbol{\alpha}_{i},\boldsymbol{\beta}_{i}=\boldsymbol{W}\big{(}\text{SiLU}(\boldsymbol{c}_{i})\big{)},$ (A32)
with weight matrix $\boldsymbol{W}\in\mathbb{R}^{2H\times H}$ initialized to all zeros and $\boldsymbol{\alpha}_{i},\boldsymbol{\beta}_{i}\in\mathbb{R}^{H}$ and do
$\boldsymbol{h}_{i}$ $\leftarrow\text{AdaLN}(\boldsymbol{h}_{i},\boldsymbol{\alpha}_{i},\boldsymbol{\beta}_{i})\,,$
$\hat{\boldsymbol{x}}_{i}$ $\leftarrow\boldsymbol{W}_{\text{readout}}\boldsymbol{h}_{i}\,,$
where $\boldsymbol{W}_{\text{readout}}\in\mathbb{R}^{3\times H}$ is a trainable weight matrix. Thus, we predict a three-dimensional vector per-atom.
B.2 SO(3) Equivariant DiT
Following the notation in Ref. [60], we denote $\mathrm{SO}(3)$-equivariant tokens as =$\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\mathbb{R}^{(L+1)^{2}\times H}\}$, where $L$ denotes the maximal degree of the spherical harmonics. We denote the features corresponding to the $\ell$-th degree as $\boldsymbol{h}_{i}^{(\ell)}\in\mathbb{R}^{(2\ell+1)\times H}$, where the $(2\ell+1)$ entries corresponds to the orders $m=-\ell,\dots,+\ell$ per degree $\ell$. We refer the reader to Ref. [60] for an in-depth introduction into equivariant features. For all our experiments we use maximal degree of $L=1$.
B.2.1 Adaptive Layer Normalization and Adaptive Scale
In the $\mathrm{SO}(3)$-equivariant case, we define adapted version of AdaLN and AdaScale, preserving equivariance. Our equivariant formulation of AdaLN is given as
$\text{EquivAdaLN}(\boldsymbol{h},\boldsymbol{\alpha},\boldsymbol{\beta})=\begin{cases}\text{LN}(\boldsymbol{h}^{(\ell)})\odot(1+\boldsymbol{\alpha}^{(0)})+\boldsymbol{\beta}\hskip 100.00015pt&\text{for }\ell=0,\\ \text{EquivLN}(\boldsymbol{h}^{(\ell)})\odot(1+\boldsymbol{\alpha}^{(\ell)})&\text{for }\ell>0,\end{cases}$ (A33)
where EquivLN is the equivariant formulation of layer normalization following Ref. [36] without trainable per-degree scales and LN is standard layer normalization without trainable scale and bias. Scaling vectors $\boldsymbol{\alpha}^{(\ell)}\in\mathbb{R}^{1\times H}$ are defined per degree $\ell$, such that input scaling vectors are tensors $\boldsymbol{\alpha}\in\mathbb{R}^{(L+1)\times H}$. Bias vectors are only defined for the invariant ($\ell=0$) component of the tokens, since adding a non-zero bias to components with $\ell>0$ would lead to a non-equivariant operation (the bias does not transform under rotations). The element wise multiplication between $(1+\boldsymbol{\alpha}^{(\ell)})\in\mathbb{R}^{1\times H}$ and tokens $h^{(\ell)}\in\mathbb{R}^{(2\ell+1)\times H}$ is “broadcasted” along the degree-axis. For $L=0$, Eq. A33 reduces to the standard adaptive layer normalization.
Equivariant adaptive scale is defined as
$\text{EquivAdaScale}(\boldsymbol{h},\boldsymbol{\gamma})=\boldsymbol{h}^{(\ell)}\odot\boldsymbol{\gamma}^{(\ell)}.$ (A34)
As for “EquivAdaLN”, we define a separate $\boldsymbol{\gamma}^{(\ell)}\in\mathbb{R}^{1\times H}$ per degree $\ell$, such that $\boldsymbol{\gamma}\in\mathbb{R}^{(L+1)\times H}$. Again, the element wise product is “broadcasted” along the degree-axis. Since no bias is involved, the invariant and equivariant parts in $\boldsymbol{h}$ can be treated equally.
Within each $\mathrm{SO}(3)$-equivariant DiTMC block, we calculate
$\boldsymbol{\alpha}_{1i},\boldsymbol{\beta}_{1i},\boldsymbol{\gamma}_{1i},\boldsymbol{\alpha}_{2i},\boldsymbol{\beta}_{2i},\boldsymbol{\gamma}_{2i}=\boldsymbol{W}\big{(}\text{SiLU}(\boldsymbol{c}_{i})\big{)},$ (A35)
where $\boldsymbol{\alpha}_{1i},\boldsymbol{\alpha}_{2i},\boldsymbol{\gamma}_{1i},\boldsymbol{\gamma}_{2i}\in\mathbb{R}^{(L+1)\times H}$ and $\boldsymbol{\beta}_{1i},\boldsymbol{\beta}_{2i}\in\mathbb{R}^{H}$. Thus, the weight matrix is given as $\boldsymbol{W}\in\mathbb{R}^{(4(L+1)+2)\times H}$ and initialized to all zeros, such that “EquivAdaLN” behaves like identity at initialization and “EquivAdaScale” returns zeros. Thus, also the $\mathrm{SO}(3)$-equivariant DiTMC block behaves like the identity function at initialization.
B.2.2 Equivariant positional embeddings
Given a set of transformations that act on a vector space $\mathbb{A}$ as $S_{g}:\mathbb{A}\mapsto\mathbb{A}$ to which we associate an abstract group $G$, a function $f:\mathbb{A}\mapsto\mathbb{B}$ is said to be equivariant w.r.t. $G$ if
$f(S_{g}x)=T_{g}f(x)\,,$ (A36)
where $T_{g}:\mathbb{B}\mapsto\mathbb{B}$ is an equivalent transformation on the output space. Thus, in order to say that $f$ is equivariant, it must hold that under transformation of the input, the output transforms “in the same way”.
Let us now recall our definition for the equivariant positional embeddings for a single degree $\ell$
$\boldsymbol{p}_{ij}^{\text{PE(3)},(\ell)}(\vec{r}_{ij})=\boldsymbol{\phi}_{\ell}(||\vec{r}_{ij}||)\,\odot\boldsymbol{Y}_{\ell}(\hat{r}_{ij})\,,$ (A37)
where $\boldsymbol{\phi}_{\ell}:\mathbb{R}\mapsto\mathbb{R}^{1\times H}$ is a radial filter function, $\hat{r}=\vec{r}/r$, and $\boldsymbol{Y}_{\ell}\in\mathbb{R}^{(2\ell+1)\times 1}$ are spherical harmonics of degree $\ell=0\dots L$. The element-wise multiplication $\odot$’ between radial filters and spherical harmonics is understood to be “broadcasting” along axes with size $1$, such that $(\boldsymbol{\phi}_{\ell}\,\odot\boldsymbol{Y}_{\ell})\in\mathbb{R}^{(2\ell+1)\times H}$. We have also made the dependence of PE(3) on the pairwise displacement vector $\vec{r}_{ij}=\vec{r}_{i}-\vec{r}_{j}$ explicit.
Lets not consider a single feature channel $c$ in PE(3), which is given as
$\boldsymbol{p}_{ijc}^{\text{PE(3)},(\ell)}(\vec{r}_{ij})=\boldsymbol{\phi}_{\ell c}(||\vec{r}_{ij}||)\,\odot\boldsymbol{Y}_{\ell}(\hat{r}_{ij})\,.$ (A38)
Rotating the input positions in Eq. A38 leads to
$\boldsymbol{p}_{ijc}^{\text{PE(3)},(\ell)}(\boldsymbol{R}\vec{r}_{ij})$ $=\boldsymbol{\phi}_{\ell c}(||\boldsymbol{R}\vec{r}_{ij}||)\,\odot\boldsymbol{Y}_{\ell}(\boldsymbol{R}\hat{r}_{ij})$ (A39)
$=\boldsymbol{\phi}_{\ell c}(||\vec{r}_{ij}||)\,\odot\boldsymbol{D}^{(\ell)}(\boldsymbol{R})\boldsymbol{Y}_{\ell}(\hat{r}_{ij}),$ (A40)
$=\boldsymbol{D}^{(\ell)}(\boldsymbol{R})\boldsymbol{p}_{ijc}^{\text{PE(3)},(\ell)}(\vec{r}_{ij})$ (A41)
where $\boldsymbol{D}^{(\ell)}\in\mathbb{R}^{(2\ell+1)\times(2\ell+1)}$ are the Wigner-D matrices for degree $\ell$ and $\boldsymbol{R}\in\mathbb{R}^{3\times 3}$ is a rotation matrix. According to Eq. A36 and Eq. A41, each channel transforms equivariant and thus $\boldsymbol{p}_{ij}^{\text{PE(3)},(\ell)}(\vec{r}_{ij})$ is also equivariant.
The concatenation of different degrees $\ell$ up to some maximal degree $L$ as given in the main body of the text
$\boldsymbol{p}_{ij}^{\text{PE(3)}}(\vec{r}_{ij})=\bigoplus_{\ell=0}^{L}\boldsymbol{\phi}_{\ell}(r_{ij})\,\odot\boldsymbol{Y}_{\ell}(\hat{r}_{ij})\,,$ (A42)
transforms under rotation as
$\boldsymbol{p}_{ij}^{\text{PE(3)}}(\boldsymbol{R}\vec{r}_{ij})=\boldsymbol{D}(\boldsymbol{R})\,\boldsymbol{p}_{ij}^{\text{PE(3)}}(\vec{r}_{ij})$ (A43)
with $\boldsymbol{D}(\boldsymbol{R})=\bigoplus_{\ell=0}^{L}\boldsymbol{D}^{(\ell)}(\boldsymbol{R})\in\mathbb{R}^{(L+1)^{2}\times(L+1)^{2}}$ being a block-diagonal matrix with the Wigner-D matrices of degree $\boldsymbol{D}^{(\ell)}(\boldsymbol{R})\in\mathbb{R}^{(2\ell+1)\times(2\ell+1)}$ along the diagonal. Therefore, according to Eq. A36 the proposed positional embeddings $\boldsymbol{p}_{ij}^{\text{PE(3)}}$ are $\mathrm{SO}(3)$-equivariant.
B.3 Invariance of the Dot-Product
In the self-attention update, the dot-product between query and key is computed as along the degree and the feature axis. Under rotation, the equivariant features behave as
$\boldsymbol{h}(\boldsymbol{R}\vec{r})=\boldsymbol{D}(\boldsymbol{R})\boldsymbol{h}(\vec{r}),$ (A44)
where $\boldsymbol{D}(\boldsymbol{R})$ is the concatenation of Wigner-D matrices from above. The inner product along the degree axis for two features $\boldsymbol{h}$ and $\boldsymbol{g}$ behaves under rotation as
$\boldsymbol{g}(\boldsymbol{R}\vec{r})^{T}\cdot\boldsymbol{h}(\boldsymbol{R}\vec{r})=\boldsymbol{g}(\vec{r})^{T}\underbrace{\boldsymbol{D}(\boldsymbol{R})^{T}\boldsymbol{D}(\boldsymbol{R})}_{=\text{Id}}\boldsymbol{h}(\vec{r})=\boldsymbol{g}(\vec{r})^{T}\cdot\boldsymbol{h}(\vec{r}),$ (A45)
where we made use of the fact that the Wigner-D matrices are orthogonal matrices. Thus, the dot-product along the degree-axis is invariant and therefore taking the dot-product along the degree and then along the feature axis is also invariant.
B.3.1 Readout
Given final equivariant features $\boldsymbol{h}_{i}\in\mathbb{R}^{(L+1)^{2}\times H}$ we use a readout layer to predict the atomic positions of the clean data sample $\boldsymbol{x}_{1}$. We employ our equivariant formulation of adaptive layer normalization and calculate
$\boldsymbol{\alpha}_{i},\boldsymbol{\beta}_{i}=\boldsymbol{W}\big{(}\text{SiLU}(\boldsymbol{c}_{i})\big{)},$ (A46)
with weight matrix $\boldsymbol{W}\in\mathbb{R}^{2H(L+1)\times H}$ initialized to all zeros and $\boldsymbol{\alpha}_{i},\boldsymbol{\beta}_{i}\in\mathbb{R}^{H(L+1)}$. We then do,
$\boldsymbol{h}_{i}$ $\leftarrow\text{EquivAdaLN}(\boldsymbol{h}_{i},\boldsymbol{\alpha}_{i},\boldsymbol{\beta}_{i})\,,$
$\boldsymbol{y}_{i}$ $\leftarrow\boldsymbol{W}_{\text{readout}}\boldsymbol{h}_{i}^{(\ell=1)}\,,$
where $\boldsymbol{W}_{\text{readout}}\in\mathbb{R}^{1\times H}$ is a trainable weight vector that is applied along the feature axis in $\boldsymbol{h}$. Since $\boldsymbol{h}_{i}^{(\ell=1)}\in\mathbb{R}^{3\times H}$ this produces per-atom vectors $\hat{\boldsymbol{y}}_{i}\in\mathbb{R}^{3}$. As $\boldsymbol{h}_{i}$ are rotationally equivariant so is $\hat{\boldsymbol{y}}_{i}\in\mathbb{R}^{3}$.
B.4 Equivariant MLP
Standard MLPs do not preserve the equivariance of the tokens. However, it is possible to define an equivariant formulation for dense layers and so-called gated non-linearities which preserve equivariance. We use them to build equivariant MLPs for the node-wise refinement after the self-attention calculation. See e.g. Ref. [60] for more details.
Appendix C Sampling
For sampling, we use a simple Euler scheme with 50 steps to sample from the associated ordinary differential equation (ODE) as described in Algorithm 2. Since during training we predict the clean sample $\boldsymbol{x}_{1}$, we re-parametrize the velocity required for the integration as
$\boldsymbol{v}^{\theta}_{t}(\boldsymbol{x}_{t},t,\mathcal{G})=\frac{\boldsymbol{x}^{\theta}_{1}(\boldsymbol{x}_{t},t,\mathcal{G})-\boldsymbol{x}_{t}}{1-t},$ (A47)
where $\boldsymbol{x}^{\theta}_{1}(\boldsymbol{x}_{t},t,\mathcal{G})$ is the original output of DiTMC.
To ensure $\mathrm{SE}(3)$ invariance of the probability path from an (approximately) $\mathrm{SO}(3)$-equivariant velocity predictor, we center the prior $\boldsymbol{x}_{0}\sim p_{0}(\boldsymbol{x})$ as well as the prediction of DiTMC in each ODE step.
Algorithm 2 ODE Sampling
1:Model $\boldsymbol{x}_{1}^{\theta}$, Graph $\mathcal{G}$, steps $N>0$
2:
3:$t_{n}\leftarrow n/N$ for $n\in\{0,\ldots,N\}$
4:$\boldsymbol{x}_{0}\sim p_{\text{prior}}(\boldsymbol{x})$triangleright Sample prior.
5:$\boldsymbol{x}_{0}\leftarrow\mathrm{Center}(\boldsymbol{x}_{0})$
6:for $n\leftarrow 0$ to $N-1$ do
7:     $\Delta t\leftarrow t_{n+1}-t_{n}$triangleright Compute step size.
8:     $\hat{\boldsymbol{x}}_{1}\leftarrow\boldsymbol{x}_{1}^{\theta}(\boldsymbol{x}_{t_{n}},t_{n},\mathcal{G})$
9:     $\hat{\boldsymbol{x}}_{1}\leftarrow\mathrm{Center}(\hat{\boldsymbol{x}}_{1})$
10:     $\boldsymbol{v}\leftarrow(\hat{\boldsymbol{x}}_{1}-\boldsymbol{x}_{t_{n}})/(1-t_{n})$
11:     $\boldsymbol{x}_{t_{n+1}}\leftarrow\boldsymbol{x}_{t_{n}}+\Delta t\cdot\boldsymbol{v}$triangleright Euler step.
12:end for
13:
14:return $\boldsymbol{x}_{1}$
Appendix D Implementation details
D.1 Data Preprocessing
For both GEOM-QM9 and GEOM-DRUGS, we use the first 30 conformers for each molecule with the lowest energies, i.e., highest Boltzmann weights. We use the train/test/val split from Geomol [65], using the same 1000 molecules for testing.
D.2 Input Featurization
Tab. A7 defines the features we use for each atom or bond. Each feature is one-hot encoded before being passed to the network.
Table A7: Atomic and bond features included in DiT-MC. All features are one-hot encoded.
Atom features Options
Chirality TETRAHEDRAL_CW, TETRAHEDRAL_CCW, UNSPECIFIED, OTHER
Number of hydrogens 0, 1, 2, 3, 4
Number of radical electrons 0, 1, 2, 3, 4
Atom type (QM9) H, C, N, O, F
Atom type (DRUGS) H, Li, B, C, N, O, F, Na, Mg, Al, Si, P, S, Cl, K, Ca, V, Cr, Mn, Cu,
Zn, Ga, Ge, As, Se, Br, Ag, In, Sb, I, Gd, Pt, Au, Hg, Bi
Aromaticity true, false
Degree 0, 1, 2, 3, 4, other
Hybridization $sp,sp^{2},sp^{3},sp^{3}d,sp^{3}d^{2}$, other
Implicit valence 0, 1, 2, 3, 4, other
Formal charge -5, -4, …, 5, other
Presence in ring of size x x = 3, 4, 5, 6, 7, 8, other
Number of rings atom is in 0, 1, 2, 3, other
Bond features Options
Bond type single, double, triple, aromatic
D.3 Evaluation Metrics
During evaluation, we follow the same procedure as described in Refs. [65, 47, 46, 32]. The root-mean-square deviation (RMSD) metric measures the average distance between atoms of a generated conformer with respect to its reference, while taking into account all possible symmetries. For $L=2K$ let $\{\hat{C}_{l}\}_{l\in\{1,\dots,L\}}$ and $\{C_{k}\}_{k\in\{1,\dots,K\}}$ be the sets of generated conformers and reference conformers respectively. The Average Minimum RMSD (AMR) and Coverage (COV) metrics for both Recall (R) and Precision (P) are defined as follows, where $\delta>0$ is the coverage threshold:
[…]
D.4 Chirality Correction
Given the four 3D coordinates around a chirality center denoted as $\mathbf{p_{1}},\mathbf{p_{2}},\mathbf{p_{3}},\mathbf{p_{4}}\in\mathbb{R}^{3}$ with $\mathbf{p_{i}}=(x_{i},y_{i},z_{i})$ for $i=1,2,3,4$, we can compute the volume $V$ of the tetrahedron as follows
$V(\mathbf{p_{1}},\mathbf{p_{2}},\mathbf{p_{3}},\mathbf{p_{4}})=\det\left(\begin{bmatrix}1&1&1&1\\ x_{1}&x_{2}&x_{3}&x_{4}\\ y_{1}&y_{2}&y_{3}&y_{4}\\ z_{1}&z_{2}&z_{3}&z_{4}\end{bmatrix}\right)=(\mathbf{p_{1}}-\mathbf{p_{4}})\cdot((\mathbf{p_{2}}-\mathbf{p_{4}})\times(\mathbf{p_{3}}-\mathbf{p_{4}}))$ (A52)
Following GeomMol [65] and ET-Flow [32] we can then compare the orientation of the volume given by $\text{sign}(V)$ with the local chirality label produced by RDKit, which corresponds to a certain orientation as well (CW = +1 and CCW = -1) [69]. If the orientation of the volume differs from the RDKit label, we correct the chirality of the conformer by reflecting its positions against the $z$-axis.
Appendix E Additional Experimental Results
E.1 Scaling Results for PE(3)
Figure A4: (A) Number of gradient steps vs. wall time for PE(3) and aPE on the GEOM-DRUGS dataset measured on Nvidia H100 GPU with batch size 128. (B) Coverage precision and recall as function of threshold $\delta$ and (C) mean AMR metrics for base (“B”) and large (“L”) model of DiTMC+PE(3).
We investigate the scaling behavior of our equivariant models in Fig. A4. Tab. A10 shows a comparison of our base and large models for GEOM drugs. Similar to its non-equivariant counterpart, our scaled PE(3)-L model yields improved results for precision metrics, while we observe only minor improvements for recall. Tab. A11, Tab. A9 and Tab. A13 further demonstrate improvements for ensemble property prediction and out-of-distribution generalization using the larger model.
E.2 Conditioning ablation
To evaluate the effectiveness of various graph conditioning strategies in DiTMC , we compare the performance of three conditioning methods – “node only”, “node & bond-pair”, and “node & all-pair” – against a baseline model without any conditioning. As reported in Tab. A8, all conditioning variants significantly reduce the Average Minimum RMSD (AMR) for both recall (AMR-R) and precision (AMR-P) across the aPE and PE(3) settings, compared to the unconditioned baseline.
Notably, the “node & all-pair” strategy achieves the best overall performance, with the lowest AMR values. These results highlight the strength of the all-pair conditioning strategy, which leverages graph geodesics to incorporate information from all atom pairs, rather than restricting conditioning to directly connected nodes or bonded pairs. This comprehensive approach captures more global structural information, thereby improving both precision and recall. See Appendix K for a more in-depth analysis.
Table A8: Ablation of conditioning strategies in DiTMC (in brackets) with absolute PE and PE(3) on GEOM-QM9. -R indicates Recall, -P indicates Precision. Best results in bold. Our results are averaged over three random seeds with standard deviation reported below.
Method COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
Mean Median Mean Median Mean Median Mean Median
DiTMC+aPE-B (no conditioning) 68.6 $\pm$1.0 91.7 $\pm$2.1 0.405 $\pm$0.005 0.325 $\pm$0.004 36.8 $\pm$0.5 36.4 $\pm$2.2 0.729 $\pm$0.006 0.703 $\pm$0.007
DiTMC+PE(3)-B (no conditioning) 71.5 $\pm$0.5 99.0 $\pm$1.4 0.358 $\pm$0.003 0.266 $\pm$0.005 40.4 $\pm$0.2 42.2 $\pm$0.4 0.683 $\pm$0.003 0.653 $\pm$0.005
DiTMC+aPE-B (node only) 96.3 $\pm$0.0 100.0 $\pm$0.0 0.079 $\pm$0.000 0.037 $\pm$0.000 93.2 $\pm$0.2 100.0 $\pm$0.0 0.112 $\pm$0.001 0.051 $\pm$0.000
DiTMC+PE(3)-B (node only) 95.5 $\pm$0.4 100.0 $\pm$0.0 0.074 $\pm$0.003 0.026 $\pm$0.001 91.1 $\pm$0.4 100.0 $\pm$0.0 0.114 $\pm$0.003 0.041 $\pm$0.001
DiTMC+aPE-B (node & bond-pair) 96.5 $\pm$0.1 100.0 $\pm$0.0 0.077 $\pm$0.001 0.035 $\pm$0.001 95.3 $\pm$0.2 100.0 $\pm$0.0 0.092 $\pm$0.001 0.046 $\pm$0.002
DiTMC+PE(3)-B (node & bond-pair) 96.1 $\pm$0.3 100.0 $\pm$0.0 0.068 $\pm$0.001 0.022 $\pm$0.001 93.6 $\pm$0.2 100.0 $\pm$0.0 0.091 $\pm$0.003 0.035 $\pm$0.001
DiTMC+aPE-B (node & all-pair) 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
DiTMC+PE(3)-B (node & all-pair) 95.7 $\pm$0.3 100.0 $\pm$0.0 0.068 $\pm$0.002 0.021 $\pm$0.001 93.4 $\pm$0.2 100.0 $\pm$0.0 0.089 $\pm$0.002 0.032 $\pm$0.001
E.3 Ensemble Properties
While RMSD-based metrics offer a geometric perspective on the quality of the generated conformers, they do not assess the chemical fidelity. To address this, we evaluate the median averaged errors of different ensemble properties between the generated and ground truth conformers. We adopt the property prediction task setup from MCF [47] and ET-Flow [32], where we draw a subset of 100 randomly sampled molecules from the test set of GEOM-DRUGS and generate $\min(2K,32)$ conformers for a molecule with $K$ ground truth conformers. Afterwards we relax the conformers using GFN2-xTB [70] and compare the Boltzmann-weighted properties of the generated and ground truth ensembles. More specifically, we employ xTB [70] to calculate the energy $E$, the dipole moment $\mu$, the HOMO-LUMO gap $\Delta\epsilon$ and the minimum energy $E_{\text{min}}$. We repeat this procedure for three subsets each sampled with a different random seed and report the averaged median absolute error and standard deviation of the different ensemble properties in Table A11.
Table A9: Results on GEOM-QM9 for different generative models (number of parameters in parentheses). -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined. Our models are marked with an asterisk “*”. Our results are averaged over three random seeds with standard deviation reported below. Other works do not report standard deviations.
Method COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
Mean Median Mean Median Mean Median Mean Median
GeoMol (0.3M) 91.5 100.0 0.225 0.193 86.7 100.0 0.270 0.241
GeoDiff (1.6M) 76.5 100.0 0.297 0.229 50.0 33.5 0.524 0.510
Tors. Diff. (1.6M) 92.8 100.0 0.178 0.147 92.7 100.0 0.221 0.195
MCF-B (64M) 95.0 100.0 0.103 0.044 93.7 100.0 0.119 0.055
DMT-B (55M) 95.2 100.0 0.090 0.036 93.8 100.0 0.108 0.049
ET-Flow (8.3M) 96.5 100.0 0.073 0.030 94.1 100.0 0.098 0.039
*DiTMC+aPE-B (9.5M) 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
*DiTMC+rPE-B (9.6M) 96.3 $\pm$0.0 100.0 $\pm$0.0 0.070 $\pm$0.001 0.027 $\pm$0.000 95.7 $\pm$0.1 100.0 $\pm$0.0 0.080 $\pm$0.000 0.035 $\pm$0.000
*DiTMC+PE(3)-B (8.6M) 95.7 $\pm$0.3 100.0 $\pm$0.0 0.068 $\pm$0.002 0.021 $\pm$0.001 93.4 $\pm$0.2 100.0 $\pm$0.0 0.089 $\pm$0.002 0.032 $\pm$0.001
Table A10: Results on GEOM-DRUGS for different generative models (number of parameters in parentheses). -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined. Our models are marked with an asterisk “*”. Our results are averaged over three random seeds with standard deviation reported below. Other works do not report standard deviations.
Method COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
Mean Median Mean Median Mean Median Mean Median
GeoMol (0.3M) 44.6 41.4 0.875 0.834 43.0 36.4 0.928 0.841
GeoDiff (1.6M) 42.1 37.8 0.835 0.809 24.9 14.5 1.136 1.090
Tors. Diff. (1.6M) 72.7 80.0 0.582 0.565 55.2 56.9 0.778 0.729
MCF-L (242M) 84.7 92.2 0.390 0.247 66.8 71.3 0.618 0.530
DMT-L (150M) 85.8 92.3 0.375 0.346 67.9 72.5 0.598 0.527
ET-Flow - SS (8.3M) 79.6 84.6 0.439 0.406 75.2 81.7 0.517 0.442
*DiTMC+aPE-B (9.5M) 79.9 $\pm$0.1 85.4 $\pm$0.3 0.434 $\pm$0.002 0.389 $\pm$0.002 76.5 $\pm$0.1 83.6 $\pm$0.3 0.500 $\pm$0.002 0.423 $\pm$0.004
*DiTMC+aPE-L (31.1M) 79.2 $\pm$0.1 84.4 $\pm$0.2 0.432 $\pm$0.003 0.386 $\pm$0.003 77.8 $\pm$0.1 85.7 $\pm$0.5 0.470 $\pm$0.001 0.387 $\pm$0.003
*DiTMC+PE(3)-B (8.6M) 80.8 $\pm$0.1 85.6 $\pm$0.5 0.427 $\pm$0.001 0.396 $\pm$0.001 75.3 $\pm$0.1 82.0 $\pm$0.2 0.515 $\pm$0.000 0.437 $\pm$0.003
*DiTMC+PE(3)-L (28.2M) 80.8 $\pm$0.3 85.6 $\pm$0.1 0.415 $\pm$0.003 0.376 $\pm$0.001 76.4 $\pm$0.2 82.6 $\pm$0.3 0.491 $\pm$0.002 0.414 $\pm$0.004
Table A11: Median absolute error of ensemble properties between generated and reference conformers. Best results in bold, second best underlined. Our models are marked with an asterisk “*”. Results for MCF, ET-Flow, and ours are averaged over three random seeds.
Method $E$ [kcal/mol] $\downarrow$ $\mu$ [D] $\downarrow$ $\Delta\epsilon$ [kcal/mol] $\downarrow$ $E_{\text{min}}$ [kcal/mol] $\downarrow$
OMEGA 0.68 0.66 0.68 0.69
GeoDiff 0.31 0.35 0.89 0.39
GeoMol 0.42 0.34 0.59 0.40
Torsional Diff. 0.22 0.35 0.54 0.13
MCF 0.68 $\pm$0.06 0.28 $\pm$0.05 0.63 $\pm$0.05 0.04 $\pm$0.00
ET-Flow 0.18 $\pm$0.01 0.18 $\pm$0.01 0.35 $\pm$0.06 0.02 $\pm$0.00
*DiTMC+aPE-B 0.17 $\pm$0.00 0.16 $\pm$0.01 0.27 $\pm$0.01 0.01 $\pm$0.00
*DiTMC+aPE-L 0.16 $\pm$0.02 0.14 $\pm$0.03 0.27 $\pm$0.01 0.01 $\pm$0.00
*DiTMC+PE(3)-B 0.18 $\pm$0.01 0.18 $\pm$0.01 0.27 $\pm$0.03 0.02 $\pm$0.00
*DiTMC+PE(3)-L 0.17 $\pm$0.01 0.14 $\pm$0.01 0.25 $\pm$0.01 0.01 $\pm$0.00
E.4 Generalization Results on GEOM-QM9
Additionally, we want to assess how well our proposed model architecture generalizes to unseen molecules. Here we evaluate our models trained on GEOM-DRUGS on the GEOM-QM9 dataset. We report the generalization performance in Table A12.
Table A12: Out-of-distribution generalization results on GEOM-QM9 for models trained on GEOM-DRUGS. -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined. Our models are marked with an asterisk “*”. Our results are averaged over three random seeds with standard deviation reported below. Other works do not report standard deviations.
Method COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
Mean Median Mean Median Mean Median Mean Median
ET-Flow 86.7 100.0 0.218 0.160 68.7 75.3 0.369 0.317
*DiTMC+aPE-B 84.3 $\pm$0.5 100.0 $\pm$0.0 0.209 $\pm$0.004 0.134 $\pm$0.003 69.4 $\pm$0.7 84.4 $\pm$3.2 0.329 $\pm$0.004 0.250 $\pm$0.008
*DiTMC+aPE-L 80.6 $\pm$0.5 100.0 $\pm$0.0 0.199 $\pm$0.004 0.115 $\pm$0.004 68.83 $\pm$0.2 86.2 $\pm$2.0 0.329 $\pm$0.002 0.224 $\pm$0.007
*DiTMC+PE(3)-B 84.8 $\pm$0.3 100.0 $\pm$0.0 0.205 $\pm$0.003 0.130 $\pm$0.005 69.1 $\pm$0.1 81.41 $\pm$2.9 0.329 $\pm$0.004 0.254 $\pm$0.015
*DiTMC+PE(3)-L 85.1 $\pm$0.2 100.0 $\pm$0.0 0.195 $\pm$0.004 0.115 $\pm$0.003 69.5 $\pm$0.6 82.4 $\pm$2.3 0.319 $\pm$0.05 0.245 $\pm$0.009
E.5 Generalization Results on GEOM-XL
We also study how well our proposed model architecture generalizes to unseen molecules with a large number of atoms. For this we adopt the GEOM-XL dataset containing a total of 102 molecules with more than 100 atoms presented in [46]. We report the generalization performance in Table A13.
Table A13: Out-of-distribution generalization results on GEOM-XL for models trained on GEOM-DRUGS. -R indicates Recall, -P indicates Precision. Best results in bold, second best underlined. Our models are marked with an asterisk “*”. Our results are averaged over three random seeds with standard deviation reported below. Other works do not report standard deviations.
Method AMR-R [Å] $\downarrow$ AMR-P [Å] $\downarrow$ # mols
Mean Median Mean Median
GeoDiff 2.92 2.62 3.35 3.15 -
GeoMol 2.47 2.39 3.30 3.14 -
Tor. Diff. 2.05 1.86 2.94 2.78 -
MCF - S 2.22 1.97 3.17 2.81 102
MCF - B 2.01 1.70 3.03 2.64 102
MCF - L 1.97 1.60 2.94 2.43 102
ET-Flow 2.31 1.93 3.31 2.84 102
*DiTMC+aPE-B 1.96 $\pm$0.00 1.60 $\pm$0.03 2.90 $\pm$0.00 2.48 $\pm$0.03 102
*DiTMC+aPE-L 1.88 $\pm$0.01 1.51 $\pm$0.02 2.81 $\pm$0.00 2.30 $\pm$0.02 102
*DiTMC+PE(3)-B 1.98 $\pm$0.01 1.67 $\pm$0.02 3.03 $\pm$0.00 2.60 $\pm$0.01 102
*DiTMC+PE(3)-L 1.85 $\pm$0.02 1.58 $\pm$0.03 2.93 $\pm$0.00 2.53 $\pm$0.03 102
Tor. Diff. 1.93 1.86 2.84 2.71 77
MCF - S 2.02 1.87 2.9 2.69 77
MCF - B 1.71 1.61 2.69 2.44 77
MCF - L 1.64 1.51 2.57 2.26 77
ET-Flow 2.00 1.80 2.96 2.63 75
*DiTMC+aPE-B 1.68 $\pm$0.00 1.47 $\pm$0.02 2.59 $\pm$0.00 2.24 $\pm$0.01 77
*DiTMC+aPE-L 1.56 $\pm$0.01 1.28 $\pm$0.01 2.47 $\pm$0.00 2.14 $\pm$0.01 77
*DiTMC+PE(3)-B 1.73 $\pm$0.01 1.55 $\pm$0.01 2.71 $\pm$0.00 2.35 $\pm$0.01 77
*DiTMC+PE(3)-L 1.57 $\pm$0.01 1.46 $\pm$0.01 2.60 $\pm$0.00 2.27 $\pm$0.02 77
Appendix F Additional Ablations
F.1 Index Positional Encoding (iPE)
Tab. A14 compares a variant including index positional encoding (iPE) from classic transformer architectures with our base model using aPE on QM9. Specifically, we use the node index and encode it via sinusoidal encodings into the tokens $\mathcal{H}$ before the first DiTMC block, similar to embedding the absolute positions via aPE. Since the graphs are generated from SMILES strings via rdkit and rdkit has to some extend a canonical ordering, this information can be used by the transformer architecture. However, index positional encoding breaks permutation equivariance (as we show in Tab. A14). This might be undesirable as permutation equivariance is one of the fundamental symmetries when learning on graphs. Since the ordering of atoms in a SMILES string is not uniquely defined, the trained network depends on the used framework for parsing the SMILES string or even a particular software version. We use rdkit (version 2024.9.5) for parsing SMILES strings to graphs.
Nevertheless, our model using iPE can effectively exploit the information contained in atom indices assigned by rdkit. A version of our base aPE model without atom-pair conditioning but iPE achieves comparable performance to our model using geodesic distances as atom-pair conditioning (all-pair conditioning). As our all-pair conditioning strategy is similarly or more effective than iPE but additionally preserves permutation equivariance, it should be preferred over iPE and we don’t use iPE in any of our other experiments.
Table A14: Ablating index positional encoding (iPE) on QM9 for different conditioning strategies (in brackets). To show the effect of atom permutations, we include results with randomly permuted atom indices (perm.). -R indicates Recall, -P indicates Precision. Best results in bold. Our results are averaged over three random seeds with standard deviation reported below.
Method COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
Mean Median Mean Median Mean Median Mean Median
DiTMC+aPE-B (node only) 96.3 $\pm$0.0 100.0 $\pm$0.0 0.079 $\pm$0.000 0.037 $\pm$0.000 93.2 $\pm$0.2 100.0 $\pm$0.0 0.112 $\pm$0.001 0.051 $\pm$0.000
DiTMC+aPE-B (node only), perm. 96.3 $\pm$0.0 100.0 $\pm$0.0 0.079 $\pm$0.000 0.037 $\pm$0.000 93.2 $\pm$0.2 100.0 $\pm$0.0 0.112 $\pm$0.001 0.051 $\pm$0.000
DiTMC+aPE+iPE-B (node only) 96.6 $\pm$0.4 100.0 $\pm$0.0 0.079 $\pm$0.002 0.037 $\pm$0.001 95.5 $\pm$0.1 100.0 $\pm$0.0 0.093 $\pm$0.001 0.046 $\pm$0.001
DiTMC+aPE+iPE-B (node only), perm. 82.3 $\pm$1.1 100.0 $\pm$0.0 0.229 $\pm$0.008 0.108 $\pm$0.005 60.0 $\pm$0.9 61.8 $\pm$1.4 0.493 $\pm$0.008 0.416 $\pm$0.011
DiTMC+aPE-B (node & all-pair) 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
DiTMC+aPE-B (node & all-pair), perm. 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
F.2 Gaussian vs. Harmonic Prior
As shown in Tab. A15, using the harmonic prior improves all metrics slightly for our models on GEOM-QM9. Using the harmonic prior however doesn’t seem to be a crucial ingredient for the success of our method, as differences between Gaussian and Harmonic prior appear diminishing. As the results in Tab. A15 verify, our method can also be used with a simple Gaussian prior effectively. For larger molecular graphs the expensive eigendecomposition required for the Harmonic prior could therefore be avoided, which helps scaling our approach more easily.
Table A15: Ablation of PE strategies and Gaussian (G) and Harmonic (H) prior. We report mean coverage (COV) at a threshold of $0.5\mbox{\AA}$, and mean average minimum RMSD (AMR) for Recall -R and Precision -P. Best results in bold. All results are averaged over three random seeds.
PE strategy COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
G H G H G H G H
DiTMC+aPE 96.2 96.1 0.074 0.073 95.2 95.4 0.087 0.085
DiTMC+rPE 96.0 96.3 0.073 0.070 95.2 95.7 0.084 0.080
DiTMC+PE(3) 95.7 95.7 0.069 0.068 93.5 93.4 0.090 0.089
Appendix G Analysis of training loss as a function of latent time
In this section, we provide details for the analysis in Fig. 2C in the main part of the paper. We investigate the effect of the PEs and self-attention formulations on the accuracy of the model. We therefore take pre-trained models on GEOM-QM9 and compute the training loss (as detailed in Algorithm 1) averaged over 1000 samples drawn randomly from the GEOM-QM9 validation set. We compute the loss for 30 logarithmically spaced values of $t_{i}=1-10^{x_{i}}$, where $x_{i}\in[-1.8,0]$ with uniform spacing. We skip the stochastic term in the loss as is done while sampling from the ODE.
As detailed in Fig. A4, we observe empirically that equivariance leads to a decreased loss close to the data distribution after training. This explains why our equivariant model more often succeeds to produce samples with increased fidelity, as depicted in figure Fig. 2B.
We further note, that absolute loss values for models trained with all our PE strategies decrease as latent time increases (see Fig. A4). This is expected, as conditional vector fields for each data sample will start to interact more strongly moving away from the data distribution. Our weighted loss (see Appendix A) effectively penalizes errors close to the data distribution during training and helps with keeping the error low in this important regime.
Figure A5: Misclassified exampled for SMILES classification experiment. We randomly pick 3 examples, which are misclassified by a classification head without any GNN layers. We show that GNN layers are essential for correct classification of isomers.
Figure A6: Loss as a function of time comparing different PE strategies. Results averaged over 1000 samples randomly drawn from the GEOM QM9 validation set. Left: loss relative to PE(3) as a baseline. In the important regime close to the data distribution, the model PE(3) has lower loss, yielding higher sample fidelity. Right: absolute loss values for all PEs. The loss decreases close to the data distribution for all models.
Appendix H SMILES Classification Experiment via Conditioning GNN
Table A16: We measure the discriminative power of our conditioning graph network on a training set of 1000 randomly sampled SMILES strings from the Geom QM9 validation set, as well as a toy dataset of 3 different SMILES strings. We investigate the required number of message passing layers, as well as using pre-trained weights from an end-to-end trained model.
GNN layers Weight init Trainable Accuracy (Geom QM9) Accuracy (Toy Data)
0 random trainable 0.887 0.333
1 random trainable 0.999 0.666
2 random trainable 1.000 1.000
2 random frozen 0.980 1.000
2 pre-trained frozen 1.000 1.000
A critical requirement of our DiTMC approach is the ability to disentangle distinct SMILES representations through conditioning. We investigate whether our proposed conditioning strategy is capable of learning the necessary information to distinguish between different SMILES strings for the generation of matching molecular conformers.
As a proxy evaluation task, we assess whether the conditioning network alone can function as a classifier of SMILES strings. To this end, we construct two training datasets: a toy dataset comprising three specific SMILES strings of a hydroxyl group moving along a carbon chain (C(O)CCCCCCCC, CC(O)CCCCCCC, CCC(O)CCCCCC), and a larger set consisting of 1000 randomly sampled SMILES strings drawn from the GEOM-QM9 validation set. Each SMILES string becomes a seperate class, so for each class there is exactly one example in the training data. The classification task is performed on the graph representations of the molecules, employing the same feature set and GNN architecture utilized in our conditioning graph neural network (Sec. 4.1, appendix D.2) plus a simple classification head.
We train models with a batch size of 3 for 5000 epochs (toy dataset) and batch size of 64 over 250 epochs (GEOM-QM9 subset). We report classification accuracy on the training sets directly to evaluate the model’s discriminative capacity. Furthermore, we explore whether conditioning weights obtained from an end-to-end trained model retain discriminatory power by freezing them and attaching a linear classification head.
Our results, as shown in Tab. A16, reveal that a simple linear classifier lacking message-passing capabilities fails to distinguish certain SMILES strings. Overall, our results indicate that a simple two-layer GNN effectively captures the necessary conditioning information through end-to-end training. Fig. A5 shows that without GNN layers, isomers will be misclassified.
Appendix I Architecture Details for Conditioning GNN
To transform SMILES representations to per-token conditioning information, we use a GNN that directly operates on the bond graph induced by the SMILES string. Our model processes molecular graphs by first embedding node and edge features, initially represented as one-hot vectors (a full list of features is provided in Tab. A7). These features are projected into a shared latent space using two-layer multilayer perceptrons (MLPs). For message passing, we employ a graph neural network architecture inspired by the processor described in the MeshGraphNet (MGN) framework [54].
The GNN maintains and updates both node and edge representations across multiple layers. Each message passing block consists of two main steps: first, the edge representations are updated based on the current edge representations and the representations of the connected nodes:
$\boldsymbol{e}_{ij}^{\prime}\leftarrow f_{e}(\boldsymbol{e}_{ij},\boldsymbol{v}_{i},\boldsymbol{v}_{j})$ (A53)
where $\mathbf{e}_{ij}$ and $\mathbf{v}_{i}$ denote the input edge and node representations, and $\mathbf{e}_{ij}^{\prime}$ are the updated edge representations. The learnable function $f_{e}$ is implemented as a two-layer MLP. Next, node representations $\mathbf{v}_{i}$ are updated to $\mathbf{v}_{i}^{\prime}$ using aggregated messages from neighboring edges:
$\boldsymbol{v}_{i}^{\prime}\leftarrow f_{v}\left(\boldsymbol{v}_{i},\sum_{j}\boldsymbol{e}_{ij}^{\prime}\right)$ (A54)
where $f_{v}$ is also a two-layer MLP, and the summation is over all edges ending at node $i$.
The final output of the described GNN is a set of node embeddings per atom, and a set of edge embeddings per bond. Those serve as atom-wise or bond-pair conditioning inputs to the DiT transformer, as detailed in Sec. 4.1.
Appendix J Autoguidance
We apply autoguidance, a recent technique for enhancing generative models by leveraging the outputs of a weaker model to guide sampling from a stronger one [71]. Originally proposed in the context of image diffusion, autoguidance works by introducing a degraded model with parameters $\hat{\theta}$ trained on the same data and conditioning as the main model with parameters $\theta$, but with additional constraints such as reduced capacity, increased noise or fewer gradient steps. During generation, the predicted vector field $\boldsymbol{v}^{\theta}(\boldsymbol{x},t)$ is guided using the vector field predicted by $\boldsymbol{v}^{\hat{\theta}}(\boldsymbol{x},t)$, where the discrepancy between the two models serves as a corrective signal. Intuitively, this nudges the generation process towards higher-quality samples by amplifying the difference between a less accurate and a more accurate model.
In our molecular conformer generation setup, we choose the parameters after training for 5 epochs without applying exponential moving average (EMA) for $\hat{\theta}$. We then use the difference in predicted vector fields from to bias the conformer generation trajectory. Following Ref. [72], we apply autoguidance only in the guidance interval $t\in[0,t_{max}]$ and additionally ablate $t_{max}$ as the upper bound of the guidance interval. It becomes straightforward to adapt Algorithm 2 to reflect these changes during sampling as a drop-in repacement for the predicted vector field. The guided velocity (for $t\in[0,t_{max}]$) is computed as $\boldsymbol{v}_{\eta}(\boldsymbol{x},t)=\eta\boldsymbol{v}^{\theta}(\boldsymbol{x},t)+(1-\eta)\boldsymbol{v}^{\hat{\theta}}(\boldsymbol{x},t)$, where $\eta$ is the guidance strength.
Fig. A7 shows that autoguidance yields enhanced fidelity of generated conformations, as indicated by improvements in precision metrics, but might negatively impact diversity (recall). Increasing guidance strength improves precision (COV-P, AMR-P) but may reduce recall (COV-R, AMR-R). A similar trend can be observed for the guidance interval, where increased $t_{max}$ leads to stronger effects of autoguidance. This effectively establishes a trade-off between precision and recall that can be tuned post-hoc after training with minimal effort.
Figure A7: Autoguidance performance on GEOM-QM9 using our DiTMC+aPE-B model. We evaluate the effect of autoguidance on conformer generation quality across various guidance intervals ($t_{max}$) on the x-axes, and guidance strength ($\eta$) as different colors. Dashed black lines indicate performance without guidance.
Appendix K Analysis of Sampling Trajectories
Figure A8: Comparison of molecular generation with node-only versus node- and all-pair conditioning on GEOM-QM9. Each row shows a prior sample, sampling trajectories, and final generated structure for both models; pairwise conditioning preserves bonds from the conditioning graph $\mathcal{G}$.
Fig. A8 compares the generative performance of DiTMC models trained on the GEOM-QM9 dataset under two different conditioning strategies: (1) node-only conditioning and (2) node plus pairwise conditioning. Each row in the figure corresponds to one example molecule selected from the test set. The molecules are chosen to maximize the root mean squared deviation (RMSD) between the final generated structures of the two models. This selection highlights cases where the differences between the conditioning schemes are most pronounced.
Within each row, we display a sequence of images: the initial prior sample, followed by the intermediate trajectory of the ODE sampling process over 50 sampling steps for the node-only conditioned model, and the resulting final structure (with predicted bonds rendered in yellow). This sequence is repeated for the node- and all-pair-conditioned model, allowing a side-by-side visual comparison of the generation dynamics and final outputs.
The results reveal a consistent pattern: models trained with node-only conditioning fail to preserve bonding patterns from the conditioning graph $\mathcal{G}$. This manifests as bond stretching or atom permutation in the final structure. In contrast, the pairwise-conditioned model produces geometries that adhere more closely to expected chemical structure and the given bonds. We note that if atoms are simply permuted by the model using only node conditioning, the generated structure might still be valid in terms of the combination of generated 3D positions and atom types. The degraded performance of node conditioning versus node and pair-conditioning can therefore in part be explained by the used RMSD and Coverage metrics, which are not invariant to permutations of atoms.
Our findings still underscore the importance of incorporating both node-level and pairwise features in molecular generative models, in particular when agreement with the given conditioning on a bond graph is essential.
Appendix L Visualization
Fig. A9, Fig. A10, and Fig. A11 provide a visual comparison of conformers generated by MCF, ET-Flow, and DiTMC against the corresponding ground-truth reference conformers for the GEOM-QM9, GEOM-DRUGS and GEOM-XL datasets, respectively. For each dataset, we randomly select six reference conformers from the test split and generate conformers using each method. Finally, we apply rotation alignment of the generated conformers with their corresponding reference conformer.
Figure A9: Comparison of conformers generated by MCF, ET-Flow, and DiTMC against ground-truth reference conformers from GEOM-QM9. The generated conformers are rotationally aligned with their corresponding reference conformer to facilitate comparison. From left to right: generated conformers from MCF, ET-Flow, DiTMC, and the corresponding ground-truth reference conformers.
Figure A10: Comparison of conformers generated by MCF, ET-Flow, and DiTMC against ground-truth reference conformers from GEOM-DRUGS. The generated conformers are rotationally aligned with their corresponding reference conformer to facilitate comparison. From left to right: generated conformers from MCF, ET-Flow, DiTMC, and the corresponding ground-truth reference conformers.
Figure A11: Comparison of conformers generated by MCF, ET-Flow, and DiTMC against ground-truth reference conformers from GEOM-XL. The generated conformers are rotationally aligned with their corresponding reference conformer to facilitate comparison. From left to right: generated conformers from MCF, ET-Flow, DiTMC, and the corresponding ground-truth reference conformers.

### v2

@@
 Our results highlight how architectural choices and symmetry priors affect sample quality and efficiency, suggesting promising directions for large-scale generative modeling of molecular structures.
-Code available at https://github.com/ML4MolSim/dit_mc.
-Keywords Generative Modeling  $\cdot$ Flow Matching  $\cdot$ Diffusion Transformers  $\cdot$ Geometric Deep Learning  $\cdot$ Molecular Simulation  $\cdot$ Conformer Generation
+Code is available at https://github.com/ML4MolSim/dit_mc.
 1 Introduction
@@
 Generative machine learning (ML) models offer a more targeted approach by allowing to directly sample from the space of promising conformations.
-Recent years have seen significant progress, enabled by the development of specialized architectures for the generation of molecules [1, 2, 3, 4, 5, 6, 7, 8, 9] and materials [10, 11, 12].
-This is in contrast to image and video synthesis, where the more generalized diffusion transformer (DiT) architecture [13] has become a leading model, consistently delivering strong performance and efficiency across diverse applications [14, 15, 16, 17].
-Adapting DiTs, which were originally developed for grid-structured image data, to continuous, irregular molecular geometries poses unique challenges, which need to be addressed to unlock the potential of this powerful architecture for molecular conformer generation and leverage prior advances from vision research.
+Recent years have seen significant progress, enabled by the development of specialized architectures for the generation of molecules [gebauer2019symmetry, masuda2020generating, ragoza2020learning, gebauer2022inverse, luo20213d, simonovsky2018graphvae, de2018molgan, hoogeboom2022equivariant, garcia2021n] and materials [jiao2023crystal, vecchio2024matfuse, zeni2025generative].
+This is in contrast to image and video synthesis, where the more generalized diffusion transformer (DiT) architecture [peebles2023DiT] has become a leading model, consistently delivering strong performance and efficiency across diverse applications [bao2023all, chen2024gentron, liu2024sora, ma2025latte].
+Adapting DiTs, which were originally developed for grid-structured image data, to continuous, irregular molecular geometries poses unique challenges, which need to be addressed to unlock the potential of this powerful architecture for molecular conformer generation.
 Key design questions include how to encode molecular connectivity and incorporate Euclidean symmetries, such as translational and rotational invariance/equivariance.
 In this work, we address these conceptual challenges and propose DiTMC, a new DiT-style architecture for molecular conformer generation.
-We introduce novel conditioning strategies based on molecular graphs, enabling 3D conformer generation.
+We introduce novel conditioning strategies based on molecular graphs, enabling the generation of 3D structures.
 Our modular architectural design allows us to systematically investigate the impact of different self-attention mechanisms within the DiT architecture on conformer generation quality and efficiency.
@@
 We find that simpler attention mechanisms are highly scalable and still perform competitively.
-Multiple of the tested DiT variants achieve state-of-the-art (SOTA) precision on established conformer generation benchmarks.
-Further, the ensembles of molecular structures generated by our models exhibit increased realism, as evidenced by the high accuracy of physical properties extracted from them.
+Multiple of the tested DiTMC variants achieve state-of-the-art precision on established conformer generation benchmarks.
+Moreover, the molecular structure ensembles generated by our models align more closely with physical reality, as evidenced by the high accuracy of physical properties extracted from them.
 To summarize, our work contains the following main contributions:
   • We propose two complementary conditioning strategies based on trainable conditioning tokens for (pairs of) atoms extracted from molecular graphs, which are designed to align with the architectural principles of DiTs.
+We propose to condition our self-attention formulation on geodesic graph distances extracted from molecular graphs and demonstrate that it significantly increases performance of our model.
   • We investigate the impact of different self-attention mechanisms, including standard (non-equivariant) and SO(3)-equivariant formulations, on model accuracy and performance.
 We find that including symmetries can improve the fidelity of generated samples at the price of increased computational cost during training and inference.
-  • Based on our insights, we present a simple, non-equivariant, yet expressive DiT architecture that achieves state-of-the-art (SOTA) precision and physical validity on established benchmarks.
+  • Based on our insights, we present a simple, non-equivariant, yet expressive DiT architecture that achieves state-of-the-art precision and physical validity on established benchmarks.
 Its performance improves with model scaling, making it a promising candidate for large-scale molecular conformer generation.
@@
 2 Related Work
-Generative Modeling
-Generative models create diverse, high-quality samples from an unknown data distribution.
-They are widely used in image generation [18, 19], text synthesis [20] and the natural sciences [21, 22, 23].
-Most recent approaches learn a probabilistic path from a simple prior to the data distribution, enabling both efficient sampling and likelihood estimation [24, 25, 26, 27, 28].
-This is achieved by different modeling paradigms, including flow matching [26] or denoising diffusion [24].
-(Diffusion) Transformers
-Originally introduced for natural language processing [29], transformer architectures have become SOTA in domains such as computer vision [18], and recently also found widespread adoption in quantum chemistry, e.g., for protein prediction [30, 31], 3D molecular generation [32, 33] or molecular dynamics simulations [34, 35, 36].
-In the context of generative modeling, diffusion transformers [13] (DiTs) have emerged as powerful tools incorporating conditioning tokens, e.g., to prompt image generation [13] or to design molecules and materials with desirable properties [37].
+Generative Modeling Generative models create diverse, high-quality samples from an unknown data distribution.
+They are widely used in image generation [dosovitskiy2021an, rombach2022high], text synthesis [gu2022vector] and the natural sciences [jumper2021highly, alakhdar2024diffusion, hsu2024score].
+Most recent approaches learn a probabilistic path from a simple prior to the data distribution, enabling both efficient sampling and likelihood estimation [ho2020denoising, kingma2021variational, lipman2023flow, albergo2023stochastic, kingma2023understanding].
+This is achieved by different modeling paradigms, including flow matching  [lipman2023flow] or denoising diffusion [ho2020denoising].
+(Diffusion) Transformers Originally introduced for natural language processing [vaswani2017attention], transformer architectures have become state-of-the-art in computer vision [dosovitskiy2021an], and recently also found widespread adoption in quantum chemistry, e.g., for protein prediction [abramson2024accurate, geffner2025proteina], 3D molecular generation [hassan2024etflow, lewis2024scalable] or molecular dynamics simulations [tholke2022equivariant, frank2024fastandstable, liao2024equiformerv].
+In the context of generative modeling, diffusion transformers [peebles2023DiT] (DiTs) have emerged as powerful tools incorporating conditioning tokens, e.g., to prompt image generation [peebles2023DiT] or to design molecules and materials with desirable properties [joshi2025all].
 This work applies prototypical DiTs to molecular conformer generation.
-Molecular Conformer Generation
-Molecular conformer generation aims to find atomic arrangements (Cartesian coordinates) consistent with a given molecular graph.
+Molecular Conformer Generation Molecular conformer generation aims to find atomic arrangements (Cartesian coordinates) consistent with a given molecular graph.
 Numerous ML approaches have been proposed for sampling conformers, all aiming to improve upon conventional methods.
-While early approaches were based on RDKit [38] or variational auto encoders [39, 40, 41], more recent advances employ diffusion and flow-based models [42, 43], including E-NFs [9], CGCF [44], GeoDiff [45], TorsionDiff [46], MCF [47], ET-Flow [32], and DMT [48].
+While early approaches were based on RDKit [riniker2015better] or variational auto encoders [mansimov2019molecular, simm2019generative, xu2021end], more recent advances employ diffusion and flow-based models [luo2021predicting, shi2021learning], including E-NFs [garcia2021n], CGCF [xu2021learning], GeoDiff [xu2022geodiff], TorsionDiff [jing2022torsional], MCF [wang2023swallowing], ET-Flow [hassan2024etflow], and DMT [liu2025nextmol].
+De-Novo Molecular Generation Instead of conditioning on a given molecular graph, several existing approaches tackle the problem of finding novel and stable molecules given only a set of atoms [huang2023mdm, hoogeboom2022equivariant, xu2023geometric].
+Bridging the gap between unconditional and conformer generation, universal models have been proposed that can solve a multitude of tasks.
+In particular, models within the Uni-Mol model family [zhou2023uni, ji2024uni], make use of graph geodesic distances.
+However, their design requires tracking high-dimensional pair-representations throughout the architecture, which makes it less efficient.
 3 Preliminaries
 3.1 Molecular Conformers
-A molecule can be associated with a graph $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$, where the nodes $\mathcal{V}$ correspond to atoms and the edges $\mathcal{E}$ represent chemical bonds between them.
+A molecule can be represented as a molecular graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the nodes $\mathcal{V}$ correspond to atoms and the edges $\mathcal{E}$ represent chemical bonds between them.
 The nodes and edges contain information about their types, bond orders, and additional structural features such as branches, rings, and stereochemistry.
-What is missing is the exact spatial arrangement of the $N=\lvert\mathcal{V}\rvert$ atoms as a 3D point cloud $\boldsymbol{x}\in\mathbb{R}^{N\times 3}$ in Euclidean space.
+The missing component is the exact spatial arrangement of the $N=\lvert\mathcal{V}\rvert$ atoms, represented as a 3D point cloud $\boldsymbol{x}\in\mathbb{R}^{N\times 3}$ in Euclidean space.
 Only the relative distances between atoms are relevant, as translating or rotating the entire point cloud $\boldsymbol{x}$ does not change the identity of the conformer.
-We frame conformer prediction as sampling from the conditional probability distribution $p(\boldsymbol{x}\,|\,\mathcal{G})$, which will guide the design of our model architectures (Sec.
-4).
+We frame conformer prediction as sampling from the $\mathrm{SE}(3)$-invariant conditional probability distribution $p(\boldsymbol{x}\,|\,\mathcal{G})$, which will guide the design of our model architectures in Sec.
+4.
 3.2 Conditional Flow Matching
-Starting from an easy-to-sample base distribution $q_{0}:\mathbb{R}^{d}\mapsto\mathbb{R}_{\geq 0}$, a generative process creates samples from a target distribution $q_{1}:\mathbb{R}^{d}\mapsto\mathbb{R}_{\geq 0}$ [27].
+Starting from an easy-to-sample base distribution $q_{0}:\mathbb{R}^{d}\mapsto\mathbb{R}_{\geq 0}$, a generative process creates samples from a target distribution $q_{1}:\mathbb{R}^{d}\mapsto\mathbb{R}_{\geq 0}$ [albergo2023stochastic].
 Here, $q_{1}$ models the molecular conformer data with $d=N\times 3$.
 We aim to learn a time-dependent vector field $u_{t}(\boldsymbol{x}):[0,1]\times\mathbb{R}^{N\times 3}\mapsto\mathbb{R}^{N\times 3}$, which defines an ordinary differential equation (ODE) whose solution pushes samples $\boldsymbol{x}_{0}\in\mathbb{R}^{N\times 3}$ from the prior to samples $\boldsymbol{x}_{1}\in\mathbb{R}^{N\times 3}$ from the data distribution.
-We describe this transformation in terms of a stochastic interpolant $\boldsymbol{x}_{t}$ [49, 26, 50]. A noisy sample at time $t\in[0,1]$ is defined as
+We describe this transformation in terms of a stochastic interpolant $\boldsymbol{x}_{t}$ [albergo2023building, lipman2023flow, liu2023flow].
+A noisy sample at time $t\in[0,1]$ is defined as
 $\boldsymbol{x}_{t}=(1-t)\cdot\boldsymbol{x}_{0}+t\cdot\boldsymbol{x}_{1}+\sigma\cdot\boldsymbol{\epsilon},$ (1)
@@
 Notably, stochastic interpolants enable transformations between arbitrary distributions and allow us to assess the performance of the generative process under varying prior distributions $q_{0}$.
-This contrasts with, e.g., score based diffusion methods [24, 51], which typically assume an isotropic Gaussian prior.
+This contrasts with, e.g., score based diffusion methods [ho2020denoising, song2021scorebased], which typically assume an isotropic Gaussian prior.
 The stochastic interpolant induces a deterministic trajectory of densities $p_{t}(\boldsymbol{x})$, governed by an ODE known as the probability flow:
@@
 If the vector field $u_{t}(\boldsymbol{x})$ was tractable to sample, the weights of a neural network (NN) $\boldsymbol{v}^{\theta}(\boldsymbol{x},t):[0,1]\times\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ could be optimized directly by minimizing
-$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim\mathcal{U}(0,1),\boldsymbol{x}\sim p_{t}(\boldsymbol{x})}\ \Big{|}\Big{|}u_{t}(\boldsymbol{x})-\boldsymbol{v}^{\theta}(\boldsymbol{x},t)\Big{|}\Big{|}.$ (3)
+[…]
 The learned vector field $\boldsymbol{v}^{\theta}$ could then be used to generate new samples from the target distribution by starting from $\boldsymbol{x}_{0}\sim q_{0}$ and integrating the probability flow ODE (Eq.
@@
 However, for arbitrary distributions $q_{0}$ and $q_{1}$, the objective in Eq.
-3 is computationally intractable [52].
+3 is computationally intractable [tong2024improving].
 Instead, we consider the expectation over interpolated point pairs from the two distributions.
 Eq.
-1 defines a conditional probability distribution $p_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}|(1-t)\cdot\boldsymbol{x}_{0}+t\cdot\boldsymbol{x}_{1},\sigma^{2})$, with conditional vector field $\boldsymbol{u}_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})=\boldsymbol{x}_{1}-\boldsymbol{x}_{0}$ [27].
+1 defines a conditional probability distribution $p_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}|(1-t)\cdot\boldsymbol{x}_{0}+t\cdot\boldsymbol{x}_{1},\sigma^{2})$, with conditional vector field $\boldsymbol{u}_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})=\boldsymbol{x}_{1}-\boldsymbol{x}_{0}$ [albergo2023stochastic].
 The ability to directly sample from the conditional probability via Eq.
@@
 As shown in Ref.
-[26], the gradients of the two losses coincide, $\nabla_{\theta}\mathcal{L}_{\text{FM}}=\nabla_{\theta}\mathcal{L}_{\text{CFM}}$, thereby recovering the vector field that defines the probability flow ODE in Eq.
+[lipman2023flow], the gradients of the two losses coincide, $\nabla_{\theta}\mathcal{L}_{\text{FM}}=\nabla_{\theta}\mathcal{L}_{\text{CFM}}$, thereby recovering the vector field that defines the probability flow ODE in Eq.
 2.
-Following prior work [53, 37], we reparametrize the training objective to predict noise-free data $\boldsymbol{x}_{0}^{\theta}$ directly.
-During inference, we invert the reparametrization to obtain $\boldsymbol{v}^{\theta}$ for sampling (see Appendix C).
+Following prior work [yim2023fast, joshi2025all], we reparametrize the training objective to predict noise-free data $\boldsymbol{x}_{0}^{\theta}$ directly.
+During inference, we invert the reparametrization to obtain $\boldsymbol{v}^{\theta}$ for sampling (see Appendix C for details).
 4 A New Diffusion Transformer for Molecular Conformer Sampling
@@
 3.1, this involves sampling from a conditional probability $p(\boldsymbol{x}\,|\,\mathcal{G})$, where $\mathcal{G}$ is the molecular graph representing atomic connectivity.
-Therefore, we choose our model to be a function $\boldsymbol{v}^{\theta}(\boldsymbol{x},t,\mathcal{G})$.
-We begin by initializing the node features using learnable embeddings $\boldsymbol{e}(z_{i})\in\mathbb{R}^{H}$, where $z_{i}\in\mathbb{N}_{+}$ denotes the atomic number of atom $i$ [35].
-These embeddings are then processed through multiple DiT blocks.
-Each DiT block transforms the current node features (input tokens) =$\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$ to a set of output tokens $\mathcal{H}^{\prime}$, which serve as input for the next block.
-Additional inputs consist of positional embeddings $\mathcal{P}$ (representing the position of nodes in Euclidean space), as well as conditioning tokens $\mathcal{C}$, for example for encoding the time ($\boldsymbol{c}^{t}$) and graph structure ($\boldsymbol{c}^{\mathcal{G}}$).
-The central operation in each DiT block is a self-attention update,
-$\boldsymbol{h}_{i}=\boldsymbol{h}_{i}+\text{ATT}(\mathcal{H},\mathcal{P},\mathcal{C}^{\text{Pair}})_{i}\,,$ (5)
-followed by a node-wise Multi-Layer Perceptron (MLP).
-Here, $\mathcal{C}^{\text{Pair}}$ are special conditioning tokens acting on pairs of atoms (see Sec.
-4.1).
-Similar to applications of DiTs in image synthesis, adaptive layer norm (AdaLN) and adaptive scale (AdaScale) [13] are used for conditioning (see Fig.
-1A) based on per-atom bias and scaling parameters
-$\alpha_{1i},\beta_{1i},\gamma_{1i},\alpha_{2i},\beta_{2i},\gamma_{2i}=\text{MLP}(\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}})\,.$ (6)
-The final output is a 3D vector (“velocity”) per atom, which is extracted from a readout layer.
-The whole architecture is summarized in Fig.
-1A.
-In the following, we describe the individual components of the architecture in more detail.
-Training details are given in Appendix A.
-4.1 Conditioning
-Each DiT block receives conditioning signals that encode information about the time step $t$, as well as atom- and pair-wise information derived from a molecular graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$.
-Pair-wise information is used during the self-attention update (see Eq.
-5 and Sec.
-4.3), whereas atom-wise information and the time conditioning signal are injected via adaptive layer norms (see Eq.
-6).
-Time conditioning
-The current time $t$ of the latent state $\boldsymbol{x}_{t}$ is encoded via a two-layer MLP as
+Therefore, we choose our model to be a function $\boldsymbol{v}^{\theta}(\boldsymbol{x},t,\mathcal{G})$, where the final output is a 3D velocity vector per atom, which is extracted from a readout layer.
+Next, we outline the key components of the DiTMC architecture, with an overview shown in Fig. 1A. Training and architectural details are provided in Appendix B and Appendix D, respectively.
+4.1 Conditioning Tokens
+We begin by defining the conditioning tokens in DiTMC.
+Each DiTMC block receives a time conditioning token $\boldsymbol{c}^{t}\in\mathbb{R}^{H}$, as well as atom-wise conditioning tokens $\mathcal{C}_{\text{atom}}^{\mathcal{G}}=\{\boldsymbol{c}_{i}^{\mathcal{G}}\in\mathbb{R}^{H}\,|\,i\in[N]\}$, and pair-wise conditioning tokens $\mathcal{C}_{\text{pair}}^{\mathcal{G}}=\{\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}\,|\,i,j\in[N]\}$ derived from $\mathcal{G}=(\mathcal{V},\mathcal{E})$.
+Time conditioning The current time $t$ of the latent state $\boldsymbol{x}_{t}$ is encoded via a two-layer MLP as
 $\boldsymbol{c}^{t}=\text{MLP}(t)\,.$ (5)
-Atom-wise conditioning
-Atom-wise conditioning tokens are obtained from a GNN inspired by the processor module of the MeshGraphNet (MGN) framework [54]
+Atom-wise conditioning Atom-wise graph conditioning tokens are obtained from a GNN inspired by the processor module of the MeshGraphNet (MGN) framework [pfaff2020learning] as
 $\boldsymbol{c}_{i}^{\mathcal{G}}=\text{GNN}_{\text{node}}(\mathcal{V},\mathcal{E})\,$, (6)
-where $\boldsymbol{c}_{i}^{\mathcal{G}}$ denotes the final node representation for atom $i$. See Appendix I for details on the GNN.
-Pair-wise conditioning
-In addition to node representations, the GNN also produces edge-level representations, which can be used to define pair-wise conditioning tokens:
-[…]
-These tokens only capture interactions between bonded atoms, i.e., when $(i,j)\in\mathcal{E}$.
-Conditioning tokens for non-bonded pairs are set to a learnable vector $\boldsymbol{\bar{c}}^{\mathcal{G}}$.
-Since self-attention operates on all atom pairs $(i,j)$, even if they are not connected by a chemical bond, we also define an alternative pair-wise conditioning method inspired by the Graphormer architecture [55]
-[…]
-where $s(i,j)$ denotes the graph geodesic (i.e., the shortest path between atoms $i$ and $j$ via edges in $\mathcal{G}$).
-This formulation allows conditioning on all atom pairs, even if they are not directly connected.
-In the following, we refer to these two variants as bond-pair (Eq.
-9) and all-pair (Eq.
-10) conditioning, respectively (see Sec.
-5 for empirical comparisons between both methods).
-4.2 Positional Embeddings ( PE )
-We inject information about the current atomic positions =$\mathcal{R}=\{\vec{r}_{1},\dots\vec{r}_{N}\,|\,\vec{r}_{i}\in\mathbb{R}^{3}\}$ of the latent state $\boldsymbol{x}_{t}$ via positional embeddings (PEs).
-To that end, we assume the positions to have zero center of mass (see section A.1).
-As mentioned in Sec.
-3.1, conformer identity is determined only by the relative positions of atoms (i.e., it is invariant under rigid translations and rotations), but ensuring that these symmetries are preserved within the network often incurs computational overhead.
-Therefore, we examine a representative range of PEs that vary in the number of Euclidean symmetries they respect by construction, which affects how the latent representations transform under translations and rotations.
-This includes absolute and relative PEs, motivated by a similar distinction made in other application domains [56, 57].
-Inspired by recent advances in geometric deep learning [58, 59, 60], we also propose a PE strategy, obeying all Euclidean symmetries.
-Our aim is to find an acceptable tradeoff between model accuracy, efficiency and scalability, by scanning the design space of PE strategies.
-Absolute Positional Embeddings ( aPE )
-Following Refs. [31, 37] they are calculated as
+where $\boldsymbol{c}_{i}^{\mathcal{G}}$ denotes the final node representation for atom $i$. See Appendix D.3 for details on the GNN.
+Pair-wise conditioning We define pair-wise graph conditioning tokens inspired by the Graphormer architecture [ying2021graphormer] as
+$\boldsymbol{c}_{ij}^{\mathcal{G}}=\text{MLP}\left(s(i,j)\right),$ (7)
+where $s(i,j)$ denotes the graph geodesic, i.e., the shortest path between atoms $i$ and $j$ in $\mathcal{G}$.
+This formulation allows conditioning on all atom pairs, including those not directly connected by a bond.
+4.2 Positional Embeddings
+To encode the atomic positions $\mathcal{R}=\{\vec{r}_{1},\dots,\vec{r}_{N}\,|\,\vec{r}_{i}\in\mathbb{R}^{3}\}$ of the latent state $\boldsymbol{x}_{t}$, we use positional embeddings (PEs).
+We examine a representative range of positional embeddings that vary in the number of Euclidean symmetries they respect by construction, which affects how the latent representations transform under translations and rotations.
+We denote the set of positional embeddings by $\mathcal{P}$, where $\mathcal{P}=\{\boldsymbol{p}_{i}\,|\,i\in[N]\}$ for the atom-wise (aPE) embeddings, and $\mathcal{P}=\{\boldsymbol{p}_{ij}\,|\,i,j\in[N],\,i\neq j\}$ for pair-wise (rPE or PE(3)) embeddings.
+Without loss of generality, we assume that the positions are centered such that the center of mass vanishes (see Appendix A).
+Absolute Positional Embeddings Following Refs. [geffner2025proteina, joshi2025all], atom-wise absolute Positional Embeddings (aPE) are calculated as
 $\boldsymbol{p}_{i}^{\text{aPE}}=\text{MLP}(\vec{r}_{i})\,$, (8)
-where $\vec{r}_{i}\in\mathbb{R}^{3}$ is the Cartesian position of the $i$-th atom.
-This kind of PE is neither rotationally nor translationally invariant, and serves as a baseline without any symmetry constraints.
-Relative Positional Embedding ( rPE )
-We use displacements vectors $\vec{r}_{ij}=\vec{r}_{i}-\vec{r}_{j}$ to build pairwise relative PEs as
-$\boldsymbol{p}_{ij}^{\text{rPE}}=\text{MLP}(\vec{r}_{ij})\,.$ (12)
-This formulation ensures translational invariance but not rotational invariance.
-Euclidean Positional Embedding ( PE(3) )
-Adapting ideas from equivariant message passing neural networks like PaiNN [61] or NequIP [62], we construct SO(3)-equivariant pairwise PEs as a concatenation of $L+1$ components
+such that $\boldsymbol{p}_{i}^{\text{aPE}}\in\mathbb{R}^{H}$.
+This kind of positional embedding does not preserve rotational nor translational invariance, and serves as a baseline without any symmetry constraints.
+Relative Positional Embeddings We use displacements vectors $\vec{r}_{ij}=\vec{r}_{i}-\vec{r}_{j}$ to build pairwise relative Positional Embeddings (rPE) as
+$\boldsymbol{p}_{ij}^{\text{rPE}}=\text{MLP}(\vec{r}_{ij})\,$, (9)
+such that $\boldsymbol{p}_{ij}^{\text{rPE}}\in\mathbb{R}^{H}$. This formulation ensures translational invariance but not rotational invariance.
+Euclidean Positional Embeddings Adapting ideas from equivariant message passing neural networks like PaiNN [schutt2021equivariant] or NequIP [batzner20223], we construct SO(3)-equivariant pairwise Euclidean Positional Embeddings (PE(3)) as a concatenation of $L+1$ components
 $\boldsymbol{p}_{ij}^{\text{PE(3)}}=\bigoplus_{\ell=0}^{L}\boldsymbol{\phi}_{\ell}(r_{ij})\,\odot\boldsymbol{Y}_{\ell}(\hat{r}_{ij})\,$, (10)
-where $\boldsymbol{\phi}_{\ell}:\mathbb{R}\mapsto\mathbb{R}^{1\times H}$ is a radial filter function, $\hat{r}=\vec{r}/r$, and $\boldsymbol{Y}_{\ell}\in\mathbb{R}^{(2\ell+1)\times 1}$ are spherical harmonics of degree $\ell=0\dots L$.
+where $\boldsymbol{\phi}_{\ell}:\mathbb{R}\mapsto\mathbb{R}^{1\times H}$ is a radial filter function, $\hat{r}=\vec{r}/r$, and $\boldsymbol{Y}_{\ell}\in\mathbb{R}^{(2\ell+1)\times 1}$ are spherical harmonics of degree $\ell=0,\dots,L$.
 The element-wise multiplication $\odot$’ between radial filters and spherical harmonics is understood to be “broadcasting” along axes with size $1$, such that $(\boldsymbol{\phi}_{\ell}\,\odot\boldsymbol{Y}_{\ell})\in\mathbb{R}^{(2\ell+1)\times H}$ and (after concatenation) $\boldsymbol{p}_{ij}^{\text{PE(3)}}\in\mathbb{R}^{(L+1)^{2}\times H}$.
-Under rotation of the input positions, these PEs transform equivariantly (see Appendix B.2.2).
+Under rotation of the input positions, these positional embeddings transform equivariantly (see Appendix E).
 Moreover, because displacement vectors are used as inputs, the embeddings are also invariant to translations.
 As a result, they respect the full set of Euclidean symmetries relevant to molecular geometry.
-4.3 Self-Attention Operation
-For ease of notation, we only describe self-attention with a single head, but employ multi-head attention [29] with $n_{\text{heads}}$ heads in our experiments.
-All self-attention blocks rely on query, key and value vectors, which are obtained from the input tokens =$\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\ |\ \boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$ as
-$\boldsymbol{q}=\boldsymbol{W}_{q}\boldsymbol{\tilde{h}}\,,\hskip 25.0pt\boldsymbol{k}=\boldsymbol{W}_{k}\boldsymbol{\tilde{h}}\,,\hskip 25.0pt\boldsymbol{v}=\boldsymbol{W}_{v}\boldsymbol{\tilde{h}}\,,$ (14)
-where $\boldsymbol{W}_{q},\boldsymbol{W}_{k},\boldsymbol{W}_{v}\in\mathbb{R}^{H\times H}$ are trainable weight matrices and $\boldsymbol{\tilde{h}}$ is either identical to $\boldsymbol{h}$ or combines it with a PE (see below).
-We define a slightly modified similarity kernel
-$\text{sim}(\boldsymbol{q},\boldsymbol{k},\boldsymbol{u})=\exp\Bigg{(}\frac{\boldsymbol{q}^{\intercal}\cdot(\boldsymbol{k}\,\odot\boldsymbol{u})}{\sqrt{H}}\Bigg{)},$ (15)
-where $\boldsymbol{u}\in\mathbb{R}^{H}$ is used to inject additional information, e.g., conditioning signals and/or positional embeddings, and $\odot$’ denotes element-wise multiplication.
-Depending on the subset of Euclidean symmetries we aim to incorporate, we adopt slightly different formulations of the self-attention mechanism, as detailed below.
-Standard Self-Attention
-For absolute and relative PEs, we slightly modify standard self-attention to allow injecting pair-wise information into the values in addition to using our modified similarity kernel:
-$\text{ATT}(\mathcal{H})_{i}=\frac{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})\cdot(\boldsymbol{v}_{j}\odot\boldsymbol{u}_{ij})}{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})}\,.$ (16)
-Queries, keys, and values are obtained with Eq.
-14 from different (position-encoded) tokens $\boldsymbol{\tilde{h}}_{i}$ depending on the chosen PEs; further, the injected pair-wise information $\boldsymbol{u}_{ij}$ differs:
-[…]
-Here $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$ are pair-wise graph conditioning tokens (see Eqs.
-9 and 10) and $\boldsymbol{p}_{i}^{\text{aPE}}\in\mathbb{R}^{H}$ and $\boldsymbol{p}_{ij}^{\text{rPE}}\in\mathbb{R}^{H}$ are the absolute and relative PEs described above (see Eqs.
-11 and 12).
-SO(3)-Equivariant Self-Attention
-To preserve all Euclidean symmetries throughout the network, every token must transform equivariantly.
-One way to achieve this is by separating out the rotational degrees of freedom, encoding them with irreducible representations of the rotation group $\mathrm{SO(3)}$.
-This introduces a “degree-axis” of size $(L+1)^{2}$, which encodes angular components of increasing order.
-The maximum degree $L$ is chosen to ensure high fidelity at a reasonable computational cost.
-For example, setting $L=1$ restricts the representation to scalars and vectors, as used in models like PaiNN [61] or TorchMDNet [34].
-An $\mathrm{SO(3)}$-equivariant formulation of self-attention is then given as
-$\text{ATT}_{\text{SO(3)}}(\mathcal{H})_{i}=\frac{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})\cdot(\hat{\boldsymbol{u}}_{ij}\otimes\boldsymbol{v}_{j})}{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})}\,,$ (18)
-where equivariant queries, keys and values can be calculated similarly to Eq.
-14 and $\otimes$’ denotes a Clebsch-Gordan (CG) tensor product contraction [60].
-The dot-product in the similarity measure is taken along both feature and degree axes, such that the overall update preserves equivariance (see Appendix for details).
-Tokens and scaling vectors are calculated as
-[…]
-where $\boldsymbol{\phi}(r_{ij})\in\mathbb{R}^{(L+1)^{2}\times H}$ is a radial filter, and the element-wise products with the pair-wise conditioning tokens $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{1\times H}$ are broadcast along the degree axis.
-Importantly, the $12\ell+$ 1 subcomponents of the radial filter for degree $\ell$ are obtained by repeating per-degree filter functions $\boldsymbol{\phi}_{\ell}(r_{ij})\in\mathbb{R}^{1\times H}$ along the degree axis to preserve equivariance (see also Eq.
-13).
-5 Experiments
-Datasets and Metrics
-We conduct our experiments on the GEOM dataset [63], comprising QM9 (133,258 small molecules) and AICures (304,466 drug-like molecules).
-Reference conformers are generated using CREST [64].
-Drug-like molecules exhibit greater structural diversity, including more rotatable bonds and multiple stereocenters.
-Data splits are taken from Ref.
-[65].
-We evaluate our models’ ability to generate accurate and diverse conformers using Average Minimum RMSD (AMR) and Coverage (COV), measuring Recall (ground-truth coverage) and Precision (generation accuracy).
-A generated conformer is considered valid if it falls within a specified RMSD threshold of any reference conformer ($\delta=0.5\mbox{\AA}$ for GEOM-QM9 and $\delta=0.75\mbox{\AA}$ for GEOM-DRUGS).
-Following prior work, we generate $2K$ conformers per test molecule with $K$ reference structures.
-Appendix D.3 provides further details on the calculation of metrics.
-Following ET-Flow [32] and GeomMol [65] we also apply chirality correction (see Appendix D.4).
-Conditioning Strategies
-To assess the impact of different graph conditioning strategies, we compare “node only”, “node & bond-pair”, and “node & all-pair” conditioning (see Sec.
-4.1 for details) against a baseline without graph conditioning.
-We find all variants to be effective, with “node & all-pair” conditioning performing best (see Appendix Tab.
-A8).
-It outperforms both, “node only” and “node & bond-pair” conditioning, highlighting the strength of our all-pair-conditioning strategy based on graph geodesics.
-This underlines the importance of deriving conditioning tokens for all atom pairs, not just those connected by edges in the molecular graph, which lack global information about the graph structure.
+4.3 DiTMC Block
+Based on the positional embedding strategies introduced in Sec.
+4.2, we can define different DiTMC blocks that preserve the extent of Euclidean symmetries encoded in the embeddings throughout the model.
+Each DiTMC block transforms a set of input tokens $\mathcal{H}$ into a set of output tokens $\mathcal{H}^{\prime}$, which serve as input for the next block.
+For aPE and rPE, we have $\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$, and for PE(3), we have $\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{(L+1)^{2}\times H}\}$.
+In each DiTMC block, we inject time-based, as well as graph-based atom-wise and pair-wise conditioning information via the conditioning tokens introduced in Sec.
+4.1.
+We use the pair-wise graph conditioning tokens $\mathcal{C}_{\text{Pair}}^{\mathcal{G}}$ during the self-attention update,
+$\boldsymbol{h}_{i}=\boldsymbol{h}_{i}+\text{ATT}(\mathcal{H},\mathcal{P},\mathcal{C}_{\text{Pair}}^{\mathcal{G}})_{i}\,$, (11)
+where we employ a standard self-attention mechanism for aPE and rPE and an SO(3)-equivariant self-attention mechanism for PE(3).
+Additionally, we use the time conditioning token $\boldsymbol{c}^{t}$ and the atom-wise graph conditioning tokens $\mathcal{C}_{\text{atom}}^{\mathcal{G}}=\{\boldsymbol{c}_{i}^{\mathcal{G}}\in\mathbb{R}^{H}\,|\,i\in[N]\}$ to obtain per-atom bias and scaling parameters,
+$\alpha_{1i},\beta_{1i},\gamma_{1i},\alpha_{2i},\beta_{2i},\gamma_{2i}=\text{MLP}(\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}})\,$, (12)
+and apply adaptive layer norm (AdaLN) and adaptive scale (AdaScale) [peebles2023DiT] for conditioning (see Fig. 1A) similar to applications of DiTs in image synthesis.
+We provide details on the non-equivariant DiTMC blocks based on aPE and rPE in Appendix D.1, and on the SO(3)-equivariant DiTMC block based on PE(3) in Appendix D.2.
 Table 1: Results on GEOM-QM9 for different generative models (parameter counts in parentheses).
@@
 Our results are averaged over three random seeds.
-See Tab.
+See Appendix Tab.
 A9 for results including standard deviations.
@@
 *DiTMC+PE(3)-B (8.6M) 95.7 100.0 0.068 0.021 93.4 100.0 0.089 0.032
-Figure 2: Analysis of equivariant (PE(3)) and non-equivariant (aPE, rPE) model formulations.
-(A) Mean coverage recall (COV-R) versus threshold $\delta$ (root mean square deviation (RMSD) to any reference conformer).
+Figure 2: Analysis of SO(3)-equivariant (PE(3)) and non-equivariant (aPE, rPE) model formulations on GEOM-QM9.
+(A) Mean Coverage Recall (COV-R) versus root mean square deviation (RMSD) threshold $\delta$ to any reference conformer.
 (B) Histogram of the minimal RMSD per generated sample.
-(C) Loss as a function of latent time $t$ relative to PE(3) loss.
-Ablating Self-Attention and Positional Embedding Strategies
-The modular structure of DiTMC allows efficient exploration of the design space through variations in the PE strategy and associated attention blocks (see Sec.
+(C) Loss as a function of latent time $t$ relative to PE(3) loss (see Appendix J for details).
+5 Experiments
+Datasets and Metrics We conduct our experiments on the GEOM dataset [axelrod2022geom], comprising QM9 (133,258 small molecules) and AICures (304,466 drug-like molecules).
+Reference conformers are generated using CREST [pracht2024crest].
+Drug-like molecules exhibit greater structural diversity, including more rotatable bonds and multiple stereocenters.
+The data splits are taken from Ref.
+[ganea2021geomol].
+We evaluate our models’ ability to generate accurate and diverse conformers using average minimum RMSD (AMR) and coverage (COV), measuring recall (ground-truth coverage) and precision (generation accuracy).
+A generated conformer is considered valid if it falls within a specified RMSD threshold of any reference conformer ($\delta=0.5\mbox{\AA }$ for GEOM-QM9 and $\delta=0.75\mbox{\AA }$ for GEOM-DRUGS).
+Following prior work, we generate $2K$ conformers per test molecule with $K$ reference structures.
+Appendix G.3 provides further details on the calculation of metrics.
+Following ET-Flow [hassan2024etflow] and GeomMol [ganea2021geomol] we also apply chirality correction (see Appendix G.4).
+Ablating Self-Attention and Positional Embedding Strategies The modular structure of DiTMC allows efficient exploration of the design space through variations in the positional embeddings and associated attention blocks (see Sec.
 4).
-All DiTMC models share hyperparameters, differing only in the choice of PEs and self-attention formulation.
-For the architecture using PE(3), we reduce the number of heads to match the parameter count across models.
-Tab.
-1 shows results on GEOM-QM9.
-All DiTMC models produce diverse, high quality samples, outperforming the current SOTA in all AMR metrics and COV-P.
-This demonstrates the broad applicability of our modular design and conditioning.
+We define three model variants, DiTMC+aPE, DiTMC+rPE, and DiTMC+PE(3), which differ only in the choice of positional embedding and self-attention formulation.
+Architectural details can be found in Appendix D.
+On GEOM-QM9, our models produce diverse, high quality samples, outperforming the current state-of-the-art across all AMR-R, AMR-P, and COV-P metrics, demonstrating the broad applicability of our modular design and conditioning (see Tab.
+1).
 We use the harmonic prior introduced in Ref.
-[66] throughout, which yields improved results (see Appendix Tab.
-A15).
-Therefore, we use it in all following experiments.
-Probing the effect of Euclidean symmetries
-Our PEs form a hierarchy based on the extent of Euclidean symmetry incorporated by construction.
+[stark2023harmonic] throughout, which yields improved results compared to the Gaussian prior (see Appendix Tab.
+A13), and therefore adopt it in all subsequent experiments.
+Notably, our models maintain competitive performance even when using the Gaussian prior, contrary to the findings in Ref.
+[hassan2024etflow].
+Probing the effect of Euclidean symmetries Our PEs form a hierarchy based on the extent of Euclidean symmetry incorporated by construction.
 This enables a systematic evaluation of how incorporating symmetry affects model behavior.
-We summarize our findings below.
+We summarize our findings on GEOM-QM9 below.
 Equivariance improves the fidelity of samples.
-We compute COV-R as a function of threshold $\delta$ for different PEs and self-attention blocks (Fig.
+We analyze COV-R Mean as a function of RMSD threshold $\delta$ for all DiTMC variants (see Fig.
 2A).
-SO(3)-equivariant attention with PE(3) outperforms the other variants at low $\delta$, indicating that many of the generated geometries closely match the ground-truth structures.
-This appears as a leftward shift in the distribution of the minimal RMSD found per generated structure (Fig.
-2B) and aligns with the observation that the PE(3) model gives better AMR-R and AMR-P values (see Tab.
+The SO(3)-equivariant DiTMC+PE(3) outperforms the non-equivariant DiTMC+aPE and DiTMC+rPE at low $\delta$, indicating that many of the generated conformers closely match the ground-truth structures.
+This appears as a leftward shift in the distribution of the minimal RMSD found per generated structure (as shown in Fig.
+2B) and aligns with the observation that DiTMC+PE(3) achieves better AMR-R and AMR-P values (as reported in Tab.
 1).
-To better understand this behavior, we analyze the loss over time $t$ and find that models with non-equivariant PEs exhibit higher error near the data distribution ($t=1$) (Fig.
+To better understand this behavior, we examine the loss over time $t$ and find that the non-equivariant models exhibit higher error near the data distribution ($t=1$) (see Fig.
 2C).
@@
 The benefit of higher fidelity comes at increased computational cost.
-During training, the equivariant PE(3) model is approximately $3.5$ times slower than models using aPE or rPE, while at inference time, the factor is $\sim 3$ (see Appendix Tab.
-A5).
-All models use the same number of layers and differ only in the number of heads per layer in order to match the total parameter count.
+During training, the equivariant DiTMC+PE(3) is approximately $3.5$ times slower than the non-equivariant DiTMC+aPE and DiTMC+rPE, and about $3$ times slower at inference (see Fig.
+3B).
+All models use the same number of layers and differ only in the number of heads per layer in order to match the total parameter count, as discussed in Appendix D.
+Conditioning Strategies To assess the impact of graph conditioning, we compare different conditioning strategies.
+As further discussed in Appendix I.2, we compare conditioning solely on atom-wise information (Eq.
+6) with an extended scheme that also incorporates pair-wise geodesic graph distances (Eq.
+7).
+In Tab.
+2, we show ablation results on GEOM-DRUGS.
+We also ablate conditioning on pairwise information extracted from our conditioning GNN for each edge corresponding to a chemical bond (see Appendix I.2).
+We find all variants to be effective, but our proposed combination of geodesic distances and atom-wise information to perform best.
+This experiment underlines the importance of deriving conditioning tokens for all atom pairs, not just those connected by edges in the molecular graph (i.e.
+by chemical bonds), which lack global information about the graph structure.
+Table 2: Ablation of conditioning strategies using DiTMC+aPE-B on GEOM-DRUGS.
+-R indicates Recall, -P indicates Precision.
+Best results in bold.
+Our results are averaged over three random seeds.
+See Appendix Tab.
+A14 for results on GEOM-QM9.
+Method COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
+Mean Median Mean Median Mean Median Mean Median
+No conditioning 16.6 4.7 1.132 1.110 5.15 1.2 1.849 1.845
+Atom-wise 72.8 77.5 0.555 0.565 55.6 53.3 0.762 0.716
+Atom-wise & Pair-wise 79.9 85.4 0.434 0.389 76.5 83.6 0.500 0.423
+Model Scaling To investigate model scaling, we define a small (“S”), base (“B”), and large (“L”) DiTMC+aPE model variant (see Appendix D).
+Each model is trained with identical training hyperparameters on the GEOM-DRUGS dataset.
+We find strong relative improvements for all metrics (up to 54.1% for COV-P Mean) when scaling DiTMC+aPE-S to DiTMC+aPE-B (see Fig.
+3A).
+Scaling DiTMC+aPE-B to DiTMC+aPE-L yields relative improvements, which however might be smaller than expected given the strong increase in performance from the small to the base model.
+Prior work has shown that to avoid diminishing returns, dataset and model sizes must be scaled together [kaplan2020scaling, hoffmann2022training].
+Scaling only the model yields limited benefits as soon as model size saturates given the amount of data.
+Strong improvements of DiTMC+aPE-L over DiTMC+aPE-B are observed in terms of sampling fidelity (as indicated by higher COV-R and COV-P Mean values at small thresholds $\delta$) but differences start to vanish at $\delta=0.75\mbox{\AA }$ (see Fig.
+4A).
+Moreover, we obtain relative improvement in terms of generalization to larger and previously unseen molecules from the GEOM-XL dataset [jing2022torsional] between 3.1% (for AMR-P Mean) and 7.3% (for AMR-P Median), i.e., DiTMC+aPE-L exhibits significantly stronger out-of-distribution performance compared to the smaller DiTMC+aPE-B (see Tab.
+5).
+Figure 3: (A) Coverage (COV) mean and Absolute Minimum RMSD (AMR) mean for DiTMC+aPE models of increasing model capacity on GEOM-DRUGS.
+(B) Training and inference time for different positional embedding (PE) and associated self-attention strategies.
+(C) Recall and precision AMR mean for the different DiTMC models on GEOM-DRUGS.
+For panels (B) and (C) we use results from the base (“B”) variant of each model.
+Drug-like Molecules For the following experiments, we restrict our analysis to the base (“B”) and large (“L”) DiTMC variants.
+In contrast to the smaller GEOM-QM9 dataset, the size of GEOM-DRUGS allows us to explore the design space of DiTMC under a more realistic setting, where training is contstrained by a fixed compute budget [hoffmann2022training].
+Our budget of 9 GPU days allows training DiTMC+aPE-L and DiTMC+rPE-L for 50 epochs, and DiTMC+PE(3)-L for 10 epochs.
+For consistency, we train the base models for the same number of epochs as their large counterparts.
+All DiTMC variants achieve state-of-the-art performance on GEOM-DRUGS for COV-P and AMR-P (see Tab.
+3).
+Importantly, even the smaller DiTMC+aPE-B and DiTMC+PE(3)-B models outperform ET-Flow-SS of similar model size across all metrics, underlining the effectiveness of our approach.
 Table 3: Results on GEOM-DRUGS for different generative models (parameter counts in parentheses).
@@
 Our results are averaged over three random seeds.
-See Tab.
+See Appendix Tab.
 A10 for results including standard deviations.
@@
 ET-Flow - SS (8.3M) 79.6 84.6 0.439 0.406 75.2 81.7 0.517 0.442
-*DiTMC+aPE-L 79.2 84.4 0.432 0.386 77.8 85.7 0.470 0.387
-*DiTMC+PE(3)-L 80.8 85.6 0.415 0.376 76.4 82.6 0.491 0.414
-Figure 3: (A) Time per gradient step for the DiTMC S, M and L variants using aPE and PE(3), measured on an Nvidia H100 GPU (batch size 128 on Geom QM9).
-(B) Coverage precision and recall as function of threshold $\delta$ and (C) mean AMR metrics for base (“B”) and large (“L”) model of DiTMC+aPE.
-Drug-like Molecules
-Finally, we evaluate our architecture on the challenging GEOM-DRUGS dataset.
-We analyze its behavior when scaling model capacity, assess physical validity via ensemble property prediction, and examine out-of-distribution performance.
-We focus on the models using aPE and PE(3), as they represent the extreme cases of not including any, and including all Euclidean symmetries in the architecture design, respectively.
-Model Scaling.
-To analyze how DiTMC behaves as model capacity scales, we define a small base (“B”) and a large (“L”) model variant (see Appendix B).
-The compute cost per gradient step for the equivariant PE(3) model is substantially higher than for the aPE model, despite having a similar number of parameters (Fig.
-3A).
-Even the base PE(3) model requires more compute than the large aPE model.
-This enables training aPE models for more epochs under the same compute budget (see Appendix Fig.
-A4).
-Scaling the capacity of aPE-L yields substantial gains in precision metrics (Fig.
-3B and C).
-For recall, sample fidelity improves with larger model capacity (as indicated by larger COV-R at small thresholds) but differences start to vanish at $\delta=0.75\mbox{\AA}$ (Fig.
-3B).
-Extrapolation performance also improves for the larger model (see Tab.
+*DiTMC+aPE-B (9.5M) 79.9 85.4 0.434 0.389 76.5 83.6 0.500 0.423
+*DiTMC+rPE-B (9.6M) 79.3 84.6 0.444 0.400 77.2 84.6 0.492 0.414
+*DiTMC+PE(3)-B (8.6M) 80.8 85.6 0.427 0.396 75.3 82.0 0.515 0.437
+*DiTMC+aPE-L (28.2M) 79.2 84.4 0.432 0.386 77.8 85.7 0.470 0.387
+*DiTMC+rPE-L (28.3M) 78.7 84.1 0.438 0.388 78.1 86.4 0.466 0.381
+*DiTMC+PE(3)-L (31.1M) 80.8 85.6 0.415 0.376 76.4 82.6 0.491 0.414
+Figure 4: (A) Coverage (COV) mean as function of Root Mean Square Deviation (RMSD) threshold $\delta$ and (B) average minimum RMSD (AMR) mean vs.
+time per conformer for DiTMC+aPE and other state-of-the-art models.
+Per model markers from left to right correspond to 5, 10, 20, and 50 sampler steps following Refs.
+[hassan2024etflow, wang2023swallowing].
+Note, that the original MCF paper reports results with two different samplers.
+Benchmark results (Tab.
+3) are obtained with DDPM sampler (1000 steps) and AMR vs.
+time results are reported for DDIM sampler (5–50 steps).
+(C) Comparison of conformers generated by MCF, ET-Flow, and DiTMC against ground-truth reference conformers from GEOM-XL.
+Generated conformers are rotationally aligned with their corresponding reference conformer.
+Coverage vs.
+RMSD Threshold.
+We further analyse the COV-R Mean and COV-P Mean as a function of RMSD threshold $\delta$ for DiTMC+aPE-B and DiTMC+aPE-L (see Fig.
+4A).
+For small thresholds ($\delta<0.4\mbox{\AA }$) both DiTMC+aPE models outperform all other methods for coverage recall and precision.
+For larger thresholds ($\rho\geq 0.4\mbox{\AA }$) MCF-L starts to outperform aPE models in terms of COV-R Mean.
+For COV-P Mean, DiTMC+aPE-B and DiTMC+aPE-L perform better than all other methods for all reasonably small thresholds ($\delta<1.2\mbox{\AA }$).
+In particular for the most relevant regime where thresholds are small, we see a strong benefit due to model scaling.
+We find similar results for DiTMC+rPE and DiTMC+PE(3) (see Appendix Fig.
+A9 and Fig.
+A10).
+Pareto Front.
+Additionally, we investigate the Pareto front of accuracy and computational efficiency, by plotting model accuracy as a function of wall clock time per generated conformer, following Refs.
+[hassan2024etflow, wang2023swallowing] (see Appendix G.5 for details).
+As measure for accuracy we consider the average minimum RMSD (AMR), since it is independent of the RMSD threshold.
+We report results of DiTMC+aPE-B and DiTMC+aPE-L for AMR-R Mean and AMR-P Mean in Fig.
+4B.
+For AMR-P Mean, both DiTMC+aPE models shift the whole Pareto front, yielding higher accuracy at lower computational cost.
+Even higher accuracies (at the cost of compute time) can be obtained by scaling DiTMC+aPE-B to DiTMC+aPE-L.
+For recall, both DiTMC+aPE models shift the Pareto front for little compute times, but most accurate results at increased cost are obtained by MCF-L.
+Similar results are obtained for DiTMC+rPE (see Appendix Fig.
+A12), but benefits for DiTMC+PE(3) are limited due to high computational cost of equivariant operations (see Appendix Fig.
+A13).
+Ensemble Properties.
+To complement the RMSD-based geometric evaluation, we report the median absolute error (MAE) of ensemble properties between generated and reference conformers, following the protocol of MCF [wang2023swallowing] and ET-Flow [hassan2024etflow] (see Appendix G.6 for details).
+Our models predict ensemble properties more accurately than all baselines, highlighting the physical validity of our generated structures (see Tab.
 4).
-Similar trends are observed for both base and large variants of PE(3) (see Appendix Fig.
-A4).
-This allows our aPE-L model to achieve SOTA results on GEOM-DRUGS for all precision metrics (see Tab.
-2).
-For recall, aPE-L achieves better results than ET-Flow, but worse than DMT-L and MCF-L.
+In particular, MCF, which shows better performance for recall metrics, is outperformed by a large margin (up to a factor of four for energy).
+The strong performance of the aPE-B and aPE-L models underlines the potential of achieving high physical validity without any geometric priors, which have been hypothesized to be one of the reasons for ET-Flow outperforming MCF in the ensemble property task [hassan2024etflow].
 Table 4: Median absolute error of ensemble properties between generated and reference conformers for GEOM-DRUGS.
-$E$, $\Delta\epsilon$, $E_{\text{min}}$ are in kcal/mol, and $\mu$ is in debye.
 Best results in bold, second best underlined; our models are marked with an asterisk (*).
 Results for MCF, ET-Flow, and ours are averaged over three random seeds.
-See Tab.
-A11 for results including standard deviations.
-Method $E$$\downarrow$ $\mu$$\downarrow$ $\Delta\epsilon$$\downarrow$ $E_{\text{min}}$$\downarrow$
-OMEGA 0.68 0.66 0.68 0.69
-Torsional Diff. 0.22 0.35 0.54 0.13
-MCF-L 0.68 0.28 0.63 0.04
-ET-Flow 0.18 0.18 0.35 0.02
-*DiTMC+aPE-B 0.17 0.16 0.27 0.01
-*DiTMC+aPE-L 0.16 0.14 0.27 0.01
-Table 4: Out-of-distribution generalization results on GEOM-XL for models trained on GEOM-DRUGS reporting median RMSD in Å.
+See Appendix Tab.
+A11 for additional results including standard deviations.
+Method $E$ [kcal/mol] $\downarrow$ $\mu$ [D] $\downarrow$ $\Delta\epsilon$ [kcal/mol] $\downarrow$ $E_{\text{min}}$ [kcal/mol] $\downarrow$
+MCF-L (242M) 0.68 0.28 0.63 0.04
+ET-Flow (8.3M) 0.18 0.18 0.35 0.02
+*DiTMC+aPE-B (9.5M) 0.17 0.16 0.27 0.01
+*DiTMC+aPE-L (28.2M) 0.16 0.14 0.27 0.01
+Generalization Performance Finally, we assess how well our models, trained on GEOM-DRUGS, generalize to larger and previously unseen molecules using the GEOM-XL dataset [jing2022torsional].
+It comprises 102 molecules with more than 100 atoms, whereas the molecules in the training set contain only 44 atoms on average.
+Following MCF [wang2023swallowing] we report results for all 102 molecules and a subset of 77 molecules.
+ET-Flow uses a slightly different subset of 75 molecules.
+Our aPE models perform on par with or better than the previously best-performing method MCF-L (see Tab.
+5), while using approximately $8$ times and $25$ times fewer parameters for aPE-L and aPE-B, respectively.
+The other baselines, such as ET-Flow, are outperformed by a larger margin (see Appendix Tab.
+A12).
+Table 5: Out-of-distribution generalization results on GEOM-XL for models trained on GEOM-DRUGS.
+-R indicates Recall, -P indicates Precision.
 Best results in bold, second best underlined; our models are marked with an asterisk (*).
 Our results are averaged over three random seeds.
-See Tab.
-A13 for results including standard deviations.
-Method AMR-R $\downarrow$ AMR-P $\downarrow$
-102 molecules
-MCF - L (242M) 1.60 2.43
-*DiTMC+aPE-B 1.60 2.49
-*DiTMC+aPE-L 1.51 2.30
-77 molecules
-MCF - L (242M) 1.51 2.26
-*DiTMC+aPE-B 1.47 2.24
-*DiTMC+aPE-L 1.28 2.14
-Ensemble properties.
-To complement RMSD-based geometric evaluation with a chemically meaningful assessment, we report the median absolute error (MAE) of ensemble properties between generated and reference conformers.
-The RMSD metric can penalize conformers that differ due to rotations or atom reordering, but are otherwise chemically equivalent.
-For our analysis we follow the protocol of MCF [47] and ET-Flow [32] (see Appendix E.3) and compare energy $E$, dipole moment $\mu$, HOMO-LUMO gap $\Delta\epsilon$, and minimum energy $E_{\text{min}}$.
-Our aPE-L model predicts ensemble properties more accurately than all baselines, highlighting the physical validity of our generated structures.
-In particular, MCF-L, which shows better performance for recall metrics, is outperformed by a large margin (up to a factor of $4$ for energy).
-Generalization performance.
-We assess the generalization of our model to larger and unseen molecules using the GEOM-XL dataset [46].
-It comprises 102 molecules of size $N>100$ atoms, whereas the molecules in the training data contain only $N=44$ atoms on average.
-Following MCF [47] we report the generalization performance for all 102 molecules and a subset of 77 molecules.
-Our models perform on par or better to the previously best-performing method MCF-L while using only a fraction of the parameters (Tab.
-4).
-Other baselines (like ET-Flow) are outperformed by a larger margin (see Appendix Tab.
-A13).
+See Appendix Tab.
+A12 for additional results including standard deviations.
+Method 75 / 77 mols 102 mols
+AMR-R [Å] $\downarrow$ AMR-P [Å] $\downarrow$ AMR-R [Å] $\downarrow$ AMR-P [Å] $\downarrow$
+Mean Median Mean Median Mean Median Mean Median
+MCF-L (242M) 1.64 1.51 2.57 2.26 1.97 1.60 2.94 2.43
+ET-Flow (8.3M) 2.00 1.80 2.96 2.63 2.31 1.93 3.31 2.84
+*DiTMC+aPE-B (9.5M) 1.68 1.47 2.59 2.24 1.96 1.60 2.90 2.48
+*DiTMC+aPE-L (28.2M) 1.56 1.28 2.47 2.14 1.88 1.51 2.81 2.30
 6 Summary and Limitations
@@
 This modular framework enables a rigorous exploration of different positional embedding and self-attention strategies, allowing us to identify scalable generative architectures that perform competitively with prior methods on standard benchmarks.
-Our models achieve SOTA results on GEOM-QM9 and GEOM-DRUGS, excelling in both precision and physical validity.
-Through ablation studies, we assess the impact of incorporating Euclidean symmetries into DiTs.
+Our models achieve state-of-the-art results on GEOM-QM9 and GEOM-DRUGS, excelling in both precision and physical validity, and show strong generalization to larger, previously unseen molecules from the GEOM-XL dataset.
+Exemplary generations for GEOM-XL are shown in Fig.
+4C, with additional samples provided in Appendix M.
+Through ablation studies, we assess the impact of incorporating Euclidean symmetries into DiTMC.
 While such symmetries improve performance, they also increase computational cost.
 Notably, simpler non-equivariant variants remain highly effective.
-These findings allow us to develop an efficient, accurate, and scalable DiT architecture suitable for large-scale conformer generation.
+These findings allow us to develop an efficient, accurate, and scalable architecture suitable for large-scale conformer generation.
 Nonetheless, some limitations persist.
@@
 Moreover, the training process depends on high-quality ground-truth conformers, which may be unavailable in some cases.
+An interesting direction for future work, is extending the scope of our model beyond conformer generation, for example to de-novo molecular generation [hoogeboom2022equivariant, xu2023geometric, alakhdar2024diffusion].
+However, this would require a re-design of our proposed conditioning strategies, as those are specifically tailored to molecular graphs as inputs.
+In terms of computational cost, DiTMC scales quadratic in the number of atoms, resulting in large computational cost for bigger systems.
+Therefore, integrating recently developed Euclidean fast attention mechanisms [frank2024euclidean], poses an interesting direction for further research.
 Finally, while our analysis advances understanding of equivariance within DiT-based generative models, drawing broader conclusions would require further study.
@@
 2022-0-00984, Development of Artificial Intelligence Technology for Personalized Plug-and-Play Explanation and Verification of Explanation).
-We further want to thank Romuald Elie for helpful discussion.
-Appendix A Training
+We further want to thank Romuald Elie, Michael Plainer, Adil Kabylda, Khaled Kahouli, Stefan Gugler, Martin Michajlow, Christoph Bornett, Johannes Maeß, Leon Werner and Maximilian Eißler for helpful discussion.
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: [N/A]
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: [N/A]
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14. Justification: [N/A]
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: [N/A]
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: [N/A]
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: [N/A]
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: [N/A]
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: [N/A]
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: [N/A]
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [No]
+  49.
+Justification: While it can be expected that the general field of AI4Science will have societal impact, our work is focused on the development of a new deep learning architecture in this field.
+To have a direct societal impact, further studies under realistic applications, i.e.
+wet-lab experiments would be required.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: [N/A]
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: [N/A]
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: [N/A]
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: [N/A]
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: [N/A]
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: [N/A]
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Appendix A From SE(3) to SO(3) Invariance
+The target data distribution of molecular conformers $p_{1}(\boldsymbol{x})$ is $\mathrm{SE}(3)$-invariant, i.e.
+it does not change under translations and rotations of the input.
+Following Ref.
+[yim2023se3diffusion], one can define an $\mathrm{SE}(3)$-invariant measure on $\mathrm{SE}(3)^{N}$ by keeping the center of mass fixed at zero, which can be achieved via the centering operation from Eq.
+A15.
+This defines a subgroup $\mathrm{SE}(3)^{N}_{0}$, called centered $\mathrm{SE}(3)$.
+It can then be shown, that one can define an $\mathrm{SE}(3)$-invariant measure on $\mathrm{SE}(3)^{N}_{0}$ by constructing an $\mathrm{SO}(3)$-invariant (rotationally invariant) measure on $\mathrm{SE}(3)^{N}_{0}$.
+As a consequence, it is then sufficient to learn an $\mathrm{SO}(3)$-equivariant vector field on the space of centered input positions (see also Ref.
+[kohler2020equivariantflows]).
+This is achieved by centering $\boldsymbol{x}_{0}$, $\boldsymbol{x}_{1}$ and $\boldsymbol{z}$ for the calculation of the interpolant.
+Moreover, the neural network output (predicted velocities) and the clean target $\boldsymbol{x}_{1}$ must be centered to have zero center of mass.
+We discuss the implications for training in Section B and for sampling in Section C.
+Appendix B Training
 Algorithm 1 describes the computation of the training loss for our flow matching objective.
-We start by sampling from the prior $\boldsymbol{x}_{0}\sim p_{0}(\boldsymbol{x})$, from the data $\boldsymbol{x}_{1}\sim p_{1}(\boldsymbol{x})$, and from a Gaussian distribution $\boldsymbol{\epsilon}\sim N(\boldsymbol{x};0,\boldsymbol{I})$.
-The interpolant is then constructed as
+We start by sampling from the prior $\boldsymbol{x}_{0}\sim p_{0}(\boldsymbol{x})$, the data distribution $\boldsymbol{x}_{1}\sim p_{1}(\boldsymbol{x})$, and a Gaussian distribution $\boldsymbol{\epsilon}\sim N(\boldsymbol{x};0,\boldsymbol{I})$.
+The interpolant is then defined as
 $\boldsymbol{x}_{t}=(1-t)\cdot\boldsymbol{x}_{0}+t\cdot\boldsymbol{x}_{1}+\sigma\cdot\boldsymbol{\epsilon}\,$, (A13)
 where $\sigma\in\mathbb{R}_{>0}$ is a non-zero noise scaling parameter.
-Instead of predicting the conditional vector field $\boldsymbol{u}_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})$ directly, we choose to reparametrize the network such that it predicts the clean sample $\boldsymbol{x}_{1}$.
+Rather than directly predicting the conditional vector field $\boldsymbol{u}_{t}(\boldsymbol{x}|\boldsymbol{x}_{0},\boldsymbol{x}_{1})$, we choose to reparametrize the network such that it predicts the clean sample $\boldsymbol{x}_{1}$.
 Similar to Ref.
-[53], we add a weighting term $1/(1-t)^{2}$ to encourage the model to accurately capture fine details close to the data distribution.
-This gives raise to the following loss function
+[yim2023fast], we add a weighting term $1/(1-t)^{2}$ to encourage the model to accurately capture fine details close to the data distribution.
+This gives rise to the following loss function
 $\mathcal{L}=\frac{1}{(1-t)^{2}}\|\boldsymbol{x}_{1}^{\theta}(\boldsymbol{x}_{t},t,\mathcal{G})-\boldsymbol{x}_{1}\|^{2}\,$, (A14)
-where $t\in(0,1)$ denotes the point of time in the interpolant $\boldsymbol{x}_{t}\in\mathbb{R}^{N\times 3}$, $\boldsymbol{x}_{1}\in\mathbb{R}^{N\times 3}$ is the clean geometry and $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denotes the molecular graph.
+where $t\in(0,1)$ denotes the timestep in the interpolant $\boldsymbol{x}_{t}\in\mathbb{R}^{N\times 3}$, $\boldsymbol{x}_{1}\in\mathbb{R}^{N\times 3}$ is the clean geometry, and $\mathcal{G}=(\mathcal{V},\mathcal{E})$ denotes the molecular graph.
 The full algorithm is summarized in Algorithm 1.
 Geometry Alignment
-Given a set of vectors $\mathcal{U}=\{\vec{u}_{1},\dots,\vec{u}_{N}\,|\,\vec{u}_{i}\in\mathbb{R}^{3}\}$ associated with a point cloud $\boldsymbol{x}\in\mathbb{R}^{N\times 3}$ we define a centering operation for the $i$-th row
-$\mathrm{Center}(\boldsymbol{x})_{i}=\vec{u}_{i}-\sum_{j=1}^{N}\vec{u}_{j}\,,$ (A22)
+Given a set of vectors $\mathcal{U}=\{\vec{u}_{1},\dots,\vec{u}_{N}\,|\,\vec{u}_{i}\in\mathbb{R}^{3}\}$ associated with a point cloud $\boldsymbol{x}\in\mathbb{R}^{N\times 3}$, we define a centering operation for the $i$-th row
+$\mathrm{Center}(\boldsymbol{x})_{i}=\vec{u}_{i}-\dfrac{1}{N}\sum_{j=1}^{N}\vec{u}_{j}\,$, (A15)
 which removes global drift in $\boldsymbol{x}$.
@@
 $\mathrm{RotationAlign}(\boldsymbol{x}_{A},\boldsymbol{x}_{B})_{i}=\boldsymbol{R}_{\text{opt}}\boldsymbol{x}_{iA},$ (A16)
-where $\boldsymbol{R}_{\text{opt}}\in\mathbb{R}^{3\times 3}$ is the optimal rotation matrix, minimizing the root mean square deviation (RMSD) between the positions of point clouds A and B.
-Here, we employ the Kabsch-Algorithm.
-For full geometry alignment “$\mathrm{GeometryAlign}(\boldsymbol{x}_{A},\boldsymbol{x}_{B})$”, we do the following
+where $\boldsymbol{R}_{\text{opt}}\in\mathbb{R}^{3\times 3}$ is the optimal rotation matrix, minimizing the root mean square deviation (RMSD) between the positions of point clouds A and B, which can be obtained via the Kabsch algorithm [kabsch1976solution].
+The full geometry alignment operation $\mathrm{GeometryAlign}(\boldsymbol{x}_{A},\boldsymbol{x}_{B})$”, is given by
 $\boldsymbol{x}_{A}\leftarrow\mathrm{Center}(\boldsymbol{x}_{A})$ (A17)
@@
 $\boldsymbol{x}_{A}\leftarrow\mathrm{RotationAlign}(\boldsymbol{x}_{A},\boldsymbol{x}_{B})$ (A19)
-In words, the operations from above first center both point clouds at the origin and then align them as much as possible via a rotation.
-This procedure also minimizes the path length of a linear interpolation between the point clouds A and B.
+In words, both point clouds are first centered at the origin and then optimally aligned by rotation.
+This procedure minimizes the path length of a linear interpolation between the point clouds A and B.
 Algorithm 1 Conditional Flow Matching Training Loss
@@
 2:
-3:$\boldsymbol{x}_{0}\sim p_{0}$, $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$, $t\sim\mathcal{U}(0,1)$, $\boldsymbol{R}\sim\mathrm{SO}(3)$
-4:$\boldsymbol{x}_{0},\boldsymbol{x}_{1}\leftarrow\mathrm{GeometryAlign}(\boldsymbol{x}_{0},\boldsymbol{x}_{1})$triangleright This centers $\boldsymbol{x}_{0}$ and $\boldsymbol{x}_{1}$ and rotation-aligns $\boldsymbol{x}_{0}$ to $\boldsymbol{x}_{1}$.
-5:$\boldsymbol{\epsilon}\leftarrow\mathrm{Center}(\boldsymbol{\epsilon})$
-6:$\boldsymbol{x}_{t}\leftarrow(1-t)\boldsymbol{x}_{0}+t\boldsymbol{x}_{1}+\sigma\boldsymbol{\epsilon}$
-7:$\boldsymbol{x}_{t}\leftarrow\mathrm{ApplyRotation}(\boldsymbol{R},\boldsymbol{x}_{t})$
-8:$\boldsymbol{x}_{1}\leftarrow\mathrm{ApplyRotation}(\boldsymbol{R},\boldsymbol{x}_{1})$
-9:$\hat{\boldsymbol{x}}_{1}\leftarrow\boldsymbol{x}_{1}^{\theta}(\boldsymbol{x}_{t},t,\mathcal{G})$
-10:$\hat{\boldsymbol{x}}_{1}\leftarrow\mathrm{Center}(\hat{\boldsymbol{x}}_{1})$
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 11:
@@
 Data Augmentation
-One can construct an $\mathrm{SE}(3)$-invariant density by learning an $\mathrm{SO}(3)$-equivariant vector field on centered $\mathrm{SE}(3)$ (see section A.1).
-However, only DiTMC with PE(3) is $\mathrm{SO}(3)$-equivariant, whereas aPE and rPE violate $\mathrm{SO}(3)$-equivariance.
+The target data distribution of molecular conformers $p_{1}(\boldsymbol{x})$ is $\mathrm{SE}(3)$-invariant, i.e.
+it does not change under translations and rotations of the input.
+One can construct an $\mathrm{SE}(3)$-invariant density by learning an $\mathrm{SO}(3)$-equivariant vector field on centered $\mathrm{SE}(3)$ (see Section A).
+However, only DiTMC+PE(3) is $\mathrm{SO}(3)$-equivariant, whereas aPE and rPE violate $\mathrm{SO}(3)$-equivariance.
 Therefore, we learn equivariance approximately during training, using data augmentation.
@@
 Noise Scaling Parameter
-We ablated the noise scaling parameter $\sigma$ for both GEOM QM9 and drugs experiments, comparing a larger value of $0.5$ and a smaller value of $0.05$.
-We set the $\sigma$ parameter to the value that empirically worked best for each dataset: $0.05$ for Geom QM9 and $0.5$ for GEOM drugs.
+We ablated the noise scaling parameter $\sigma$ on GEOM-QM9 and GEOM-DRUGS, comparing a larger value of $0.5$ and a smaller value of $0.05$.
+We set $\sigma$ to the value that empirically worked best for each dataset: $0.05$ for GEOM-QM9 and $0.5$ for GEOM-DRUGS.
 Optimizer and Hyperparameters
-We use the AdamW optimizer (weight decay 0.01) with initial learning rate of $\mu_{0}=10^{-5}$.
-It is increased up to $\mu_{\text{max}}$ via a linear learning rate warmup up over the first 1% of training steps.
-Afterwards, it is decreased to $\mu_{\text{min}}$ via a cosine decay schedule.
-We use $\mu_{\text{max}}=3\times 10^{-4}$ and $\mu_{\text{min}}=0$ for GEOM-QM9 and $\mu_{\text{max}}=1\times 10^{-4}$ and $\mu_{\text{min}}=1\times 10^{-5}$ for GEOM-DRUGS.
-We use a batch size of 128 for all data sets and models.
-All models on QM9 are trained for 250 epochs in total.
-For GEOM-DRUGS we determine the maximal number of epochs from the epochs the larger model “L” can perform within a fixed compute budget of 9 days.
-Thus, we train the PE(3) variants for 10 epochs and the aPE variants for 50 epochs.
-Fig.
-A4A shows the number of gradient steps per day that can be performed with the different DiTMC variants.
-A.1 From SE(3) to SO(3) Invariance
-The target data distribution of molecular conformers $p_{1}(\boldsymbol{x})$ is $\mathrm{SE}(3)$-invariant, i.e.
-it does not change under translations and rotations of the input.
+We use the AdamW optimizer (weight decay 0.01) with batch size of 128 and learning rate of $\mu_{\text{max}}=3\times 10^{-4}$ for GEOM-QM9 and $\mu_{\text{max}}=1\times 10^{-4}$ for GEOM-DRUGS.
+First, we increase the initial learning rate of $\mu_{0}=10^{-5}$ up to $\mu_{\text{max}}$ via a linear learning rate warmup over the first 1% of training steps.
+Afterwards, the learning rate is decreased via a cosine decay schedule to $\mu_{\text{min}}=0$ for GEOM-QM9 and $\mu_{\text{min}}=1\times 10^{-5}$ for GEOM-DRUGS.
+Compute Budget and Training Times
+All models on GEOM-QM9 are trained for 250 epochs, which requires 2 days of training on Nvidia H100 GPU for aPE and rPE models and almost 4 days for PE(3).
+For GEOM-DRUGS, we fix the total compute budget per model to nine days on a single NVIDIA H100 GPU, due to computational constraints.
+Within this budget, we can train aPE-L and rPE-L for 50 epochs and PE(3)-L for 10 epochs.
+To ensure consistency within each PE strategy, the base (“B”) model are trained for the same number of epochs as the corresponding large (“L”) model.
+Appendix C Sampling
+For sampling, we use a simple Euler scheme with 50 steps to sample from the associated ordinary differential equation (ODE) as described in Algorithm 2.
+Since during training we predict the clean sample $\boldsymbol{x}_{1}$, we re-parametrize the velocity required for the integration as
+$\boldsymbol{v}^{\theta}_{t}(\boldsymbol{x}_{t},t,\mathcal{G})=\frac{\boldsymbol{x}^{\theta}_{1}(\boldsymbol{x}_{t},t,\mathcal{G})-\boldsymbol{x}_{t}}{1-t},$ (A21)
+where $\boldsymbol{x}^{\theta}_{1}(\boldsymbol{x}_{t},t,\mathcal{G})$ is the original output of DiTMC.
+To ensure $\mathrm{SE}(3)$ invariance of the probability path from an (approximately) $\mathrm{SO}(3)$-equivariant velocity predictor, we center the prior $\boldsymbol{x}_{0}\sim p_{0}(\boldsymbol{x})$ and the prediction of DiTMC in each ODE step.
+Algorithm 2 ODE Sampling
+1:Model $\boldsymbol{x}_{1}^{\theta}$, Graph $\mathcal{G}$, steps $N>0$
+2:
+[…]
+[…]
+[…]
+6:for $n\leftarrow 0$ to N1N-1 do
+7:  $\Delta t\leftarrow t_{n+1}-t_{n}$ $\triangleright$ Compute step size.
+8:  $\hat{\boldsymbol{x}}_{1}\leftarrow\boldsymbol{x}_{1}^{\theta}(\boldsymbol{x}_{t_{n}},t_{n},\mathcal{G})$
+9:  $\hat{\boldsymbol{x}}_{1}\leftarrow\mathrm{Center}(\hat{\boldsymbol{x}}_{1})$
+10:  $\boldsymbol{v}\leftarrow(\hat{\boldsymbol{x}}_{1}-\boldsymbol{x}_{t_{n}})/(1-t_{n})$
+11:  $\boldsymbol{x}_{t_{n+1}}\leftarrow\boldsymbol{x}_{t_{n}}+\Delta t\cdot\boldsymbol{v}$ $\triangleright$ Euler step.
+12:end for
+13:
+14:return $\boldsymbol{x}_{1}$
+Appendix D Architectural Details
+In this section, we describe the architecture of DiTMC and the conditioning GNN in more detail.
+All DiTMC models rely on conditioning tokens $\mathcal{C}$, for the time $\boldsymbol{c}^{t}\in\mathbb{R}^{H}$, per-atom $\boldsymbol{c}_{i}^{\mathcal{G}}\in\mathbb{R}^{H}$, and per atom-pair $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$.
+See main text Section 4.1 for more details.
 Following Ref.
-[67], one can define an $\mathrm{SE}(3)$-invariant measure on $\mathrm{SE}(3)^{N}$ by keeping the center of mass fixed at zero, which can be achieved via the centering operation from Eq.
-A22.
-This defines a subgroup $\mathrm{SE}(3)^{N}_{0}$, called centered $\mathrm{SE}(3)$.
-It can then be shown, that one can define an $\mathrm{SE}(3)$-invariant measure on $\mathrm{SE}(3)^{N}_{0}$ by constructing an $\mathrm{SO}(3)$-invariant (rotationally invariant) measure on $\mathrm{SE}(3)^{N}_{0}$.
-As a consequence, it is then sufficient to learn an $\mathrm{SO}(3)$-equivariant vector field on the space of centered input positions (see also Ref.
-[68]).
-This is achieved by centering $\boldsymbol{x}_{0}$, $\boldsymbol{x}_{1}$ and $\boldsymbol{z}$ for the calculation of the interpolant.
-Moreover, the neural network output (predicted velocities) and the clean target $\boldsymbol{x}_{1}$ must be centered to have zero center of mass.
-We discuss the implications for sampling in section C.
-A.2 Compute and Training Times
+[peebles2023DiT], we use adaptive layer norm (AdaLN) and adaptive scale (AdaScale) to include per-atom conditioning tokens based on time $t$ and molecular graph information.
+To that end, we construct conditioning tokens
+$\boldsymbol{c}_{i}=\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}}.$ (A22)
 Tab.
-A5 compares training and inference time for our models (see also Fig.
-3A for a visual comparison).
-Despite having a similar parameter count, the equivariant PE(3)-B model is approximately $3.5$ times slower than models using aPE or rPE, while at inference time, the factor is $\sim 3$.
-The effect is even more pronounced for our large models, where the equivariant model is almost 5 times slower during training compared to its non-equivariant counterpart with a similar number of parameters.
-Even our large non-equivariant model is significantly faster to train and sample from than the equivariant base model with far fewer parameters.
-All our models on GEOM QM9 are trained for 250 epochs, which requires 2 days of training on Nvidia H100 GPU for aPE and rPE models and almost 4 days for PE(3).
-Due to compute constraints, our non-equivariant (DiTMC+aPE) models are trained for 50 epochs on GEOM drugs.
-The equivariant models are trained for 10 epochs on GEOM drugs resulting in a similar compute budget for training in terms of GPU hours.
-Our maximal compute budget per model is 9 days on a single Nvidia H100 GPU.
-Table A5: Architectural details as well as training and inference time (in ms) for different PE strategies on QM9.
-Times are measured with batch size 128 on a single Nvidia H100 GPU.
+A6 contains details on the MLPs used throughout our architecture, while Tab.
+A7 summarizes architectural details, as well as training and inference times for all DiTMC variants.
+Note that the number of attention heads in DiTMC+PE(3) is adjusted to match the total parameter count of the corresponding DiTMC+aPE and DiTMC+rPE variants, while all other hyperparameters are identical.
+Table A6: Architectural details for MLPs used in the model.
+The feature dimension is given as $H=n_{\text{heads}}\cdot d_{\text{head}}$ where $n_{\text{head}}$ is the number of heads and $d_{\text{head}}$ is the number of features per head.
+Name Layers Hidden Dim Out Dim Activation Input
+DiTMC Block 2 $4H$ $H$ GELU Tokens
+SO(3) DiTMC Block 2 $4H$ $H$
+gated
+GELU
+Tokens
+Time and Atom Cond. 1 $6H$ SiLU $\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}}$
+Bond pair 2 $H$ $H$ SiLU $\boldsymbol{c}_{ij}^{\mathcal{G}}$
+Time embedding 2 $H$ $H$ SiLU $t\in[0,1]$
+Shortest-hop embedding 2 $H$ $H$ SiLU Hop distance
+aPE embedding 2 $H$ $H$ SiLU Abs. positions $\vec{r}_{i}$
+rPE embedding 2 $H$ $H$ SiLU Rel. positions $\vec{r}_{ij}$
+GNN embedding 2 $H$ $H$ SiLU Node/edge features
+Table A7: Architectural details for different PE strategies on GEOM-QM9 and GEOM-DRUGS.
+Times are measured on GEOM-QM9 with batch size 128 on a single Nvidia H100 GPU.
 $T$ means number of transformer layers, $n_{\text{heads}}$ number of heads, $d_{\text{head}}$ number of features per head in the attention update, and $T_{MGN}$ number of layers for the conditioning mesh graph net.
@@
 Model $T$ $n_{\text{heads}}$ $d_{\text{head}}$ $T_{MGN}$ $H$ Train [ms] Infer [ms]
+DiTMC+aPE-S (1M) 2 4 32 1 128 5.8 2.0
 DiTMC+aPE-B (9.5M) 6 8 32 2 256 19.2 8.1
@@
 DiTMC+aPE-L (28.2M) 8 12 32 3 384 32.7 9.5
+DiTMC+rPE-L (28.3M) 8 12 32 3 384 33.5 10.1
 DiTMC+PE(3)-L (31.1M) 8 10 32 3 320 151.6 41.8
-Appendix B Architecture Details
-Tab. A5 shows the architectural details for our base and large model variants. Tab. A6 contains details about the MLPs we use throughout our architecture.
-Table A6: Architecture details for MLPs used in the model.
-The feature dimension is given as $H=n_{\text{heads}}\cdot d_{\text{head}}$ where $n_{\text{head}}$ is the number of heads and $d_{\text{head}}$ is the number of features per head.
-The use of gated GELU ensures equivariance (see section Sec.
-B.4).
-Name Layers Hidden Dim Out Dim Activation Input
-DiT Block 2 $4H$ $H$ GELU Tokens
-SO(3) DiT Block 2 $4H$ $H$ gated GELU Tokens
-Time and Atom Cond. 1 $6H$ SiLU $\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}}$
-Bond pair 2 $H$ $H$ SiLU $\boldsymbol{c}_{ij}^{\mathcal{G}}$
-Time embedding 2 $H$ $H$ SiLU $t\in[0,1]$
-Shortest-hop embedding 2 $H$ $H$ SiLU Hop distance
-aPE embedding 2 $H$ $H$ SiLU Abs. ositions $\vec{r}_{i}$
-rPE embedding 2 $H$ $H$ SiLU Rel. positions $\vec{r}_{ij}$
-GNN embedding 2 $H$ $H$ SiLU Node/edge features
-We first describe the building blocks of our non-equivariant DiTMC implementation and then its equivariant counterpart.
-Both architectures rely on conditioning tokens $\mathcal{C}$, for the time $\boldsymbol{c}^{t}\in\mathbb{R}^{H}$, per-node $\boldsymbol{c}_{i}^{\mathcal{G}}\in\mathbb{R}^{H}$, per bond-pair $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$ and for all-pairs $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$.
-See main text section 4.1 for more details.
-Following Ref.
-[13], we use adaptive layer norm (AdaLN) and adaptive scale (AdaScale) to include per-node conditioning based on time $t$ and molecular graph information.
-To that end, we construct conditioning tokens
-$\boldsymbol{c}_{i}=\boldsymbol{c}^{t}+\boldsymbol{c}_{i}^{\mathcal{G}}.$ (A28)
-B.1 Non Equivariant DiT
-In the non-equivariant DiTMC formulations with aPE and rPE, we have the following set of tokens =$\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$.
-B.1.1 Adaptive Layer Normalization and Adaptive Scale
+D.1 Non-Equivariant DiTMC
+In the non-equivariant DiTMC formulations based on aPE and rPE, we have the following set of tokens $\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$.
+Initial token representations are obtained via $\boldsymbol{h}_{i}=\textbf{e}(z_{i})+\boldsymbol{p}_{i}^{\text{aPE}}$ for aPE and $\boldsymbol{h}_{i}=\boldsymbol{e}(z_{i})$ for rPE, where $\boldsymbol{e}(z_{i})\in\mathbb{R}^{H}$ denotes a learnable embedding based on the atomic number $z_{i}\in\mathbb{N}_{+}$ of atom $i$ [frank2024fastandstable].
+D.1.1 Self-Attention Operation
+For ease of notation, we only describe self-attention with a single head, but employ multi-head attention [vaswani2017attention] with $n_{\text{heads}}$ heads in our experiments.
+All self-attention blocks rely on query, key and value vectors, which are obtained from the input tokens $\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\ |\ \boldsymbol{h}_{i}\in\mathbb{R}^{H}\}$ as
+$\boldsymbol{q}=\boldsymbol{W}_{q}\boldsymbol{h}\,,\hskip 25.0pt\boldsymbol{k}=\boldsymbol{W}_{k}\boldsymbol{h}\,,\hskip 25.0pt\boldsymbol{v}=\boldsymbol{W}_{v}\boldsymbol{h}\,$, (A23)
+where $\boldsymbol{W}_{q},\boldsymbol{W}_{k},\boldsymbol{W}_{v}\in\mathbb{R}^{H\times H}$ are trainable weight matrices. We define a slightly modified similarity kernel
+$\text{sim}(\boldsymbol{q},\boldsymbol{k},\boldsymbol{u})=\exp\Bigg(\frac{\boldsymbol{q}^{\intercal}\cdot(\boldsymbol{k}\,\odot\boldsymbol{u})}{\sqrt{H}}\Bigg),$ (A24)
+where $\boldsymbol{u}\in\mathbb{R}^{H}$ is used to inject additional information, e.g., conditioning signals and/or positional embeddings, and $\odot$’ denotes element-wise multiplication.
+For absolute and relative PEs, we slightly modify standard self-attention to allow injecting pair-wise information into the values in addition to using our modified similarity kernel
+$\text{ATT}(\mathcal{H},\mathcal{P},\mathcal{C}_{\text{Pair}}^{\mathcal{G}})_{i}=\frac{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})\cdot(\boldsymbol{v}_{j}\odot\boldsymbol{u}_{ij})}{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})}\,$, (A25)
+where queries, keys, and values are obtained with Eq. A23 and the injected pair-wise information $\boldsymbol{u}_{ij}$ depends on the positional embedding strategy with
+[…]
+Here $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{H}$ are pair-wise graph conditioning tokens (see Eq.
+7) and $\boldsymbol{p}_{i}^{\text{aPE}}\in\mathbb{R}^{H}$ and $\boldsymbol{p}_{ij}^{\text{rPE}}\in\mathbb{R}^{H}$ are the absolute and relative PEs described above (see Eqs.
+8 and 9).
+D.1.2 Adaptive Layer Normalization and Adaptive Scale
 In the standard, non-equivariant setting, we can follow the standard approach of other DiT architectures. We define adaptive layer norm as
@@
 The weight matrix $\boldsymbol{W}$ is initialized to all zeros, such that “AdaLN” behaves like identity at initialization.
-“AdaScale” damps all input tokens to zero at initialization such that the whole DiT block behaves like the identity function at initialization.
-B.1.2 Readout
+“AdaScale” damps all input tokens to zero at initialization such that the whole DiTMC block behaves like the identity function at initialization.
+D.1.3 Readout
 Given final tokens $\boldsymbol{h}$ after performing updates via $T$ DiTMC blocks, we use a readout layer to predict the atomic positions of the clean data sample $\boldsymbol{x}_{1}$.
@@
 where $\boldsymbol{W}_{\text{readout}}\in\mathbb{R}^{3\times H}$ is a trainable weight matrix. Thus, we predict a three-dimensional vector per-atom.
-B.2 SO(3) Equivariant DiT
+D.2 SO(3) Equivariant DiTMC
 Following the notation in Ref.
-[60], we denote $\mathrm{SO}(3)$-equivariant tokens as =$\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\mathbb{R}^{(L+1)^{2}\times H}\}$, where $L$ denotes the maximal degree of the spherical harmonics.
+[unkee3x], we denote $\mathrm{SO}(3)$-equivariant tokens as $\mathcal{H}=\{\boldsymbol{h}_{1},\dots,\boldsymbol{h}_{N}\,|\,\boldsymbol{h}_{i}\in\mathbb{R}^{(L+1)^{2}\times H}\}$, where $L$ denotes the maximal degree of the spherical harmonics.
 We denote the features corresponding to the $\ell$-th degree as $\boldsymbol{h}_{i}^{(\ell)}\in\mathbb{R}^{(2\ell+1)\times H}$, where the $(2\ell+1)$ entries corresponds to the orders $m=-\ell,\dots,+\ell$ per degree $\ell$.
 We refer the reader to Ref.
-[60] for an in-depth introduction into equivariant features.
+[unkee3x] for an in-depth introduction into equivariant features.
+For the initial token representation $\boldsymbol{h}_{i}\in\mathbb{R}^{(L+1)^{2}\times H}$ we set $\boldsymbol{h}_{i}^{(0)}=\boldsymbol{e}(z_{i})$, where $\boldsymbol{e}(z_{i})\in\mathbb{R}^{1\times H}$ denotes a learnable embedding based on the atomic number $z_{i}\in\mathbb{N}_{+}$ for atom $i$ [frank2024fastandstable].
+All higher order features $\boldsymbol{h}_{i}^{(\ell)}$ with degree $\ell>0$ are initialized with zero.
 For all our experiments we use maximal degree of $L=1$.
-B.2.1 Adaptive Layer Normalization and Adaptive Scale
-In the $\mathrm{SO}(3)$-equivariant case, we define adapted version of AdaLN and AdaScale, preserving equivariance. Our equivariant formulation of AdaLN is given as
+D.2.1 Self-Attention Operation
+Our equivariant version of self-attention uses the same transformations for queries, keys and values like the non-equivariant counterpart, as well as the modified similarity kernel (see Appendix subsubsection D.2.1).
+However, to preserve all Euclidean symmetries throughout the network, every token must transform equivariantly.
+One way to achieve this is by separating out the rotational degrees of freedom, encoding them with irreducible representations of the rotation group $\mathrm{SO(3)}$.
+This introduces a “degree-axis” of size $(L+1)^{2}$, which encodes angular components of increasing order.
+The maximum degree $L$ is chosen to ensure high fidelity at a reasonable computational cost.
+For example, setting $L=1$ restricts the representation to scalars and vectors, as used in models like PaiNN [schutt2021equivariant] or TorchMDNet [tholke2022equivariant].
+An $\mathrm{SO(3)}$-equivariant formulation of self-attention is then given as
+$\text{ATT}_{\text{SO(3)}}(\mathcal{H})_{i}=\frac{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})\cdot(\hat{\boldsymbol{u}}_{ij}\otimes\boldsymbol{v}_{j})}{\sum_{j=1}^{N}\text{sim}(\boldsymbol{q}_{i},\boldsymbol{k}_{j},\boldsymbol{u}_{ij})}\,$, (A31)
+where equivariant queries, keys and values can be calculated similarly to Eq.
+A23 and $\otimes$’ denotes a Clebsch-Gordan (CG) tensor product contraction [unkee3x].
+The dot-product in the similarity measure is taken along both feature and degree axes, such that the overall update preserves equivariance (see Appendix Appendix F for details).
+Tokens and scaling vectors are calculated as
+$\hskip 15.0pt\boldsymbol{u}_{ij}=\boldsymbol{\phi}(r_{ij})\odot\boldsymbol{c}_{ij}^{\mathcal{G}}\,,\hskip 15.0pt\hat{\boldsymbol{u}}_{ij}=\boldsymbol{p}_{ij}^{\text{PE(3)}}\odot\boldsymbol{c}_{ij}^{\mathcal{G}}\,$, (A32)
+where $\boldsymbol{\phi}(r_{ij})\in\mathbb{R}^{(L+1)^{2}\times H}$ is a radial filter, and the element-wise products with the pair-wise conditioning tokens $\boldsymbol{c}_{ij}^{\mathcal{G}}\in\mathbb{R}^{1\times H}$ are broadcast along the degree axis.
+Importantly, the $2\ell+1$ subcomponents of the radial filter for degree $\ell$ are obtained by repeating per-degree filter functions $\boldsymbol{\phi}_{\ell}(r_{ij})\in\mathbb{R}^{1\times H}$ along the degree axis to preserve equivariance (see also Eq.
+10).
+Since, standard MLPs do not preserve SO(3)-equivariance, we use equivariant MLPs for the node-wise refinement after the self-attention calculation via an equivariant formulation for dense layers and so-called gated non-linearities (see e.g.
+Ref.
+[unkee3x] for more details).
+D.2.2 Adaptive Layer Normalization and Adaptive Scale
+In the $\mathrm{SO}(3)$-equivariant case, we define an adapted version of AdaLN and AdaScale, preserving equivariance. Our equivariant formulation of AdaLN is given as
 $\text{EquivAdaLN}(\boldsymbol{h},\boldsymbol{\alpha},\boldsymbol{\beta})=\begin{cases}\text{LN}(\boldsymbol{h}^{(\ell)})\odot(1+\boldsymbol{\alpha}^{(0)})+\boldsymbol{\beta}\hskip 100.00015pt&\text{for }\ell=0,\\ \text{EquivLN}(\boldsymbol{h}^{(\ell)})\odot(1+\boldsymbol{\alpha}^{(\ell)})&\text{for }\ell>0,\end{cases}$ (A33)
 where EquivLN is the equivariant formulation of layer normalization following Ref.
-[36] without trainable per-degree scales and LN is standard layer normalization without trainable scale and bias.
+[liao2024equiformerv] without trainable per-degree scales and LN is standard layer normalization without trainable scale and bias.
 Scaling vectors $\boldsymbol{\alpha}^{(\ell)}\in\mathbb{R}^{1\times H}$ are defined per degree $\ell$, such that input scaling vectors are tensors $\boldsymbol{\alpha}\in\mathbb{R}^{(L+1)\times H}$.
@@
 Thus, also the $\mathrm{SO}(3)$-equivariant DiTMC block behaves like the identity function at initialization.
-B.2.2 Equivariant positional embeddings
-Given a set of transformations that act on a vector space $\mathbb{A}$ as $S_{g}:\mathbb{A}\mapsto\mathbb{A}$ to which we associate an abstract group $G$, a function $f:\mathbb{A}\mapsto\mathbb{B}$ is said to be equivariant w.r.t.
-$G$ if
-$f(S_{g}x)=T_{g}f(x)\,,$ (A36)
-where $T_{g}:\mathbb{B}\mapsto\mathbb{B}$ is an equivalent transformation on the output space.
-Thus, in order to say that $f$ is equivariant, it must hold that under transformation of the input, the output transforms “in the same way”.
-Let us now recall our definition for the equivariant positional embeddings for a single degree $\ell$
-$\boldsymbol{p}_{ij}^{\text{PE(3)},(\ell)}(\vec{r}_{ij})=\boldsymbol{\phi}_{\ell}(||\vec{r}_{ij}||)\,\odot\boldsymbol{Y}_{\ell}(\hat{r}_{ij})\,,$ (A37)
-where $\boldsymbol{\phi}_{\ell}:\mathbb{R}\mapsto\mathbb{R}^{1\times H}$ is a radial filter function, $\hat{r}=\vec{r}/r$, and $\boldsymbol{Y}_{\ell}\in\mathbb{R}^{(2\ell+1)\times 1}$ are spherical harmonics of degree $\ell=0\dots L$.
-The element-wise multiplication $\odot$’ between radial filters and spherical harmonics is understood to be “broadcasting” along axes with size $1$, such that $(\boldsymbol{\phi}_{\ell}\,\odot\boldsymbol{Y}_{\ell})\in\mathbb{R}^{(2\ell+1)\times H}$.
-We have also made the dependence of PE(3) on the pairwise displacement vector $\vec{r}_{ij}=\vec{r}_{i}-\vec{r}_{j}$ explicit.
-Lets not consider a single feature channel $c$ in PE(3), which is given as
-$\boldsymbol{p}_{ijc}^{\text{PE(3)},(\ell)}(\vec{r}_{ij})=\boldsymbol{\phi}_{\ell c}(||\vec{r}_{ij}||)\,\odot\boldsymbol{Y}_{\ell}(\hat{r}_{ij})\,.$ (A38)
-Rotating the input positions in Eq. A38 leads to
-$\boldsymbol{p}_{ijc}^{\text{PE(3)},(\ell)}(\boldsymbol{R}\vec{r}_{ij})$ $=\boldsymbol{\phi}_{\ell c}(||\boldsymbol{R}\vec{r}_{ij}||)\,\odot\boldsymbol{Y}_{\ell}(\boldsymbol{R}\hat{r}_{ij})$ (A39)
-$=\boldsymbol{\phi}_{\ell c}(||\vec{r}_{ij}||)\,\odot\boldsymbol{D}^{(\ell)}(\boldsymbol{R})\boldsymbol{Y}_{\ell}(\hat{r}_{ij}),$ (A40)
-$=\boldsymbol{D}^{(\ell)}(\boldsymbol{R})\boldsymbol{p}_{ijc}^{\text{PE(3)},(\ell)}(\vec{r}_{ij})$ (A41)
-where $\boldsymbol{D}^{(\ell)}\in\mathbb{R}^{(2\ell+1)\times(2\ell+1)}$ are the Wigner-D matrices for degree $\ell$ and $\boldsymbol{R}\in\mathbb{R}^{3\times 3}$ is a rotation matrix.
-According to Eq.
-A36 and Eq.
-A41, each channel transforms equivariant and thus $\boldsymbol{p}_{ij}^{\text{PE(3)},(\ell)}(\vec{r}_{ij})$ is also equivariant.
-The concatenation of different degrees $\ell$ up to some maximal degree $L$ as given in the main body of the text
-$\boldsymbol{p}_{ij}^{\text{PE(3)}}(\vec{r}_{ij})=\bigoplus_{\ell=0}^{L}\boldsymbol{\phi}_{\ell}(r_{ij})\,\odot\boldsymbol{Y}_{\ell}(\hat{r}_{ij})\,,$ (A42)
-transforms under rotation as
-$\boldsymbol{p}_{ij}^{\text{PE(3)}}(\boldsymbol{R}\vec{r}_{ij})=\boldsymbol{D}(\boldsymbol{R})\,\boldsymbol{p}_{ij}^{\text{PE(3)}}(\vec{r}_{ij})$ (A43)
-with $\boldsymbol{D}(\boldsymbol{R})=\bigoplus_{\ell=0}^{L}\boldsymbol{D}^{(\ell)}(\boldsymbol{R})\in\mathbb{R}^{(L+1)^{2}\times(L+1)^{2}}$ being a block-diagonal matrix with the Wigner-D matrices of degree $\boldsymbol{D}^{(\ell)}(\boldsymbol{R})\in\mathbb{R}^{(2\ell+1)\times(2\ell+1)}$ along the diagonal.
-Therefore, according to Eq.
-A36 the proposed positional embeddings $\boldsymbol{p}_{ij}^{\text{PE(3)}}$ are $\mathrm{SO}(3)$-equivariant.
-B.3 Invariance of the Dot-Product
-In the self-attention update, the dot-product between query and key is computed as along the degree and the feature axis. Under rotation, the equivariant features behave as
-$\boldsymbol{h}(\boldsymbol{R}\vec{r})=\boldsymbol{D}(\boldsymbol{R})\boldsymbol{h}(\vec{r}),$ (A44)
-where $\boldsymbol{D}(\boldsymbol{R})$ is the concatenation of Wigner-D matrices from above.
-The inner product along the degree axis for two features $\boldsymbol{h}$ and $\boldsymbol{g}$ behaves under rotation as
-$\boldsymbol{g}(\boldsymbol{R}\vec{r})^{T}\cdot\boldsymbol{h}(\boldsymbol{R}\vec{r})=\boldsymbol{g}(\vec{r})^{T}\underbrace{\boldsymbol{D}(\boldsymbol{R})^{T}\boldsymbol{D}(\boldsymbol{R})}_{=\text{Id}}\boldsymbol{h}(\vec{r})=\boldsymbol{g}(\vec{r})^{T}\cdot\boldsymbol{h}(\vec{r}),$ (A45)
-where we made use of the fact that the Wigner-D matrices are orthogonal matrices.
-Thus, the dot-product along the degree-axis is invariant and therefore taking the dot-product along the degree and then along the feature axis is also invariant.
-B.3.1 Readout
+D.2.3 Readout
 Given final equivariant features $\boldsymbol{h}_{i}\in\mathbb{R}^{(L+1)^{2}\times H}$ we use a readout layer to predict the atomic positions of the clean data sample $\boldsymbol{x}_{1}$.
@@
 As $\boldsymbol{h}_{i}$ are rotationally equivariant so is $\hat{\boldsymbol{y}}_{i}\in\mathbb{R}^{3}$.
-B.4 Equivariant MLP
-Standard MLPs do not preserve the equivariance of the tokens.
-However, it is possible to define an equivariant formulation for dense layers and so-called gated non-linearities which preserve equivariance.
-We use them to build equivariant MLPs for the node-wise refinement after the self-attention calculation.
-See e.g.
-Ref.
-[60] for more details.
-Appendix C Sampling
-For sampling, we use a simple Euler scheme with 50 steps to sample from the associated ordinary differential equation (ODE) as described in Algorithm 2.
-Since during training we predict the clean sample $\boldsymbol{x}_{1}$, we re-parametrize the velocity required for the integration as
-$\boldsymbol{v}^{\theta}_{t}(\boldsymbol{x}_{t},t,\mathcal{G})=\frac{\boldsymbol{x}^{\theta}_{1}(\boldsymbol{x}_{t},t,\mathcal{G})-\boldsymbol{x}_{t}}{1-t},$ (A47)
-where $\boldsymbol{x}^{\theta}_{1}(\boldsymbol{x}_{t},t,\mathcal{G})$ is the original output of DiTMC.
-To ensure $\mathrm{SE}(3)$ invariance of the probability path from an (approximately) $\mathrm{SO}(3)$-equivariant velocity predictor, we center the prior $\boldsymbol{x}_{0}\sim p_{0}(\boldsymbol{x})$ as well as the prediction of DiTMC in each ODE step.
-Algorithm 2 ODE Sampling
-1:Model $\boldsymbol{x}_{1}^{\theta}$, Graph $\mathcal{G}$, steps $N>0$
-2:
-3:$t_{n}\leftarrow n/N$ for $n\in\{0,\ldots,N\}$
-4:$\boldsymbol{x}_{0}\sim p_{\text{prior}}(\boldsymbol{x})$triangleright Sample prior.
-5:$\boldsymbol{x}_{0}\leftarrow\mathrm{Center}(\boldsymbol{x}_{0})$
-6:for $n\leftarrow 0$ to $N-1$ do
-7:     $\Delta t\leftarrow t_{n+1}-t_{n}$triangleright Compute step size.
-8:     $\hat{\boldsymbol{x}}_{1}\leftarrow\boldsymbol{x}_{1}^{\theta}(\boldsymbol{x}_{t_{n}},t_{n},\mathcal{G})$
-9:     $\hat{\boldsymbol{x}}_{1}\leftarrow\mathrm{Center}(\hat{\boldsymbol{x}}_{1})$
-10:     $\boldsymbol{v}\leftarrow(\hat{\boldsymbol{x}}_{1}-\boldsymbol{x}_{t_{n}})/(1-t_{n})$
-11:     $\boldsymbol{x}_{t_{n+1}}\leftarrow\boldsymbol{x}_{t_{n}}+\Delta t\cdot\boldsymbol{v}$triangleright Euler step.
-12:end for
-13:
-14:return $\boldsymbol{x}_{1}$
-Appendix D Implementation details
-D.1 Data Preprocessing
+D.3 Conditioning GNN
+Our conditioning GNN directly operates on the molecular graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ to obtain graph-based information for atom-wise conditioning.
+First, we processes the molecular graph $\mathcal{G}$ to derive node and edge input features, initially represented as one-hot vectors (a full list of features is provided in Tab.
+A8).
+These features then are projected into a shared latent space using two-layer multilayer perceptrons (MLPs).
+Finally, we employ a GNN architecture inspired by the processor described in the MeshGraphNet (MGN) framework [pfaff2020learning], which refines the features via message passing.
+The conditioning GNN maintains and updates both node and edge representations across multiple layers.
+Each message passing block consists of two main steps: first, the edge representations are updated based on the current edge representations and the representations of the connected nodes:
+$\boldsymbol{e}_{ij}^{\prime}\leftarrow f_{e}(\boldsymbol{e}_{ij},\boldsymbol{v}_{i},\boldsymbol{v}_{j})$ (A37)
+where $\mathbf{e}_{ij}$ and $\mathbf{v}_{i}$ denote the input edge and node representations, and $\mathbf{e}_{ij}^{\prime}$ are the updated edge representations.
+The learnable function $f_{e}$ is implemented as a two-layer MLP.
+Next, node representations $\mathbf{v}_{i}$ are updated to $\mathbf{v}_{i}^{\prime}$ using aggregated messages from neighboring edges:
+$\boldsymbol{v}_{i}^{\prime}\leftarrow f_{v}\left(\boldsymbol{v}_{i},\sum_{j}\boldsymbol{e}_{ij}^{\prime}\right)$ (A38)
+where $f_{v}$ is also a two-layer MLP, and the summation is over all edges ending at node $i$.
+The final output of the described conditioning GNN is a set of node embeddings per atom, and a set of edge embeddings per bond.
+We use the final node embeddings as atom-wise graph conditioning tokens in DiTMC as detailed in Sec.
+4.1.
+We want to highlight that one can also use the final edge embedding as pair-wise graph conditioning tokens in DiTMC.
+We discuss this further in Appendix I.2.
+Appendix E Equivariance Proof for Euclidean Positional Embeddings
+Given a set of transformations that act on a vector space $\mathbb{A}$ as $S_{g}:\mathbb{A}\mapsto\mathbb{A}$ to which we associate an abstract group $G$, a function $f:\mathbb{A}\mapsto\mathbb{B}$ is said to be equivariant w.r.t.
+$G$ if
+$f(S_{g}x)=T_{g}f(x)\,$, (A39)
+where $T_{g}:\mathbb{B}\mapsto\mathbb{B}$ is an equivalent transformation on the output space.
+Thus, in order to say that $f$ is equivariant, it must hold that under transformation of the input, the output transforms “in the same way”.
+Let us now recall our definition for the equivariant positional embeddings for a single degree $\ell$
+[…]
+where $\boldsymbol{\phi}_{\ell}:\mathbb{R}\mapsto\mathbb{R}^{1\times H}$ is a radial filter function, $\hat{r}=\vec{r}/r$, and $\boldsymbol{Y}_{\ell}\in\mathbb{R}^{(2\ell+1)\times 1}$ are spherical harmonics of degree $\ell=0\dots L$.
+The element-wise multiplication $\odot$’ between radial filters and spherical harmonics is understood to be “broadcasting” along axes with size $1$, such that $(\boldsymbol{\phi}_{\ell}\,\odot\boldsymbol{Y}_{\ell})\in\mathbb{R}^{(2\ell+1)\times H}$.
+We have also made the dependence of PE(3) on the pairwise displacement vector $\vec{r}_{ij}=\vec{r}_{i}-\vec{r}_{j}$ explicit.
+Lets not consider a single feature channel $c$ in PE(3), which is given as
+[…]
+Rotating the input positions in Eq. A41 leads to
+[…]
+[…]
+$=\boldsymbol{D}^{(\ell)}(\boldsymbol{R})\boldsymbol{p}_{ijc}^{\text{PE(3)},(\ell)}(\vec{r}_{ij})$ (A44)
+where $\boldsymbol{D}^{(\ell)}\in\mathbb{R}^{(2\ell+1)\times(2\ell+1)}$ are the Wigner-D matrices for degree $\ell$ and $\boldsymbol{R}\in\mathbb{R}^{3\times 3}$ is a rotation matrix.
+According to Eq.
+A39 and Eq.
+A44, each channel transforms equivariant and thus $\boldsymbol{p}_{ij}^{\text{PE(3)},(\ell)}(\vec{r}_{ij})$ is also equivariant.
+The concatenation of different degrees $\ell$ up to some maximal degree $L$ as given in the main body of the text
+$\boldsymbol{p}_{ij}^{\text{PE(3)}}(\vec{r}_{ij})=\bigoplus_{\ell=0}^{L}\boldsymbol{\phi}_{\ell}(r_{ij})\,\odot\boldsymbol{Y}_{\ell}(\hat{r}_{ij})\,$, (A45)
+transforms under rotation as
+$\boldsymbol{p}_{ij}^{\text{PE(3)}}(\boldsymbol{R}\vec{r}_{ij})=\boldsymbol{D}(\boldsymbol{R})\,\boldsymbol{p}_{ij}^{\text{PE(3)}}(\vec{r}_{ij})$ (A46)
+with $\boldsymbol{D}(\boldsymbol{R})=\bigoplus_{\ell=0}^{L}\boldsymbol{D}^{(\ell)}(\boldsymbol{R})\in\mathbb{R}^{(L+1)^{2}\times(L+1)^{2}}$ being a block-diagonal matrix with the Wigner-D matrices of degree $\boldsymbol{D}^{(\ell)}(\boldsymbol{R})\in\mathbb{R}^{(2\ell+1)\times(2\ell+1)}$ along the diagonal.
+Therefore, according to Eq.
+A39 the proposed positional embeddings $\boldsymbol{p}_{ij}^{\text{PE(3)}}$ are $\mathrm{SO}(3)$-equivariant.
+Appendix F Invariance Proof for the Dot-Product
+In the self-attention update, the dot-product between query and key is computed as along the degree and the feature axis. Under rotation, the equivariant features behave as
+$\boldsymbol{h}(\boldsymbol{R}\vec{r})=\boldsymbol{D}(\boldsymbol{R})\boldsymbol{h}(\vec{r}),$ (A47)
+where $\boldsymbol{D}(\boldsymbol{R})$ is the concatenation of Wigner-D matrices from above.
+The inner product along the degree axis for two features $\boldsymbol{h}$ and $\boldsymbol{g}$ behaves under rotation as
+$\boldsymbol{g}(\boldsymbol{R}\vec{r})^{T}\cdot\boldsymbol{h}(\boldsymbol{R}\vec{r})=\boldsymbol{g}(\vec{r})^{T}\underbrace{\boldsymbol{D}(\boldsymbol{R})^{T}\boldsymbol{D}(\boldsymbol{R})}_{=\text{Id}}\boldsymbol{h}(\vec{r})=\boldsymbol{g}(\vec{r})^{T}\cdot\boldsymbol{h}(\vec{r}),$ (A48)
+where we made use of the fact that the Wigner-D matrices are orthogonal matrices.
+Thus, the dot-product along the degree-axis is invariant and therefore taking the dot-product along the degree and then along the feature axis is also invariant.
+Appendix G Implementation details
+G.1 Data Preprocessing
 For both GEOM-QM9 and GEOM-DRUGS, we use the first 30 conformers for each molecule with the lowest energies, i.e., highest Boltzmann weights.
-We use the train/test/val split from Geomol [65], using the same 1000 molecules for testing.
-D.2 Input Featurization
-Tab. A7 defines the features we use for each atom or bond. Each feature is one-hot encoded before being passed to the network.
+We use the train/test/val split from Geomol [ganea2021geomol], using the same 1000 molecules for testing.
+G.2 Input Featurization
+Tab. A8 defines the features we use for each atom or bond. Each feature is computed using RDKit [landrum2013rdkit] and one-hot encoded before being passed to the network.
 Table A8: Atomic and bond features included in DiT-MC. All features are one-hot encoded.
@@
 Bond type single, double, triple, aromatic
-D.3 Evaluation Metrics
+G.3 Evaluation Metrics
 During evaluation, we follow the same procedure as described in Refs.
-[65, 47, 46, 32].
+[ganea2021geomol, wang2023swallowing, jing2022torsional, hassan2024etflow].
 The root-mean-square deviation (RMSD) metric measures the average distance between atoms of a generated conformer with respect to its reference, while taking into account all possible symmetries.
@@
 The average minimum RMSD (AMR) and coverage (COV) metrics for both recall (R) and precision (P) are defined as follows, where $\delta>0$ is the coverage threshold:
+$\text{COV-R}(C,\hat{C},\delta)$ $:=\frac{1}{K}\left|\left\{k\in\{1,\dots,K\}\mid\exists_{l\in\{1,\dots,L\}}\text{RMSD}(\hat{C}_{l},C_{k})<\delta\right\}\right|$ (A49)
+$\text{COV-P}(C,\hat{C},\delta)$ $:=\frac{1}{L}\left|\left\{l\in\{1,\dots,L\}\mid\exists_{k\in\{1,\dots,K\}}\text{RMSD}(\hat{C}_{l},C_{k})<\delta\right\}\right|$ (A50)
 […]
 […]
+G.4 Chirality Correction
+Given the four 3D coordinates around a chirality center denoted as $\mathbf{p_{1}},\mathbf{p_{2}},\mathbf{p_{3}},\mathbf{p_{4}}\in\mathbb{R}^{3}$ with $\mathbf{p_{i}}=(x_{i},y_{i},z_{i})$ for $i=1,2,3,4$, we can compute the oriented volume $OV$ of the tetrahedron via
 […]
-[…]
-D.4 Chirality Correction
-Given the four 3D coordinates around a chirality center denoted as $\mathbf{p_{1}},\mathbf{p_{2}},\mathbf{p_{3}},\mathbf{p_{4}}\in\mathbb{R}^{3}$ with $\mathbf{p_{i}}=(x_{i},y_{i},z_{i})$ for $i=1,2,3,4$, we can compute the volume $V$ of the tetrahedron as follows
-$V(\mathbf{p_{1}},\mathbf{p_{2}},\mathbf{p_{3}},\mathbf{p_{4}})=\det\left(\begin{bmatrix}1&1&1&1\\ x_{1}&x_{2}&x_{3}&x_{4}\\ y_{1}&y_{2}&y_{3}&y_{4}\\ z_{1}&z_{2}&z_{3}&z_{4}\end{bmatrix}\right)=(\mathbf{p_{1}}-\mathbf{p_{4}})\cdot((\mathbf{p_{2}}-\mathbf{p_{4}})\times(\mathbf{p_{3}}-\mathbf{p_{4}}))$ (A52)
-Following GeomMol [65] and ET-Flow [32] we can then compare the orientation of the volume given by $\text{sign}(V)$ with the local chirality label produced by RDKit, which corresponds to a certain orientation as well (CW = +1 and CCW = -1) [69].
+$=\frac{1}{6}\cdot(\mathbf{p_{1}}-\mathbf{p_{4}})\cdot((\mathbf{p_{2}}-\mathbf{p_{4}})\times(\mathbf{p_{3}}-\mathbf{p_{4}}))\,.$
+Following GeomMol [ganea2021geomol] and ET-Flow [hassan2024etflow], we can then compare the orientation of the volume given by $\text{sign}(OV)$ with the local chirality label produced by RDKit, which corresponds to a certain orientation as well (CW = +1 and CCW = -1) [pattanaik2020message].
 If the orientation of the volume differs from the RDKit label, we correct the chirality of the conformer by reflecting its positions against the $z$-axis.
-Appendix E Additional Experimental Results
-E.1 Scaling Results for PE(3)
-Figure A4: (A) Number of gradient steps vs.
-wall time for PE(3) and aPE on the GEOM-DRUGS dataset measured on Nvidia H100 GPU with batch size 128.
-(B) Coverage precision and recall as function of threshold $\delta$ and (C) mean AMR metrics for base (“B”) and large (“L”) model of DiTMC+PE(3).
-We investigate the scaling behavior of our equivariant models in Fig.
-A4.
-Tab.
-A10 shows a comparison of our base and large models for GEOM drugs.
-Similar to its non-equivariant counterpart, our scaled PE(3)-L model yields improved results for precision metrics, while we observe only minor improvements for recall.
-Tab.
-A11, Tab.
-A9 and Tab.
-A13 further demonstrate improvements for ensemble property prediction and out-of-distribution generalization using the larger model.
-E.2 Conditioning ablation
-To evaluate the effectiveness of various graph conditioning strategies in DiTMC , we compare the performance of three conditioning methods – “node only”, “node & bond-pair”, and “node & all-pair” – against a baseline model without any conditioning.
-As reported in Tab.
-A8, all conditioning variants significantly reduce the Average Minimum RMSD (AMR) for both recall (AMR-R) and precision (AMR-P) across the aPE and PE(3) settings, compared to the unconditioned baseline.
-Notably, the “node & all-pair” strategy achieves the best overall performance, with the lowest AMR values.
-These results highlight the strength of the all-pair conditioning strategy, which leverages graph geodesics to incorporate information from all atom pairs, rather than restricting conditioning to directly connected nodes or bonded pairs.
-This comprehensive approach captures more global structural information, thereby improving both precision and recall.
-See Appendix K for a more in-depth analysis.
-Table A8: Ablation of conditioning strategies in DiTMC (in brackets) with absolute PE and PE(3) on GEOM-QM9.
--R indicates Recall, -P indicates Precision.
-Best results in bold.
-Our results are averaged over three random seeds with standard deviation reported below.
-Method COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
-Mean Median Mean Median Mean Median Mean Median
-DiTMC+aPE-B (no conditioning) 68.6 $\pm$1.0 91.7 $\pm$2.1 0.405 $\pm$0.005 0.325 $\pm$0.004 36.8 $\pm$0.5 36.4 $\pm$2.2 0.729 $\pm$0.006 0.703 $\pm$0.007
-DiTMC+PE(3)-B (no conditioning) 71.5 $\pm$0.5 99.0 $\pm$1.4 0.358 $\pm$0.003 0.266 $\pm$0.005 40.4 $\pm$0.2 42.2 $\pm$0.4 0.683 $\pm$0.003 0.653 $\pm$0.005
-DiTMC+aPE-B (node only) 96.3 $\pm$0.0 100.0 $\pm$0.0 0.079 $\pm$0.000 0.037 $\pm$0.000 93.2 $\pm$0.2 100.0 $\pm$0.0 0.112 $\pm$0.001 0.051 $\pm$0.000
-DiTMC+PE(3)-B (node only) 95.5 $\pm$0.4 100.0 $\pm$0.0 0.074 $\pm$0.003 0.026 $\pm$0.001 91.1 $\pm$0.4 100.0 $\pm$0.0 0.114 $\pm$0.003 0.041 $\pm$0.001
-DiTMC+aPE-B (node & bond-pair) 96.5 $\pm$0.1 100.0 $\pm$0.0 0.077 $\pm$0.001 0.035 $\pm$0.001 95.3 $\pm$0.2 100.0 $\pm$0.0 0.092 $\pm$0.001 0.046 $\pm$0.002
-DiTMC+PE(3)-B (node & bond-pair) 96.1 $\pm$0.3 100.0 $\pm$0.0 0.068 $\pm$0.001 0.022 $\pm$0.001 93.6 $\pm$0.2 100.0 $\pm$0.0 0.091 $\pm$0.003 0.035 $\pm$0.001
-DiTMC+aPE-B (node & all-pair) 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
-DiTMC+PE(3)-B (node & all-pair) 95.7 $\pm$0.3 100.0 $\pm$0.0 0.068 $\pm$0.002 0.021 $\pm$0.001 93.4 $\pm$0.2 100.0 $\pm$0.0 0.089 $\pm$0.002 0.032 $\pm$0.001
-E.3 Ensemble Properties
-While RMSD-based metrics offer a geometric perspective on the quality of the generated conformers, they do not assess the chemical fidelity.
-To address this, we evaluate the median averaged errors of different ensemble properties between the generated and ground truth conformers.
-We adopt the property prediction task setup from MCF [47] and ET-Flow [32], where we draw a subset of 100 randomly sampled molecules from the test set of GEOM-DRUGS and generate $\min(2K,32)$ conformers for a molecule with $K$ ground truth conformers.
-Afterwards we relax the conformers using GFN2-xTB [70] and compare the Boltzmann-weighted properties of the generated and ground truth ensembles.
-More specifically, we employ xTB [70] to calculate the energy $E$, the dipole moment $\mu$, the HOMO-LUMO gap $\Delta\epsilon$ and the minimum energy $E_{\text{min}}$.
-We repeat this procedure for three subsets each sampled with a different random seed and report the averaged median absolute error and standard deviation of the different ensemble properties in Table A11.
+G.5 Pareto Front
+For all models, we generate conformers using 5, 10, 20 and 50 sampling steps on a single A100 GPU with a batch size of 128, following Refs.
+[hassan2024etflow, wang2023swallowing].
+The wall-clock time per generated sample is obtained by measuring the average time per batch and dividing by the batch size.
+As done in the original paper, we adopt DDIM sampling for MCF-S, MCF-B and MCF-L.
+G.6 Ensemble Properties
+We adopt the property prediction task setup from MCF [wang2023swallowing] and ET-Flow [hassan2024etflow], where we draw a subset of 100 randomly sampled molecules from the test set of GEOM-DRUGS and generate $\min(2K,32)$ conformers for a molecule with $K$ ground truth conformers.
+Afterwards we relax the conformers using GFN2-xTB [bannwarth2019xtb] and compare the Boltzmann-weighted properties of the generated and ground truth ensembles.
+More specifically, we employ xTB [bannwarth2019xtb] to calculate the energy $E$, the dipole moment $\mu$, the HOMO–LUMO gap $\Delta\epsilon$ and the minimum energy $E_{\text{min}}$.
+We repeat this procedure for three subsets each sampled with a different random seed and report the averaged median absolute error and standard deviation of the different ensemble properties.
+Appendix H Additional Experimental Results
+H.1 Results on GEOM-QM9
+We report the results on GEOM-QM9 including standard deviations for all DiTMC models in Tab. A9.
 Table A9: Results on GEOM-QM9 for different generative models (number of parameters in parentheses).
 -R indicates Recall, -P indicates Precision.
-Best results in bold, second best underlined.
-Our models are marked with an asterisk “*”.
+Best results in bold, second best underlined; our models are marked with an asterisk “*”.
 Our results are averaged over three random seeds with standard deviation reported below.
@@
 *DiTMC+PE(3)-B (8.6M) 95.7 $\pm$0.3 100.0 $\pm$0.0 0.068 $\pm$0.002 0.021 $\pm$0.001 93.4 $\pm$0.2 100.0 $\pm$0.0 0.089 $\pm$0.002 0.032 $\pm$0.001
+H.2 Results on GEOM-DRUGS
+We report the results on GEOM-DRUGS including standard deviations for all DiTMC models in Tab. A10. The median absolute error of ensemble properties on GEOM-DRUGS is shown in Tab. A11.
+Additional results for the coverage vs. RMSD threshold analysis, including results for DiTMC+rPE and DiTMC+PE(3), are provided in Fig. A8, Fig. A9, and Fig. A10.
+Additional results for the Pareto front analysis, including results for DiTMC+rPE and DiTMC+PE(3), are provided in Fig. A11, Fig. A12, and Fig. A13.
 Table A10: Results on GEOM-DRUGS for different generative models (number of parameters in parentheses).
 -R indicates Recall, -P indicates Precision.
-Best results in bold, second best underlined.
-Our models are marked with an asterisk “*”.
+Best results in bold, second best underlined; our models are marked with an asterisk “*”.
 Our results are averaged over three random seeds with standard deviation reported below.
@@
 Tors. Diff. (1.6M) 72.7 80.0 0.582 0.565 55.2 56.9 0.778 0.729
+MCF-S (13M) 79.4 87.5 0.512 0.492 57.4 57.6 0.761 0.715
+MCF-B (64M) 84.0 91.5 0.427 0.402 64.0 66.2 0.667 0.605
 MCF-L (242M) 84.7 92.2 0.390 0.247 66.8 71.3 0.618 0.530
@@
 *DiTMC+aPE-B (9.5M) 79.9 $\pm$0.1 85.4 $\pm$0.3 0.434 $\pm$0.002 0.389 $\pm$0.002 76.5 $\pm$0.1 83.6 $\pm$0.3 0.500 $\pm$0.002 0.423 $\pm$0.004
-*DiTMC+aPE-L (31.1M) 79.2 $\pm$0.1 84.4 $\pm$0.2 0.432 $\pm$0.003 0.386 $\pm$0.003 77.8 $\pm$0.1 85.7 $\pm$0.5 0.470 $\pm$0.001 0.387 $\pm$0.003
+*DiTMC+rPE-B (9.6M) 79.3 $\pm$0.1 84.6 $\pm$0.2 0.444 $\pm$0.002 0.400 $\pm$0.002 77.2 $\pm$0.1 84.6 $\pm$0.2 0.492 $\pm$0.001 0.414 $\pm$0.002
 *DiTMC+PE(3)-B (8.6M) 80.8 $\pm$0.1 85.6 $\pm$0.5 0.427 $\pm$0.001 0.396 $\pm$0.001 75.3 $\pm$0.1 82.0 $\pm$0.2 0.515 $\pm$0.000 0.437 $\pm$0.003
-*DiTMC+PE(3)-L (28.2M) 80.8 $\pm$0.3 85.6 $\pm$0.1 0.415 $\pm$0.003 0.376 $\pm$0.001 76.4 $\pm$0.2 82.6 $\pm$0.3 0.491 $\pm$0.002 0.414 $\pm$0.004
+*DiTMC+aPE-L (28.2M) 79.2 $\pm$0.1 84.4 $\pm$0.2 0.432 $\pm$0.003 0.386 $\pm$0.003 77.8 $\pm$0.1 85.7 $\pm$0.5 0.470 $\pm$0.001 0.387 $\pm$0.003
+*DiTMC+rPE-L (28.3M) 78.7 $\pm$0.1 84.1 $\pm$0.4 0.438 $\pm$0.002 0.388 $\pm$0.005 78.1 $\pm$0.1 86.4 $\pm$0.3 0.466 $\pm$0.001 0.381 $\pm$0.003
+*DiTMC+PE(3)-L (31.1M) 80.8 $\pm$0.3 85.6 $\pm$0.1 0.415 $\pm$0.003 0.376 $\pm$0.001 76.4 $\pm$0.2 82.6 $\pm$0.3 0.491 $\pm$0.002 0.414 $\pm$0.004
 Table A11: Median absolute error of ensemble properties between generated and reference conformers.
-Best results in bold, second best underlined.
-Our models are marked with an asterisk “*”.
-Results for MCF, ET-Flow, and ours are averaged over three random seeds.
+Best results in bold, second best underlined; our models are marked with an asterisk “*”.
+Results for MCF, ET-Flow, and ours are averaged over three random seeds with standard deviation reported below.
+Other works do not report standard deviations.
 Method $E$ [kcal/mol] $\downarrow$ $\mu$ [D] $\downarrow$ $\Delta\epsilon$ [kcal/mol] $\downarrow$ $E_{\text{min}}$ [kcal/mol] $\downarrow$
-OMEGA 0.68 0.66 0.68 0.69
-GeoDiff 0.31 0.35 0.89 0.39
-GeoMol 0.42 0.34 0.59 0.40
-Torsional Diff. 0.22 0.35 0.54 0.13
-MCF 0.68 $\pm$0.06 0.28 $\pm$0.05 0.63 $\pm$0.05 0.04 $\pm$0.00
-ET-Flow 0.18 $\pm$0.01 0.18 $\pm$0.01 0.35 $\pm$0.06 0.02 $\pm$0.00
-*DiTMC+aPE-B 0.17 $\pm$0.00 0.16 $\pm$0.01 0.27 $\pm$0.01 0.01 $\pm$0.00
-*DiTMC+aPE-L 0.16 $\pm$0.02 0.14 $\pm$0.03 0.27 $\pm$0.01 0.01 $\pm$0.00
-*DiTMC+PE(3)-B 0.18 $\pm$0.01 0.18 $\pm$0.01 0.27 $\pm$0.03 0.02 $\pm$0.00
-*DiTMC+PE(3)-L 0.17 $\pm$0.01 0.14 $\pm$0.01 0.25 $\pm$0.01 0.01 $\pm$0.00
-E.4 Generalization Results on GEOM-QM9
-Additionally, we want to assess how well our proposed model architecture generalizes to unseen molecules.
-Here we evaluate our models trained on GEOM-DRUGS on the GEOM-QM9 dataset.
-We report the generalization performance in Table A12.
-Table A12: Out-of-distribution generalization results on GEOM-QM9 for models trained on GEOM-DRUGS.
+GeoDiff (1.6M) 0.31 0.35 0.89 0.39
+GeoMol (0.3M) 0.42 0.34 0.59 0.40
+Torsional Diff. (1.6M) 0.22 0.35 0.54 0.13
+MCF-L (242M) 0.68 $\pm$0.06 0.28 $\pm$0.05 0.63 $\pm$0.05 0.04 $\pm$0.00
+ET-Flow (8.3M) 0.18 $\pm$0.01 0.18 $\pm$0.01 0.35 $\pm$0.06 0.02 $\pm$0.00
+*DiTMC+aPE-B (9.5M) 0.17 $\pm$0.00 0.16 $\pm$0.01 0.27 $\pm$0.01 0.01 $\pm$0.00
+*DiTMC+aPE-L (28.2M) 0.16 $\pm$0.02 0.14 $\pm$0.03 0.27 $\pm$0.01 0.01 $\pm$0.00
+*DiTMC+rPE-B (9.6M) 0.16 $\pm$0.03 0.16 $\pm$0.03 0.29 $\pm$0.02 0.02 $\pm$0.00
+*DiTMC+rPE-L (28.3M) 0.16 $\pm$0.01 0.15 $\pm$0.02 0.28 $\pm$0.06 0.01 $\pm$0.00
+*DiTMC+PE(3)-B (8.6M) 0.18 $\pm$0.01 0.18 $\pm$0.01 0.27 $\pm$0.03 0.02 $\pm$0.00
+*DiTMC+PE(3)-L (31.1M) 0.17 $\pm$0.01 0.14 $\pm$0.01 0.25 $\pm$0.01 0.01 $\pm$0.00
+H.3 Results on GEOM-XL
+We report the results on GEOM-XL including standard deviations for all DiTMC models in Tab. A12.
+Table A12: Out-of-distribution generalization results on GEOM-XL for models trained on GEOM-DRUGS.
 -R indicates Recall, -P indicates Precision.
-Best results in bold, second best underlined.
-Our models are marked with an asterisk “*”.
+Best results in bold, second best underlined; our models are marked with an asterisk “*”.
 Our results are averaged over three random seeds with standard deviation reported below.
 Other works do not report standard deviations.
+Method 75 / 77 mols 102 mols
+AMR-R [Å] $\downarrow$ AMR-P [Å] $\downarrow$ AMR-R [Å] $\downarrow$ AMR-P [Å] $\downarrow$
+Mean Median Mean Median Mean Median Mean Median
+Tor. Diff. (1.6M) 1.93 1.86 2.84 2.71 2.05 1.86 2.94 2.78
+MCF-S (13M) 2.02 1.87 2.90 2.69 2.22 1.97 3.17 2.81
+MCF-B (64M) 1.71 1.61 2.69 2.44 2.01 1.70 3.03 2.64
+MCF-L (242M) 1.64 1.51 2.57 2.26 1.97 1.60 2.94 2.43
+ET-Flow (8.3M) 2.00 1.80 2.96 2.63 2.31 1.93 3.31 2.84
+*DiTMC+aPE-B (9.5M) 1.68 $\pm$0.00 1.47 $\pm$0.02 2.59 $\pm$0.00 2.24 $\pm$0.01 1.96 $\pm$0.00 1.60 $\pm$0.03 2.90 $\pm$0.00 2.48 $\pm$0.03
+*DiTMC+aPE-L (28.2M) 1.56 $\pm$0.01 1.28 $\pm$0.01 2.47 $\pm$0.00 2.14 $\pm$0.01 1.88 $\pm$0.01 1.51 $\pm$0.02 2.81 $\pm$0.00 2.30 $\pm$0.02
+*DiTMC+rPE-B (9.6M) 1.69 $\pm$0.01 1.41 $\pm$0.03 2.52 $\pm$0.00 2.11 $\pm$0.00 1.97 $\pm$0.01 1.61 $\pm$0.01 2.86 $\pm$0.00 2.33 $\pm$0.01
+*DiTMC+rPE-L (28.3M) 1.66 $\pm$0.03 1.37 $\pm$0.01 2.47 $\pm$0.00 2.18 $\pm$0.02 1.96 $\pm$0.02 1.61 $\pm$0.02 2.82 $\pm$0.00 2.42 $\pm$0.02
+*DiTMC+PE(3)-B (8.6M) 1.73 $\pm$0.01 1.55 $\pm$0.01 2.71 $\pm$0.00 2.35 $\pm$0.01 1.98 $\pm$0.01 1.67 $\pm$0.02 3.03 $\pm$0.00 2.60 $\pm$0.01
+*DiTMC+PE(3)-L (31.1M) 1.57 $\pm$0.01 1.46 $\pm$0.01 2.60 $\pm$0.00 2.27 $\pm$0.02 1.85 $\pm$0.02 1.58 $\pm$0.03 2.93 $\pm$0.00 2.53 $\pm$0.03
+Appendix I Additional Ablations
+I.1 Gaussian vs. Harmonic Prior
+As shown in Tab.
+A13, using the harmonic prior improves all metrics slightly for our models on GEOM-QM9.
+Using the harmonic prior however doesn’t seem to be a crucial ingredient for the success of our method, as differences between Gaussian and Harmonic prior appear diminishing.
+As the results in Tab.
+A13 verify, our method can also be used with a simple Gaussian prior effectively.
+For larger molecular graphs the expensive eigendecomposition of the graph Laplacian required for the Harmonic prior could therefore be avoided, which helps scaling our approach more easily.
+Table A13: Ablation of PE strategies and Gaussian (G) and Harmonic (H) prior on GEOM-QM9.
+We report mean coverage (COV) at a threshold of $0.5\mbox{\AA }$, and mean average minimum RMSD (AMR) for Recall -R and Precision -P.
+Best results in bold.
+All results are averaged over three random seeds.
+Method COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
+G H G H G H G H
+DiTMC+aPE-B 96.2 96.1 0.074 0.073 95.2 95.4 0.087 0.085
+DiTMC+rPE-B 96.0 96.3 0.073 0.070 95.2 95.7 0.084 0.080
+DiTMC+PE(3)-B 95.7 95.7 0.069 0.068 93.5 93.4 0.090 0.089
+I.2 Conditioning Strategies on GEOM-QM9
+To evaluate the effectiveness of various graph conditioning strategies in DiTMC , we compare the performance of different conditioning methods against a baseline model without any conditioning.
+In addition to conditioning strategies discussed in Sec.
+4.1, we note that our conditioning GNN also produces edge-level representations, which can be used to define pair-wise graph conditioning tokens as
+$\textbf{bond-pair:}\quad\boldsymbol{c}_{ij}^{\mathcal{G}}=\begin{cases}\text{GNN}_{\text{edge}}(\mathcal{V},\mathcal{E})&\forall(i,j)\in\mathcal{E}\\ \boldsymbol{\bar{c}}^{\mathcal{G}}&\forall(i,j)\not\in\mathcal{E}\,.\end{cases}$ (A54)
+These tokens only capture interactions between bonded atoms, i.e., when $(i,j)\in\mathcal{E}$.
+Conditioning tokens for non-bonded pairs are set to a learnable vector $\boldsymbol{\bar{c}}^{\mathcal{G}}$.
+Self-attention still operates on all atom pairs $(i,j)$, even if they are not connected by a chemical bond.
+Specifically, we ablate the following conditioning strategies:
+  • node only conditioning using only atom-wise graph conditioning tokens $\boldsymbol{c}_{i}^{\mathcal{G}}$ (Eq. 6).
+  • node & bond-pair conditioning using both atom-wise graph conditioning tokens $\boldsymbol{c}_{i}^{\mathcal{G}}$ (Eq.
+6) and pair-wise graph conditioning tokens $\boldsymbol{c}_{ij}^{\mathcal{G}}$ derived from edge-level representations of the conditioning GNN as discussed above (Eq.
+A54).
+  • node & all-pair conditioning using both atom-wise graph conditioning tokens $\boldsymbol{c}_{i}^{\mathcal{G}}$ (Eq.
+6) and pair-wise graph conditioning tokens $\boldsymbol{c}_{ij}^{\mathcal{G}}$ based on geodesic graph distances (Eq.
+7).
+As reported in Tab.
+A14, all our proposed conditioning strategies significantly reduce the average minimum RMSD (AMR) for both recall (AMR-R) and precision (AMR-P) using DiTMC+aPE-B, compared to the unconditioned baseline.
+Notably, the “node & all-pair” strategy achieves the best overall performance, with the lowest AMR values.
+These results highlight the strength of the all-pair conditioning strategy, which leverages graph geodesics to incorporate information from all atom pairs, rather than restricting conditioning to directly connected nodes or bonded pairs.
+This comprehensive approach captures more global structural information, thereby improving both precision and recall.
+See Appendix L for a more in-depth analysis.
+Table A14: Ablation of conditioning strategies using DiTMC+aPE-B on GEOM-QM9.
+-R indicates Recall, -P indicates Precision.
+Best results in bold.
+Our results are averaged over three random seeds with standard deviation reported below.
 Method COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
 Mean Median Mean Median Mean Median Mean Median
-ET-Flow 86.7 100.0 0.218 0.160 68.7 75.3 0.369 0.317
-*DiTMC+aPE-B 84.3 $\pm$0.5 100.0 $\pm$0.0 0.209 $\pm$0.004 0.134 $\pm$0.003 69.4 $\pm$0.7 84.4 $\pm$3.2 0.329 $\pm$0.004 0.250 $\pm$0.008
-*DiTMC+aPE-L 80.6 $\pm$0.5 100.0 $\pm$0.0 0.199 $\pm$0.004 0.115 $\pm$0.004 68.83 $\pm$0.2 86.2 $\pm$2.0 0.329 $\pm$0.002 0.224 $\pm$0.007
-*DiTMC+PE(3)-B 84.8 $\pm$0.3 100.0 $\pm$0.0 0.205 $\pm$0.003 0.130 $\pm$0.005 69.1 $\pm$0.1 81.41 $\pm$2.9 0.329 $\pm$0.004 0.254 $\pm$0.015
-*DiTMC+PE(3)-L 85.1 $\pm$0.2 100.0 $\pm$0.0 0.195 $\pm$0.004 0.115 $\pm$0.003 69.5 $\pm$0.6 82.4 $\pm$2.3 0.319 $\pm$0.05 0.245 $\pm$0.009
-E.5 Generalization Results on GEOM-XL
-We also study how well our proposed model architecture generalizes to unseen molecules with a large number of atoms.
-For this we adopt the GEOM-XL dataset containing a total of 102 molecules with more than 100 atoms presented in [46].
-We report the generalization performance in Table A13.
-Table A13: Out-of-distribution generalization results on GEOM-XL for models trained on GEOM-DRUGS.
--R indicates Recall, -P indicates Precision.
-Best results in bold, second best underlined.
-Our models are marked with an asterisk “*”.
-Our results are averaged over three random seeds with standard deviation reported below.
-Other works do not report standard deviations.
-Method AMR-R [Å] $\downarrow$ AMR-P [Å] $\downarrow$ # mols
-Mean Median Mean Median
-GeoDiff 2.92 2.62 3.35 3.15 -
-GeoMol 2.47 2.39 3.30 3.14 -
-Tor. Diff. 2.05 1.86 2.94 2.78 -
-MCF - S 2.22 1.97 3.17 2.81 102
-MCF - B 2.01 1.70 3.03 2.64 102
-MCF - L 1.97 1.60 2.94 2.43 102
-ET-Flow 2.31 1.93 3.31 2.84 102
-*DiTMC+aPE-B 1.96 $\pm$0.00 1.60 $\pm$0.03 2.90 $\pm$0.00 2.48 $\pm$0.03 102
-*DiTMC+aPE-L 1.88 $\pm$0.01 1.51 $\pm$0.02 2.81 $\pm$0.00 2.30 $\pm$0.02 102
-*DiTMC+PE(3)-B 1.98 $\pm$0.01 1.67 $\pm$0.02 3.03 $\pm$0.00 2.60 $\pm$0.01 102
-*DiTMC+PE(3)-L 1.85 $\pm$0.02 1.58 $\pm$0.03 2.93 $\pm$0.00 2.53 $\pm$0.03 102
-Tor. Diff. 1.93 1.86 2.84 2.71 77
-MCF - S 2.02 1.87 2.9 2.69 77
-MCF - B 1.71 1.61 2.69 2.44 77
-MCF - L 1.64 1.51 2.57 2.26 77
-ET-Flow 2.00 1.80 2.96 2.63 75
-*DiTMC+aPE-B 1.68 $\pm$0.00 1.47 $\pm$0.02 2.59 $\pm$0.00 2.24 $\pm$0.01 77
-*DiTMC+aPE-L 1.56 $\pm$0.01 1.28 $\pm$0.01 2.47 $\pm$0.00 2.14 $\pm$0.01 77
-*DiTMC+PE(3)-B 1.73 $\pm$0.01 1.55 $\pm$0.01 2.71 $\pm$0.00 2.35 $\pm$0.01 77
-*DiTMC+PE(3)-L 1.57 $\pm$0.01 1.46 $\pm$0.01 2.60 $\pm$0.00 2.27 $\pm$0.02 77
-Appendix F Additional Ablations
-F.1 Index Positional Encoding (iPE)
+DiTMC+aPE-B (no conditioning) 68.6 $\pm$1.0 91.7 $\pm$2.1 0.405 $\pm$0.005 0.325 $\pm$0.004 36.8 $\pm$0.5 36.4 $\pm$2.2 0.729 $\pm$0.006 0.703 $\pm$0.007
+DiTMC+aPE-B (node only) 96.3 $\pm$0.0 100.0 $\pm$0.0 0.079 $\pm$0.000 0.037 $\pm$0.000 93.2 $\pm$0.2 100.0 $\pm$0.0 0.112 $\pm$0.001 0.051 $\pm$0.000
+DiTMC+aPE-B (node & bond-pair) 96.5 $\pm$0.1 100.0 $\pm$0.0 0.077 $\pm$0.001 0.035 $\pm$0.001 95.3 $\pm$0.2 100.0 $\pm$0.0 0.092 $\pm$0.001 0.046 $\pm$0.002
+DiTMC+aPE-B (node & all-pair) 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
+I.3 Index Positional Encoding (iPE)
 Tab.
-A14 compares a variant including index positional encoding (iPE) from classic transformer architectures with our base model using aPE on QM9.
+A15 compares a variant including index positional encoding (iPE) from classic transformer architectures with DiTMC+aPE-B on GEOM-QM9.
 Specifically, we use the node index and encode it via sinusoidal encodings into the tokens $\mathcal{H}$ before the first DiTMC block, similar to embedding the absolute positions via aPE.
-Since the graphs are generated from SMILES strings via rdkit and rdkit has to some extend a canonical ordering, this information can be used by the transformer architecture.
+Since the molecular graphs are generated from SMILES strings via RDKit and RDKit has to some extend a canonical ordering, this information can be used by the transformer architecture.
 However, index positional encoding breaks permutation equivariance (as we show in Tab.
-A14).
+A15).
 This might be undesirable as permutation equivariance is one of the fundamental symmetries when learning on graphs.
@@
 We use RDKit (version 2024.9.5) for parsing SMILES strings to graphs.
-Nevertheless, our model using iPE can effectively exploit the information contained in atom indices assigned by rdkit.
-A version of our base aPE model without atom-pair conditioning but iPE achieves comparable performance to our model using geodesic distances as atom-pair conditioning (all-pair conditioning).
-As our all-pair conditioning strategy is similarly or more effective than iPE but additionally preserves permutation equivariance, it should be preferred over iPE and we don’t use iPE in any of our other experiments.
-Table A14: Ablating index positional encoding (iPE) on QM9 for different conditioning strategies (in brackets).
+Nevertheless, our DiTMC+aPE-B using iPE can effectively exploit the information contained in atom indices assigned by RDKit.
+A version of DiTMC+aPE-B without atom-pair conditioning but iPE achieves comparable performance to DiTMC+aPE-B using geodesic distances as atom-pair conditioning (pairwise conditioning).
+As our pairwise conditioning strategy is similarly or more effective than iPE but additionally preserves permutation equivariance, it should be preferred over iPE and we don’t use iPE in any of our other experiments.
+Table A15: Ablating index positional encoding (iPE) on GEOM-QM9 for different conditioning strategies (in brackets).
 To show the effect of atom permutations, we include results with randomly permuted atom indices (perm.).
@@
 DiTMC+aPE+iPE-B (node only), perm. 82.3 $\pm$1.1 100.0 $\pm$0.0 0.229 $\pm$0.008 0.108 $\pm$0.005 60.0 $\pm$0.9 61.8 $\pm$1.4 0.493 $\pm$0.008 0.416 $\pm$0.011
-DiTMC+aPE-B (node & all-pair) 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
-DiTMC+aPE-B (node & all-pair), perm. 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
-F.2 Gaussian vs. Harmonic Prior
-As shown in Tab.
-A15, using the harmonic prior improves all metrics slightly for our models on GEOM-QM9.
-Using the harmonic prior however doesn’t seem to be a crucial ingredient for the success of our method, as differences between Gaussian and Harmonic prior appear diminishing.
-As the results in Tab.
-A15 verify, our method can also be used with a simple Gaussian prior effectively.
-For larger molecular graphs the expensive eigendecomposition required for the Harmonic prior could therefore be avoided, which helps scaling our approach more easily.
-Table A15: Ablation of PE strategies and Gaussian (G) and Harmonic (H) prior.
-We report mean coverage (COV) at a threshold of $0.5\mbox{\AA}$, and mean average minimum RMSD (AMR) for Recall -R and Precision -P.
-Best results in bold.
-All results are averaged over three random seeds.
-PE strategy COV-R [%] $\uparrow$ AMR-R [Å] $\downarrow$ COV-P [%] $\uparrow$ AMR-P [Å] $\downarrow$
-G H G H G H G H
-DiTMC+aPE 96.2 96.1 0.074 0.073 95.2 95.4 0.087 0.085
-DiTMC+rPE 96.0 96.3 0.073 0.070 95.2 95.7 0.084 0.080
-DiTMC+PE(3) 95.7 95.7 0.069 0.068 93.5 93.4 0.090 0.089
-Appendix G Analysis of training loss as a function of latent time
+DiTMC+aPE-B (node & pairwise) 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
+DiTMC+aPE-B (node & pairwise), perm. 96.1 $\pm$0.3 100.0 $\pm$0.0 0.074 $\pm$0.001 0.030 $\pm$0.001 95.4 $\pm$0.1 100.0 $\pm$0.0 0.085 $\pm$0.001 0.037 $\pm$0.000
+Appendix J Analysis of training loss as a function of latent time
 In this section, we provide details for the analysis in Fig.
 2C in the main part of the paper.
-We investigate the effect of the PEs and self-attention formulations on the accuracy of the model.
+We investigate the effect of the positional embeddings and self-attention formulations on the accuracy of the model.
 We therefore take pre-trained models on GEOM-QM9 and compute the training loss (as detailed in Algorithm 1) averaged over 1000 samples drawn randomly from the GEOM-QM9 validation set.
@@
 As detailed in Fig.
-A4, we observe empirically that equivariance leads to a decreased loss close to the data distribution after training.
+A5, we observe empirically that equivariance leads to a decreased loss close to the data distribution after training.
 This explains why our equivariant model more often succeeds to produce samples with increased fidelity, as depicted in figure Fig.
@@
 We further note, that absolute loss values for models trained with all our PE strategies decrease as latent time increases (see Fig.
-A4).
+A5).
 This is expected, as conditional vector fields for each data sample will start to interact more strongly moving away from the data distribution.
-Our weighted loss (see Appendix A) effectively penalizes errors close to the data distribution during training and helps with keeping the error low in this important regime.
-Figure A5: Misclassified exampled for SMILES classification experiment.
-We randomly pick 3 examples, which are misclassified by a classification head without any GNN layers.
-We show that GNN layers are essential for correct classification of isomers.
+Our weighted loss (see Appendix B) effectively penalizes errors close to the data distribution during training and helps with keeping the error low in this important regime.
 Figure A5: Loss as a function of time comparing different PE strategies.
-Results averaged over 1000 samples randomly drawn from the GEOM QM9 validation set.
+Results averaged over 1000 samples randomly drawn from the GEOM-QM9 validation set.
 Left: loss relative to PE(3) as a baseline.
@@
 The loss decreases close to the data distribution for all models.
-Appendix H SMILES Classification Experiment via Conditioning GNN
-Table A16: We measure the discriminative power of our conditioning graph network on a training set of 1000 randomly sampled SMILES strings from the Geom QM9 validation set, as well as a toy dataset of 3 different SMILES strings.
-We investigate the required number of message passing layers, as well as using pre-trained weights from an end-to-end trained model.
-GNN layers Weight init Trainable Accuracy (Geom QM9) Accuracy (Toy Data)
-0 random trainable 0.887 0.333
-1 random trainable 0.999 0.666
-2 random trainable 1.000 1.000
-2 random frozen 0.980 1.000
-2 pre-trained frozen 1.000 1.000
-A critical requirement of our DiTMC approach is the ability to disentangle distinct SMILES representations through conditioning.
-We investigate whether our proposed conditioning strategy is capable of learning the necessary information to distinguish between different SMILES strings for the generation of matching molecular conformers.
+Appendix K SMILES Classification Experiment via Conditioning GNN
+The molecular graphs in our dataset are represented as SMILES strings.
+A critical requirement of DiTMC is the ability to distinguish distinct molecular graphs or SMILES representations through conditioning.
+In this section, we investigate whether our conditioning GNN is capable of learning the necessary information to distinguish between different SMILES strings for the generation of matching molecular conformers.
 As a proxy evaluation task, we assess whether the conditioning network alone can function as a classifier of SMILES strings.
@@
 Each SMILES string becomes a seperate class, so for each class there is exactly one example in the training data.
-The classification task is performed on the graph representations of the molecules, employing the same feature set and GNN architecture utilized in our conditioning graph neural network (Sec.
-4.1, appendix D.2) plus a simple classification head.
-We train models with a batch size of 3 for 5000 epochs (toy dataset) and batch size of 64 over 250 epochs (GEOM-QM9 subset).
+The classification task is performed on the graph representations of the molecules, employing the same feature set and GNN architecture utilized in our conditioning GNN (see Appendix D.3 and Appendix G.2) followed by a simple classification head.
+We train different models with a batch size of 3 for 5000 epochs on the curated toy dataset and batch size of 64 over 250 epochs on the GEOM-QM9 subset.
 We report classification accuracy on the training sets directly to evaluate the model’s discriminative capacity.
@@
 Fig.
-A5 shows that without GNN layers, isomers will be misclassified.
-Appendix I Architecture Details for Conditioning GNN
-To transform SMILES representations to per-token conditioning information, we use a GNN that directly operates on the bond graph induced by the SMILES string.
-Our model processes molecular graphs by first embedding node and edge features, initially represented as one-hot vectors (a full list of features is provided in Tab.
-A7).
-These features are projected into a shared latent space using two-layer multilayer perceptrons (MLPs).
-For message passing, we employ a graph neural network architecture inspired by the processor described in the MeshGraphNet (MGN) framework [54].
-The GNN maintains and updates both node and edge representations across multiple layers.
-Each message passing block consists of two main steps: first, the edge representations are updated based on the current edge representations and the representations of the connected nodes:
-$\boldsymbol{e}_{ij}^{\prime}\leftarrow f_{e}(\boldsymbol{e}_{ij},\boldsymbol{v}_{i},\boldsymbol{v}_{j})$ (A53)
-where $\mathbf{e}_{ij}$ and $\mathbf{v}_{i}$ denote the input edge and node representations, and $\mathbf{e}_{ij}^{\prime}$ are the updated edge representations.
-The learnable function $f_{e}$ is implemented as a two-layer MLP.
-Next, node representations $\mathbf{v}_{i}$ are updated to $\mathbf{v}_{i}^{\prime}$ using aggregated messages from neighboring edges:
-$\boldsymbol{v}_{i}^{\prime}\leftarrow f_{v}\left(\boldsymbol{v}_{i},\sum_{j}\boldsymbol{e}_{ij}^{\prime}\right)$ (A54)
-where $f_{v}$ is also a two-layer MLP, and the summation is over all edges ending at node $i$.
-The final output of the described GNN is a set of node embeddings per atom, and a set of edge embeddings per bond.
-Those serve as atom-wise or bond-pair conditioning inputs to the DiT transformer, as detailed in Sec.
-4.1.
-Appendix J Autoguidance
-We apply autoguidance, a recent technique for enhancing generative models by leveraging the outputs of a weaker model to guide sampling from a stronger one [71].
-Originally proposed in the context of image diffusion, autoguidance works by introducing a degraded model with parameters $\hat{\theta}$ trained on the same data and conditioning as the main model with parameters $\theta$, but with additional constraints such as reduced capacity, increased noise or fewer gradient steps.
-During generation, the predicted vector field $\boldsymbol{v}^{\theta}(\boldsymbol{x},t)$ is guided using the vector field predicted by $\boldsymbol{v}^{\hat{\theta}}(\boldsymbol{x},t)$, where the discrepancy between the two models serves as a corrective signal.
-Intuitively, this nudges the generation process towards higher-quality samples by amplifying the difference between a less accurate and a more accurate model.
-In our molecular conformer generation setup, we choose the parameters after training for 5 epochs without applying exponential moving average (EMA) for $\hat{\theta}$.
-We then use the difference in predicted vector fields from to bias the conformer generation trajectory.
-Following Ref.
-[72], we apply autoguidance only in the guidance interval $t\in[0,t_{max}]$ and additionally ablate $t_{max}$ as the upper bound of the guidance interval.
-It becomes straightforward to adapt Algorithm 2 to reflect these changes during sampling as a drop-in repacement for the predicted vector field.
-The guided velocity (for $t\in[0,t_{max}]$) is computed as $\boldsymbol{v}_{\eta}(\boldsymbol{x},t)=\eta\boldsymbol{v}^{\theta}(\boldsymbol{x},t)+(1-\eta)\boldsymbol{v}^{\hat{\theta}}(\boldsymbol{x},t)$, where $\eta$ is the guidance strength.
-Fig.
-A7 shows that autoguidance yields enhanced fidelity of generated conformations, as indicated by improvements in precision metrics, but might negatively impact diversity (recall).
-Increasing guidance strength improves precision (COV-P, AMR-P) but may reduce recall (COV-R, AMR-R).
-A similar trend can be observed for the guidance interval, where increased $t_{max}$ leads to stronger effects of autoguidance.
-This effectively establishes a trade-off between precision and recall that can be tuned post-hoc after training with minimal effort.
-Figure A7: Autoguidance performance on GEOM-QM9 using our DiTMC+aPE-B model.
-We evaluate the effect of autoguidance on conformer generation quality across various guidance intervals ($t_{max}$) on the x-axes, and guidance strength ($\eta$) as different colors.
-Dashed black lines indicate performance without guidance.
-Appendix K Analysis of Sampling Trajectories
-Figure A8: Comparison of molecular generation with node-only versus node- and all-pair conditioning on GEOM-QM9.
+A6 shows that without GNN layers, isomers will be misclassified.
+Table A16: We measure the discriminative power of our conditioning graph network on a training set of 1000 randomly sampled SMILES strings from the GEOM-QM9 validation set, as well as a toy dataset of 3 different SMILES strings.
+We investigate the required number of message passing layers, as well as using pre-trained weights from an end-to-end trained model.
+GNN layers Weight init Trainable Accuracy (GEOM-QM9) Accuracy (Toy Data)
+0 random trainable 0.887 0.333
+1 random trainable 0.999 0.666
+2 random trainable 1.000 1.000
+2 random frozen 0.980 1.000
+2 pre-trained frozen 1.000 1.000
+Figure A6: Misclassified exampled for SMILES classification experiment.
+We randomly pick 3 examples, which are misclassified by a classification head without any GNN layers.
+We show that GNN layers are essential for correct classification of isomers.
+Appendix L Analysis of Sampling Trajectories
+Figure A7: Comparison of molecular generation with node-only versus node- and pair-wise (all-pair) conditioning on GEOM-QM9.
 Each row shows a prior sample, sampling trajectories, and final generated structure for both models; pairwise conditioning preserves bonds from the conditioning graph $\mathcal{G}$.
 Fig.
-A8 compares the generative performance of DiTMC models trained on the GEOM-QM9 dataset under two different conditioning strategies: (1) node-only conditioning and (2) node plus pairwise conditioning.
+A7 compares the generative performance of DiTMC models trained on the GEOM-QM9 dataset under two different conditioning strategies: (1) node-only conditioning and (2) node plus pairwise conditioning, where we use the all-pair conditioning based on geodesic graph distances.
 Each row in the figure corresponds to one example molecule selected from the test set.
@@
 Within each row, we display a sequence of images: the initial prior sample, followed by the intermediate trajectory of the ODE sampling process over 50 sampling steps for the node-only conditioned model, and the resulting final structure (with predicted bonds rendered in yellow).
-This sequence is repeated for the node- and all-pair-conditioned model, allowing a side-by-side visual comparison of the generation dynamics and final outputs.
+This sequence is repeated for the node- and pair-wise conditioned model, allowing a side-by-side visual comparison of the generation dynamics and final outputs.
 The results reveal a consistent pattern: models trained with node-only conditioning fail to preserve bonding patterns from the conditioning graph $\mathcal{G}$.
 This manifests as bond stretching or atom permutation in the final structure.
-In contrast, the pairwise-conditioned model produces geometries that adhere more closely to expected chemical structure and the given bonds.
+In contrast, the model, that is conditioned on pairwise geodesic distances, produces geometries that adhere more closely to expected chemical structure and the given bonds.
 We note that if atoms are simply permuted by the model using only node conditioning, the generated structure might still be valid in terms of the combination of generated 3D positions and atom types.
@@
 Our findings still underscore the importance of incorporating both node-level and pairwise features in molecular generative models, in particular when agreement with the given conditioning on a bond graph is essential.
-Appendix L Visualization
+Appendix M Visualization
 Fig.
-A9, Fig.
-A10, and Fig.
-A11 provide a visual comparison of conformers generated by MCF, ET-Flow, and DiTMC against the corresponding ground-truth reference conformers for the GEOM-QM9, GEOM-DRUGS and GEOM-XL datasets, respectively.
+A14, Fig.
+A15, and Fig.
+A16 provide a visual comparison of conformers generated by MCF, ET-Flow, and DiTMC against the corresponding ground-truth reference conformers for the GEOM-QM9, GEOM-DRUGS and GEOM-XL datasets, respectively.
 For each dataset, we randomly select six reference conformers from the test split and generate conformers using each method.
 Finally, we apply rotation alignment of the generated conformers with their corresponding reference conformer.
+Appendix N Code and Data Availability
+The code and data to reproduce the main results of this paper, can be downloaded from here: https://doi.org/10.5281/zenodo.15489212, or via github: https://github.com/ML4MolSim/dit_mc.
+Figure A8: Coverage (COV) for precision (“-P”) and recall (“-R”) as a function of RMSD threshold $\delta$ for DiTMC+aPE and other state-of-the-art methods on GEOM-DRUGS.
+The vertical dashed line denotes the RMSD threshold $\delta=0.75$ commonly used for evaluation on the GEOM-DRUGS dataset.
+Figure A9: Coverage (COV) for precision (“-P”) and recall (“-R”) as a function of RMSD threshold $\delta$ for DiTMC+rPE and other state-of-the-art methods on GEOM-DRUGS.
+The vertical dashed line denotes the RMSD threshold $\delta=0.75$ commonly used for evaluation on the GEOM-DRUGS dataset.
+Figure A10: Coverage (COV) for precision (“-P”) and recall (“-R”) as a function of RMSD threshold $\delta$ for DiTMC+PE(3) and other state-of-the-art methods on GEOM-DRUGS.
+The vertical dashed line denotes the RMSD threshold $\delta=0.75$ commonly used for evaluation on the GEOM-DRUGS dataset.
+Figure A11: Average minimum RMSD (AMR) for precision (“-P”) and recall (“-R”) as a function of wall clock time per generated conformer.
+For each model, markers from left to right correspond to an increasing number of sampling steps during generation.
+Here, we follow Refs.
+[hassan2024etflow, wang2023swallowing] and use 5, 10, 20, and 50 sampler steps.
+Figure A12: Average minimum RMSD (AMR) for precision (“-P”) and recall (“-R”) as a function of wall clock time per generated conformer.
+For each model, markers from left to right correspond to an increasing number of sampling steps during generation.
+Here, we follow Refs.
+[hassan2024etflow, wang2023swallowing] and use 5, 10, 20, and 50 sampler steps.
+Figure A13: Average minimum RMSD (AMR) for precision (“-P”) and recall (“-R”) as a function of wall clock time per generated conformer.
+For each model, markers from left to right correspond to an increasing number of sampling steps during generation.
+Here, we follow Refs.
+[hassan2024etflow, wang2023swallowing] and use 5, 10, 20, and 50 sampler steps.
 Figure A14: Comparison of conformers generated by MCF, ET-Flow, and DiTMC against ground-truth reference conformers from GEOM-QM9.